A Complete Bibliography of Publications in International Journal of Image and Graphics

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: https://www.math.utah.edu/~beebe/

13 October 2023 Version 1.61

Title word cross-reference

1 [PPS21, WZW07]. 180^{∘} [HZG07]. 2 [DPR18, DPR20, EAK^{+}06, HU04, KVSM11, KNN10, KNP23, RR05, SSF23, TMSS04, YPYW04]. 3 [BYNZ05, BME07, BH10, DDSP13, CCL^{+}07, CH08, DPR20, GCO^{+}06, HKP^{+}19, HSEI03, HFI15, HZH15, HSCW01, HN02, HWH01, HU04, JKM21, Käm08, KZCP07, KKW06, LWT02, LGS^{+}01, MMP^{+}14, MII17, MA12, NVN^{+}12, NBB05, PW05, PFQ^{+}15, PBHM03, SAA^{+}07, SH20, Sha22, SLN04, SW05, SD01, SL09, TSY^{+}03, TCB06, TMB04, VDPT19, WSO^{+}03, YPYW04, YCL13, YF20, YMA03, dSMP15]. 3D [KA13]. c [KS20, MA02]. K [FHR23, FR16]. L [DSA15]. N [FR16, FD08]. p [BJ16]. S [KSJ21].

-Based [MCT23]. -Dimensional [MII17]. -Means [KS20, MA02]. -Nearest [FHR23]. -Norm [BJ16]. -Point [FD08]. -Systems [DSA15]. -Transform [KSJ21].

1 [PB23]. 121 [Raj23]. 19 [Aro22, GSSM22].

2 [GSSM22, YCWZ09]. 256 [PPS21].

3-Stage [RRR23].

7 [PK01, WCRW01].

AAM [SG11]. Abdomen [SS16]. Abnormalities [SJ21, SM21]. Abnormality [SMR15]. ABSolute [Lee02].

Accelerated [Bha17, KA13, ZSXL08]. Accelerating [CH08]. Accurate [FRC20, NFP10, SKL21, TMB04]. Achievements [VDKP05]. Acquired [DC05]. Acquiring [LF08]. Action [APC17, FD06, IMM22, ND21, VDPT19]. Actions [GS03]. Activation [BR22]. Active [AAS04, DPR06, GGC18, LQDZ15, SD01, WL12, dSP16]. Activities [WHC13]. Activity [ANC17, KKK22]. Acupuncture [HWL^{+}06]. Acyclic [TA16]. AdaBoost [LYL12]. AdaBoosted [PS22]. Adaptation [GLN17, TT06]. Adaptive [Abd17, BH20, BR22, BJ16, CWFW03, CZS^{+}19, DG09, DC05, EFY^{+}05, GMYL10, JJ22, Liu02a, LLW^{+}11, LSES07, NIT13, PCLH21, PAD11, PS12, RRR23, SSF23, SV20, SV21, SR23a, SU15, SK09, VJ23, WCFT03, YG20, ZWM23, SLBK02]. Addition [LPZ06]. Addressing [CML07]. Adhesion [WZF18]. Adjacent [Far17]. Adult [ZZWG06]. Advance [HXCW13]. Advances [HLB01, NPZ01]. AE [JM22, JM23]. Aerial [JKW^{+}15, SMA^{+}18, ZS06]. Aerially [LLLY18]. Affected [MMSC11]. Affine [CZ01]. Agent [FD06, PFQ^{+}15]. Agent-Based [PFQ^{+}15]. AGF [SKC20]. Aging [SWC07]. Agriculture [Kit18]. AI [CG23]. AI-Based [CG23]. Aid [BR22, SL10]. Aided [SS16]. Album [CHZ^{+}03]. Alexander [VSK18]. Algorithm [AM07, AK09, AK23, ASM07, BGX06, CD22, DAM08, DTJ10, EZ21, EZ22, ETB22, GWZ07, GGC18, GS22, GKS21, GP16, HHCS19, HEB19, HK22, HZZ11, JKW^{+}15, JSW12, KB10, KSG05, KAE21, KGH01, KM22, KKK22, LK23, LPZ06, LYP^{+}08, LWZ17, LWCC11, MG21, MCS13, MA21, Mon19, NSS23, Nno19a, PAD11, PDH^{+}21, RR21, RS21, RR11, Ray14, SSF23, SMA^{+}18, SSS04, SPM13, SG21, SK09, SDY19, VG09, VR09, WCXL05, WCB12, WCZC10, ZS12, KL02, WS02].

Algorithm-Based [RR21, SSS04]. Algorithmic [NM20]. Algorithms [BDD23, FZYH08, GA01, II23, JERK23, MM17, MR06, PS01, Sar04, SM14, ZLK05]. Allocation [LZ09]. Alpha [BME07]. Alpha-Shapes [BME07]. Alphabets [UPR11]. Alzheimer [KGSG22, TAJ23]. AM [LZF15]. among [LSZ01]. Amounts [PGW04]. Analysis [AHMS22, ADV02, BR21, BJ13, BB22, Cha08, Dut19, FGT15, GS03, GP16, HJ13, HLB01, HMM05, HL09, LK02, LW07, LGD23, LWY14, LWCC11, MR12, NPZ01, NNNS18, OJD20, PDH^{+}21, Raj20, RPR23, RP23, RK23, RS21, SBDN22, SV22, Sha22, SRG22, SL10, SPC02, SK22, TS01, TAJ23, TSJ20, VR09, WYZR18, XHB08, XTLY07, ZZ18]. Analytic [KZCP07]. Analytical [FD08, FLH^{+}07, HZG07, VK22, Zou08]. Analyzer [GGC18]. Anatomical [MMSC11]. Angiography [HKP^{+}19]. Angle [CYV17, KM21, BJ04]. Angle-Based [KM21]. Angles [MII17]. Animation [AH03, JAPP10, NFP10, YF20, ZLK05]. Anisotropic [BME07, NK20, Nno21, YL16]. ANN [VS13]. Annotation [BYN22, CHZ^{+}03, HTC22, Hel18, LWY14]. Antlion [GS23]. Appearance [WXS^{+}18, dSP16]. Appearance-Based [WXS^{+}18]. Apple [SGTG22]. Application [BH20, CDD05, FZYH08, Kit18, LWZ17, MH10, NVN^{+}12, NESW06, PCV10, RA13, SMA^{+}18, Sha22, SW05, SWZM03, TS01, TTL07, UPR11, XL07, KTK02]. Applications [BBB^{+}05, BG01, ES10, FGT15, LC12, LGD23, Pal01, PPS21, PAD11, Liu02a]. Applied [GRDM19, OJD20, SD01, YCVdS17, dSP16, EK02]. Approach [AS01, ASR20, BM14, BR22, BF11, CWFW03, CYZ09, DK05, DP20, DG09, DG10, Dut19, FLW^{+}12, GSSM22, JRR21a, KZCP07, KK15, KMA22, KP23, LW07, LZF15, LWT02, LG03, MG23, MHN08, MS04, NR20, Pap07, PIG21, PD17, RTP^{+}23, Ran12, RR21, RLKB01, SNL12, SSS04, SDK18, SG13, SRK22a, VG09, WHC13, WXS07, XTLY07, YPYW04, YDZ11, YZ04, ZDT05, KKB02, SL02]. Approaches [AU07, GA23, Mas19]. Approximately [AK09]. Approximating [BAR20, BYNZ05]. Approximation [MKS^{+}09, RR13, SSS04]. Arabic [GKV23, GGTC22a, GGTC22b, OJD20]. Arbitrary [NFP10]. Arc [PDB12, SP20]. Archaeological [BYN22]. Architecture [DPR18, DPR20, FHR23, JM22, JJ22, KM12, Nno17a, PK22, PBK23, PFQ^{+}15, VP22]. Architectures [GGS21, Nno17b]. Area [DPR18, PS22, Ska06]. Areas [SBDN22]. Arguments [RRR23]. Arithmetic [DPR20]. Arnold [JKM21]. Array [LW06]. Arrhythmia [PSJ20, TSJ20]. Art [THBR05]. Arteries [HSEI03]. Artifacts [ASV13]. Artificial [DTJ10, ETB22, GS23, KS20]. Artistic [BSS07, LYT06, LXX^{+}01]. Assembly [WGPC03]. Assess [Sch03]. Assessing [TS02]. Assessment [ESJaO23, Mas19, NMR11, WZPW11, YW11]. Assist [JRR21a]. Assistance [TR10]. Assisted [SG21]. Assistive [KB15]. Asymmetry [GS03]. Asynchronous [GSP^{+}08]. Atlas [MF22]. Atmospheric [PIG21]. Atoms [TS01]. Attack [LY07b]. Attacks [LLS05]. Attention [Raj23, SM23]. Attention-Based [Raj23]. Attentive [IMM22]. Attenuation [FLH^{+}07]. attribute [RFT04]. Augmentation [AP23]. Augmented [MRBL05, PSDM22]. Authentication [AAA^{+}19, BVG22, BB22, CN21, DR20, FR16, KZ09, LPS15, LME23, MRH23, PBC05, SSA13, SYLC05, TK03, Way01]. Author [Ano01, Ano02, Ano03a, Ano04a, Ano05a, Ano06, Ano07a, Ano08a, Ano09, Ano10, Ano11, Ano13, Ano14, Ano15, Ano16, Ano17, Ano18, Ano19, Ano20, Ano21, Ano12]. Autoencoder [JM22]. Automata [PAG^{+}21, BH02]. Automated [BR22, FRC20, KS21d, MB04, MR04, TM04, VRH^{+}23]. Automatic [AK04, AA11, BCB16, BYN22, DPR06, ESJaO23, FB02, GOO^{+}16, HJ13, HEB19, HTC22, HMM05, HL13, JKW^{+}15, Käm08, KO04, KK04, LWZ17, LWY14, MMSC11, PP21, Sar04, SLN04, STB10, SB20, SRBZ01, TG08, WF05, WZF18, YLPC08, ZS06, KKB02]. Automation [SGKA23]. Autonomous [SMA^{+}18, ZLG^{+}07]. AVC [ASM07]. Averaging [PB16]. axial [TMB04].

B [DG10, DS12, SPK08]. B-Spline [DS12, SPK08]. B-Tree [DG10]. Back [BM14]. Background [GGTC22a, Ref15, RCH02]. band [RS22]. Bandwidth [KNN10]. Bangla [PDH^{+}21]. Bank [PGW04]. Bar [WZW07]. Based [AM07, Abd17, AK09, ABK^{+}23, AAL^{+}21, ANC17, Al 06, AIBA14, APC17, AKE19, AD20, ASR20, AL16, BCB16, BMS06, BM14, BG14, BNE18, Bin19, BKKP22, BF11, BSS07, BRAS21, CVdSdS16, CNS^{+}15, CBU23, Cha08, CHT09, CZS^{+}19, CDD05, CG23, CN21, CBCT07, DPSA23, DM13, DR20, DLZ^{+}11, DG09, DS19, DMMS07, DDV03, Dut19, ES10, FPHL06, Far17, FMH12, Fen01, FZYH08, FJS11, FLW^{+}12, GA22, GLX09, GSP^{+}08, GS23, GJ20, GS22, GY06, GJ22, GSSM22, HJ13, HTCZ21, HTC22, HKP^{+}19, HK22, HZH15, Hel12, HHL^{+}08, HLM23, HRH^{+}10, Hu05, HAO08, II15, JS21, JM22, JM23, JZH14, JU13, JRR20, JKM21, KRS17, KDAB19, KT07, KAE21, KH20, KKKH14, KRS19, KS20, KM21, KMA22, KM22, KJJ^{+}23, KNP23, KKK22, LNW06, LK02, LN12, LME08, LSZ01, LPZ06, LY07b]. Based [LSSH12, LZF15, LPL14, LLZ03, LHS05, LZZ07, LLW^{+}11, LZZ^{+}13, LWY14, LME23, LWCC11, MMP^{+}14, dV20, MS17, MG21, MR06, MT13, MBN13, MCT23, ND21, NVN^{+}12, NPZ01, NMM10, NBRR10, NA09, NR20, Nno17a, Nno17b, Nno19a, Nno20, Nno22, Nno23, dOGC^{+}19, PK22, PBK23, PC03a, PS22, PFQ^{+}15, PAD11, PDH^{+}21, PK14, RS16, Raj20, Raj23, RP23, RR21, RSG22, RR11, RR13, Ray14, SSF23, SK15, SGB20, SVM18, SV20, SV21, Sal22, Sal05, SMA^{+}18, SP23, SNL12, SSS04, SR23a, SP11, SSFC09, Sha17, SSP21, SU15, SB22, SK09, SSA13, SR16, SM23, SL09, SCM^{+}23, SBYG04, SRK22b, SRY21, SKL15, TYM12, TYKO14, TB18, TSAS22, TMSS04, TAJ23, THW^{+}07, TA16, TR10, TTK04, TSAS21, VS22, VDBK21, VR09, WCFT03, WF05, WCC^{+}06, WZPW11, WHC13, WYZR18, WD23, WXS^{+}18]. Based [XHB08, XPZ03, XLW12, YCVdS17, YL10, YW11, YL12, YW12, YZ04, YMA03, YHZ06, YL16, ZYNC06, ZSCG07, ZL12, ZZ18, ZLK05, ZZWG06, ZLY09, ZS12, ZWM23, ZSS08, ZDT05, ZYZH05, ZQJ09, dSMP15, Aro22, BH02, EK02, GGS21, HEB19, JSJ22, KKG23, Kon03, LG03, PB23, PL02, RDPP21, Sha22, SS02, SK21, TMB04, VRH^{+}23, VJ23]. Baseline [YPYW04]. Basis [BG14, KS21d]. Bayesian [ZLL^{+}07]. BCS [JM23]. BCS-AE [JM23]. Beam [KTO04]. Beat [ADV02]. Bed [Dut19]. Behavior [FD06]. Belief [RR21]. Benign [KMA22]. Best [HZZ11]. Beta [Sch03]. Beta-Binomial [Sch03]. Between [BBG15]. Bharatanatyam [RDS23]. Bi [SRK22b]. Bi-histogram [SRK22b]. Biased [LSKC05]. Bidiagonal [BR09]. Bidiagonal-Singular [BR09]. Bidirectional [KS21a]. Bilateral [BBG15, Nno20, PG21, SV20, SV21, Sal22]. BiLSTM [PH23]. Binarization [CW05, KS21b, LSKC05, Nno22]. Binary [BAR20, CCAS11, FMH12, GPPA04, LSKC05, NVN^{+}12, dOGC^{+}19, SAS22, ZSCG07]. Binary-Uncoded [CCAS11]. Binomial [Sch03]. Bio [ABK^{+}23, DTJ10, LFZL06]. Bio-Inspiration-Based [ABK^{+}23]. Bio-Inspired [DTJ10]. Bio-Native [LFZL06]. Biogeography [HK22]. Biogeography-Based [HK22]. Biologically [DLZ^{+}11, SLSZ09]. Biomedical [KJJ^{+}23]. Biometric [BB22, CN21, Kon03, LME23, MRH23, Sch03, SP21, VK22, Way01]. Biometric-Based [CN21]. Biometrics [CTI^{+}18, HYC05]. Biorthogonal [PK14]. Bit [CCL^{+}07, Hu05, SPM13]. Bit-Pairing [CCL^{+}07]. Bit-Rate [Hu05]. Bitmapped [KH13]. Bladder [WLW^{+}07]. Blended [SBDN22]. Blind [BF11, DTJ10, KB15, LY07b, MR16, SL11, SPK08, TR10, RMG18]. Block [Abd17, ASM07, BFT11, BVG22, GPPA04, LSZ01, PIG21, SPM13, SM14, SSA13, KL02, JM23]. Block-CS [JM23]. Block-Tree [SPM13]. Block-Wise [SSA13]. Blockchain [SP23]. Blood [EPSF23, HKP^{+}19, JH10]. Blowfish [KB10]. Blurring [LSKC05]. Board [AK14, YZ02]. Bodies [SBH08]. Body [LYL12, LLZ03, PFQ^{+}15, WXS07, YCL13]. Bones [BCB16]. Boolean [dSMP15]. Border [SKL21]. Both [FLW^{+}12]. Bottleneck [DWZ^{+}05]. Bottom [AS01]. Bottom-Up [AS01]. Bound [SL10]. Boundaries [Luk04]. Boundary [BZ10, BT05, DG09, GRDM19, KRS17, TS01, TTK04]. Boundary-Based [TTK04]. Braille [SU15, SDK18, UPR11]. Brain [BR21, BH10, GKS21, RPR23, RSG22, SAA^{+}07, SRG22, TAJ23, TSAS21, VRH^{+}23, WCZC10, ZZY^{+}22]. Branch [BBS^{+}21, Yan08]. Branch-Cut [BBS^{+}21]. Breast [GGS21, HK22, SB22, VDBK21]. Brief [SRG22, SB20]. Brightness [MCT23, Mon19]. Browsers [Zha22]. BSP [WS02]. Buffer [CC06]. Building [LWT02].

Caching [PC03b]. calculus [LPS02].

Caliber [HKP^{+}19]. Calibration [TYM12].

Camera [BT05, DAM08, FD08, IKY02, MBN13, TG08, WZW07, WPGF09, YMG04, HU04]. Camera-Imaged [BT05]. Cameras [JSW12]. Camouflaged [Mon20]. Cancelable [CTI^{+}18]. Cancer [Kha22, KKT^{+}23, SM23, WLW^{+}07]. Capsule [PH23, VRH^{+}23]. Caption [CD22, CS18]. Captioning [NCM^{+}21]. Capture [CMR08, TMSS04]. Captured [LLLY18, MBN13]. Capturing [BSS07, PW07, Sar04]. Carcinoma [GGS21]. Cardiac [ANC17, FLH^{+}07, PSJ20, Sha22]. Cardiomegaly [SR23a]. Carlo [XWB06]. Cascade [DR20]. Case [JRR20, PK22]. Case-Based [JRR20]. Casting [LN12]. Casual [LSMC16]. Cat [JKM21]. Catadioptric [FD08]. Categorization [Hel12, XL07]. CBI [JRR21a]. Cell [KD22, WZF18]. Cells [KMA22, SJ12]. Cellular [PAG^{+}21]. Cerebro [GKS21]. Certainty [SR23a]. Certainty-Based [SR23a]. CG [MH10]. Chained [FHR23]. Chains [AL16]. Challenge [CM23]. Challenges [CG21, GA23, KS21c, SR23b]. Chamfer [Cha08]. Channel [JS21, PCLH21, SYLC05]. Chaos [RR11, SK21]. Chaos-Based [RR11, SK21]. Chaotic [PPS21, SCM^{+}23, TSAS22]. Character [AK14, BNE18, CDD05, CW05, JS18, MBN13, SNL12, SU15]. Character-Based [ES10]. Characterization [AG09, GJ22, MS04]. Characters [CC06, DK05, JLS09]. Chebyshev [VR09]. Check [PGW04]. Chess [CG23]. Chest [Aro22]. Chinese [HWL^{+}06]. Chromatic [DC05]. Chromosomes [AD20]. Circles [AK04]. Circular [Haw07, PDB12, KO04]. CkMeans [HEB19]. Class [SRK22a, TA16, XL07]. Classes [CGP03]. Classical [KS21a]. Classification [ABK^{+}23, Al 06, APC17, AS01, AD20, Aro21, BH10, DP20, DAW01, HK22, HL13, JERK23, JS18, KKT^{+}23, KS20, KMA22, MNB13, MG23, NESW06, Nno19b, PSDM22, PH23, PS22, PAG^{+}21, RPR23, SJ21, SRK22a, SM21, TAJ23, TSZ^{+}17, TA16, TSAS21, VRH^{+}23, SS02]. Classifications [TSZ^{+}17]. Classifier [DSG23, GJ22, JRR21b, Kha22, KS20, PS22, RDKS22, XLW12]. Classifiers [AKE19, DP20, SSR20]. Classifying [VDBK21]. Clickpaths [KHYV02]. Clipping [Raj19]. Clock [GP16]. Closed [MS04]. Clothes [MTVM01]. Clothing [BMH08]. Clothoid [LLL01]. Cloud [KJJ^{+}23, SV22]. Clouds [SCY12]. Cluster [BS04, LN12]. Cluster-Based [LN12]. Clustering [AM07, BMS06, BKKP22, CVdSdS16, HEB19, HRH^{+}10, KSG05, KS20, NIT13, SNL12, VH07, YDZ11, SLBK02]. clusters [MA02]. Cluttered [TTK04]. CMA [SL11]. CNN [Aro22, BR22, ESJaO23, GGS21, MG21, PBK23, VP22, ZZ18]. CNN-Based [MG21, Aro22, GGS21]. Co [KSG05, NA09]. Co-Evolutionary [KSG05]. Co-Occurrence [NA09]. Code [WZW07]. Codebook [AM07]. Coded [BMK19]. Coders [LSZ01]. Codes [SGM11]. Codification [CCL^{+}07]. Coding [Abd17, BFT11, BVG22, CCAS11, CZ01, HHL^{+}08, KM12, PS01, STPG23, SSP21, SCM^{+}23, SWZM03]. Coefficient [Bin19, Sal22, YL16, YW16]. Coefficients [MCT23, PG21, YG20]. Coherence [HKP^{+}19, WZZG04, WLW^{+}07]. Coherent [KGM09]. Coiflet [DMMS07]. Collaborative [CGG^{+}06, GCO^{+}06, TSZ^{+}17, ZLG^{+}07]. Collection [JN14]. Collections [HTC22]. Colon [Kha22]. Colonoscopy [DPC^{+}10]. Color [ASJ20, BS19, CNvW^{+}14, CNS^{+}15, FR16, GCM23, Has21, HHFA22, HRH^{+}10, HL13, LME08, MR16, Nno17b, PAD11, RMG18, TYKO14, TTL07, VG09, VH07, WSL02, WCY11, WZPW11, WG01, YHZ06, ABH^{+}02]. Color-Texture-Edge [YHZ06]. Colored [Ref15]. Colorization [DF11]. Colors [Has21]. Colour [Nno23]. Combination [CHT09, SAS22]. Combinations [RDKS22]. Combined [JBC13, LLS05, PM17]. Combining [BS19, DAM08, LZW06, MMSC11, TMSS04, WYL12]. Commodity [Cai16, YPYW04]. Common [AK09, FD08]. Communication [HFI15, LK23]. Comparative [CYV17, LGD23, OJD20, SV22, WAI21]. Comparing [FGT15, KH13]. Comparison [BBG15, LSZ01, NM20, SM21, RFT04]. Compatibility [Far17]. Compensation [CA06, FLH^{+}07, LSMC16]. Competitive [SM23]. Complementary [WC11]. Complete [Raj20, ZI21, ABH^{+}02]. Complex [GM16, KT07, KKW06, KMSP17, WGPC03, WSZ09]. Component [AHMS22, BR21, SL10]. Components [FGT15]. Comprehensive [LGD23, SAA^{+}07]. Compressed [GBG13, IIU22, LHT^{+}11, PF02, SGB20, WPGF09, YCWZ09]. Compression [ASJ20, BFT11, BM11, CA06, CFC05, CCAS11, GJ20, HS06, Hu05, JM23, KJJ^{+}23, LSZ01, LHS05, PS01, RS16, Raj20, RM21, SP11, SPM13, SM14, SSP21, SK09, SRM09, SR16, SS21, TSAS22, Won09, LPS02]. Compression-Encryption [JM23]. Computation [CHY07, Hos07, LBP05, dV20, Ska06, Ska08, ZQJ09]. Computational [Hel12, MG23, XWB06]. Computed [II23, KKT^{+}23, NNNS18]. Computer [BP23, BB21, DP22, HZH15, JZH14, LF15, PDH^{+}21, SS16]. Computing [GR01, GA01, KJJ^{+}23]. Concept [DG10, RR13]. Concurrent [FGT15]. Conditional [RDPP21]. Confident [YLPC08]. Confocal [GCO^{+}06]. Congestion [PK21]. Congruency [LPS15]. Conic [SK15]. Connectivity [TAJ23, ZZY^{+}22]. Constrained [CS08, MSC13]. Constraints [BAR20, BJ16]. Constructing [WLZ10]. Construction [FJ12]. Constructive [KTK02]. Contactless [ANC17]. Content [AAA^{+}19, ASR20, BKKP22, DDV03, Fen01, HTCZ21, HLB01, KM21, KM22, LZS01, NPZ01, PK01, WCFT03]. Content-Based [ASR20, BKKP22, DDV03, Fen01, HTCZ21, KM21, NPZ01, WCFT03]. Context [Hel12]. Context-Based [Hel12]. Contexts [LZW06]. Continuous [CYZ09, MA12, SNdL18]. Contour [GGC18, GMLASVV18, LQDZ15, MR12, OSN03, PDH^{+}21, SD01, TS02, TTK04]. Contour-Point [GMLASVV18]. Contourlet [JU13, NRK20]. Contours [DPR06, WL12]. Contrast [MT13, MCT23, Nno23, SDY19, TT06]. Contribution [KS21d]. Contributions [NM20]. Control [MM17, SSH17, SKL15, ZDT05]. Convergence [VS22]. Conversion [OCL^{+}01, SDK18]. Convolution [EPSF23, LME23]. Convolutional [AP23, Aro21, HEB19, HLM23, KGSG22, KM22, ND21, PB23, SAG^{+}22, STPG23, SRK22a]. Cooperative [BHM21, KSG05, LGS^{+}01]. Coordinate [ZYZH05]. Coordinates [Ska06, Ska08]. Copula [AL16]. Copy [CG21, Raj23, SAS22, SR23b]. Copy-Move [CG21]. Core [CWFW03, TS01]. Corneal [JAG21]. Corner [AIBA14]. Coronary [HSEI03]. Correcting [TYO02]. Correction [HJ13, LN12, Nno20, Nno23, WAI22]. Correlated [IMM22]. Correlation [BB22]. Correspondence [GMLASVV18, MA21, OSN03, KK04]. Corridors [PK21]. Cosegmentation [ZL12]. Cosine [SK21]. Coupled [NRK20]. Courtesy [PGW04]. Courts [WHC13]. CoV [GSSM22]. Covariance [YCVdS17]. Cover [GA05]. COVID [GSSM22, Aro22, GSSM22]. COVID-19 [GSSM22, Aro22, GSSM22]. CPUs [Cai16]. Crack [YL10]. Cracks [HJ13]. Creasing [RDPP21]. Creating [MH10]. Creation [HMM05]. Critical [BG01, II23, RPR23]. Crop [MG23]. Crow [EZ22, RR21]. Crowd [FRC20]. Crypto [SMSK18, SYLC05]. Crypto-stego [SMSK18]. Cryptography [LK23]. CS [JM23]. CT [Aro22, CZS^{+}19, GBG13, LZZ^{+}13, LQDZ15, NVN^{+}12, SKL21, TSY^{+}03, THW^{+}07]. CT/MRI [LQDZ15]. CTA [CHY07]. Cubicle [SMR15]. Cuckoo [EZ21]. CUDA [KA13]. Cued [JAPP10]. Culling [DDSP13]. Curvature [II15, PDB12]. Curve [BB21, FLW^{+}12, JKM21, LZZ^{+}13, RR05, RR13, Sar04, ZYZH05, Zou08]. Curve-Based [LZZ^{+}13]. Curved [BYNZ05]. Curvelet [BKKP22, NK11, PG21, SS16]. Curves [DG09, DS12, SSS04, Sha17]. Customized [CNvW^{+}14]. Cut [BBS^{+}21, MMSC11, ZLY09]. Cycle [RS21].

D [BYNZ05, BME07, BH10, DDSP13, CCL^{+}07, CH08, DPR18, DPR20, EAK^{+}06, GCO^{+}06, HKP^{+}19, HSEI03, HFI15, HZH15, HSCW01, HN02, HWH01, HU04, JKM21, Käm08, KZCP07, KKW06, KVSM11, KNN10, KNP23, LWT02, LGS^{+}01, MMP^{+}14, MA12, NVN^{+}12, NBB05, PW05, PPS21, PFQ^{+}15, PBHM03, RR05, SSF23, SAA^{+}07, SH20, Sha22, SLN04, SW05, SD01, SL09, TSY^{+}03, TCB06, TMSS04, TMB04, VDPT19, WSO^{+}03, WZW07, YPYW04, YCL13, YF20, YMA03, dSMP15]. D-DTCWT [DPR18]. D-DWT [SSF23]. Dark [PCLH21]. Darwinian [EZ21, EZ22]. Data [AP23, AAA^{+}19, CBCT07, GA05, JN14, KK15, KKT^{+}23, KA13, LS03, LHS05, LZZ07, MMP^{+}14, PW05, Pap07, PCV10, Pou02, RLKB01, SP23, SV22, SMSK18, SDK18, SPC02, TYM12, UHW02, VDKP05, Zha22]. Data-Driven [KK15]. Data-Hiding [VDKP05]. Data-Mining [Pou02]. Database [NIT13]. Databases [PC03a]. Dataset [PK22]. Datasets [SK22]. Datastructure [KGH01]. Daubechies [KT07]. DCT [LY07b, MCT23]. DCT-Based [LY07b]. DCTCW [MCT23]. De-Creasing [RDPP21]. De-Hazing [Nno19a, Nno19b, Nno21]. De-Noising [WD23]. De-Speckling [Sal22]. Dealing [FLW^{+}12]. Decentralized [ZLG^{+}07]. Decision [FD06, JRR20, Lee02, MR06]. Decision-Level [MR06]. Decomposition [BR09, CHT09, GPPA04, LC12, Nno21, RRR23]. decompression [LPS02]. Decryption [KB10]. Deep [ABK^{+}23, BDD23, BR22, BB22, CBU23, DSG23, FHR23, GBP23, GA23, HTCZ21, JJ22, KGSG22, KGSA22, KMA22, KM22, MRH23, MF22, PK22, PD17, PB23, Raj23, RP23, RR21, SAG^{+}22, SR23a, SGKA23, SG21, SRK22a, SCM^{+}23, VS22, VRH^{+}23]. Deep-Learning [BB22]. DeepLearning [KKG23]. Defect [HL13]. Defined [MTH19]. Definition [HZZ11, LLW^{+}11]. Defogging [NK20]. Deformable [AAS04, CC06, DA04, GO04, KS06, NFP10, PBHM03, PR04, SWR^{+}01, TM04, YZ04, ZSS08]. Deformation [CYZ09, WLZ10, ZSAS06]. Deformed [SBDN22]. Deforming [ZW09]. Degraded [BHM21, GMYL10, JLS09, LZZ07, Nno22]. Dehazing [Pal21, PCLH21]. Denoising [AHMS22, KT07, Kit17, Kit18, KS21d, NRK20, PG21, PK14, VSK18, YL16, YW16]. Dense [Cai16, YCWZ09]. DenseNet [Raj23]. Density [AAS04, HK22, MNB13, SB22, VDBK21]. Dependent [ESS03, KKB02]. Depth [HZH15, PSM04, YPYW04]. Depth-Image [PSM04]. Derivative [Nno19a]. Derivative-Based [Nno19a]. Derivatives [SMR15]. Dermatological [JRR21a]. Dermatology [JRR20]. Dermatoscopic [KMA22]. Dermo [WSO^{+}03].

Dermo-Epidermal [WSO^{+}03]. Description [Hel12, PK01, TSAS21, SL02]. Descriptive [DP23]. Descriptor [CVdSdS16, GMLASVV18, GJ22]. Descriptors [AL10, NNNS18, PSDM22, SP11, VDPT19, WCRW01]. Design [AM07, AKS^{+}08, ASR20, CA06, CGG^{+}06, HC10, LF08, LGS^{+}01, PBK23, RS21, Sar04, YCL13]. Designing [EBA15, MTVM01]. Despeckling [KSJ21, SV20, SV21]. Detail [PZZ^{+}01, SWC07, ZLL^{+}07]. Details [STPG23]. Detect [GKS21, GSSM22, MCS13]. Detecting [CML07, WHC13]. Detection [ANC17, AA11, AW05, BH20, Bin19, BME07, CM23, CG21, CZW^{+}17, CMR08, DG09, DTJ10, GKV23, GA23, GGC18, GRDM19, GFZ14, GGS21, HHL^{+}08, HAO08, IKY02, JY22, JH10, KGSG22, KGSA22, KRS17, Kha22, LYP^{+}08, Lie01, MHN08, Mon20, NSS23, NA09, Nno19b, PH23, Pal21, PK21, Raj23, Ref15, RCH02, SSS04, SAS22, SLN04, SJK^{+}21, SG21, SB22, SM23, SR23b, SMR15, TSJ20, VP22, WAI22, WCXL05, WYZR18, Wei13, YL10, YL12, YW12, ZZWG06, ZWM23, dSMP15, KO04, KK04]. Detector [AIBA14, KS10]. Detectors [JC01]. Determination [BBB^{+}05]. Devanagari [JS18, PD17, SNL12]. Developed [II23]. Developing [DPC^{+}10, Dut19]. Development [BFT11]. Deviation [LXX^{+}01, RFT04]. Device [AK15, CMR08, MBN13, Sch03]. Device-Based [MBN13]. Devices [Nno19a]. Dewarping [WAI21]. Diabetic [GGC18, NM20, NSS23, SG21]. Diagnoses [JRR21a]. Diagnosis [CBU23, JAG21, JRR21b, KKT^{+}23, NM20, SR23a, SS16, SJK^{+}21, SSR20, VRH^{+}23, WLW^{+}07, ZZY^{+}22]. Diatoms [FB02]. Dichromats [MTH19]. Different [AKE19, GGS21, JM22, RDKS22, SV22]. Differential [Nno17a, Nno22, Nno23, YL16]. Diffuse [MMSC11]. Diffusion [CDD05, GLN17, NK20, Nno21, PS12, YL16, YW16, ZSS08]. Digit [PM17]. Digital [AAA^{+}19, AS08, AMB16, BJ13, CBCT07, GCM23, GGS19, GP16, HAO08, JKM21, MT13, PB16, RR13, SSS04, SSBD23, SR23b]. Digitally [DC05]. Digits [NESW06]. Dimensional [GOO^{+}16, HSEI03, KGM09, KSJ21, MII17, NG07, WZZG04, WXS07]. Dimensionality [RDS23]. Direct [MMP^{+}14, PB16, TA16, UHW02]. Directed [WCB12]. Direction [YW14]. Directional [KM21, NR20, WZZG04, YW14]. Directions [CG21]. Disaster [KGSA22]. Discovery [KHYV02]. Discrete [JBC13, KSJ21, KDAB19, LLL01, MA12, RSG22, SK21]. Discriminant [BB22, JS18]. Discriminating [ATA14]. Discrimination [BP23]. Discriminative [SC16]. Discriminatory [SWZM03]. Disease [CBU23, FHR23, JRR21b, KGSG22, MG23, MMSC11, SSR20, TAJ23]. Diseases [JAG21, SS16, SJK^{+}21, SGTG22]. Displacement [MH10]. Display [AKS^{+}08, AK14, Käm08]. Displays [BS04, HFI15]. Distance [ASR20, Cha08, GA01, ND21, SB15, Wei14, YW14, ZDT05]. Distance-Based [ZDT05]. Distances [MII17]. Distant [YL10]. Distortion [BFT11, BT05, FLW^{+}12, GS22, MM17, TYO02]. Distributed [CGG^{+}06, DPR20, KM12]. Distribution [GGTC22b, Sch03, SRK22b, MA02]. Divergence [ZZ18]. DMesh [PSM04]. DNA [SCM^{+}23]. Document [BHM21, CW05, DS19, FLW^{+}12, KB15, MHN08, Nno22, RTP^{+}23, WAI21, WAI22]. Documents [AW05, GGTC22b, RDKS22, Ref15, SSGM15]. Domain [AS08, BVG22, BS19, BRAS21, CA06, CFC05, DM13, GA22, GM16, LWLY08, MR16, NK11, NRK20, PBC05, RYLZ06, SGB20, SK21, YCWZ09]. Domains [APC17]. Dominant [DG09]. Dose [CZS^{+}19]. Double [KP23]. down [SS21]. Downsampling [SK09]. DPSO [NSS23]. Drawing [AU07, CM23, PP01]. Dress [RI05b]. Driven [CYZ09, KK15, TC14, WL12]. DSIT [LK23]. DTCWT [BRAS21, DPR18]. DTT [SPM13]. Dual [DL13b, LSMC16, LSES07]. Dual-Planar [LSMC16]. Ductal [GGS21]. Due [GS03]. During [RK22]. DWT [DPR20, LHS05, PBC05, SSF23, SGB20]. DWT-Based [LHS05, SGB20]. DWT/IDWT [DPR20]. Dynamic [CFC05, CMR08, JY22, JZH14, JBC13, LK23, MF22, Nno22, SKL15, THW^{+}07, WCC^{+}06, YKA08]. dynamical [KTK02]. Dysfunction [GS03].

Ear [LME23]. Early [RTP^{+}23]. Earth [Wei14]. EBCD [AIBA14]. Echo [GBG13]. Echocardiogram [CBU23]. ECNN [VJ23]. ECNN- [VJ23]. Edge [AIBA14, BBG15, CM23, GRDM19, HAO08, KS10, Nno22, OCL^{+}01, SDY19, WC11, WCB12, YHZ06, dSMP15, BH02]. Edge-Based [AIBA14, Nno22]. Edge-Directed [WCB12]. Editing [LTJY05, XPZ03]. Editorial [Ano03b, Ano03c, Ano03d, Ano03e, Ano04b, Ano04c, Ben04, LLMR07, Liu02b]. Education [LFZL06]. Effect [BB20, RI05b]. Effective [AK23, Por09, VDPT19, YG20]. Effectiveness [SSN07]. Effects [ASJ20, JN14]. Efficacy [VS22]. Efficient [BDD23, BYNZ05, BHM21, DP20, DG10, DDV03, Hos07, HON03, KNN10, KGH01, KNP23, MZE10, OCL^{+}01, PIG21, SRM09, SR16, TSAS22, VP22, VG09]. Eigen [Bha17]. Eigenspace [RI05a, RI05b]. Eigenspaces [LHW02]. Elastic [WF05, ZYNC06]. Elderly [Has21]. Element [SW05]. Elements [ADV02]. Elevation [PP21]. Elimination [MNB13, PIG21]. ellipse [KO04]. Elliptic [JKM21]. Embedded [BFT11, GLX09, Nno19a, SM14, ZZY^{+}22]. Emission [CFC05]. Emotion [DP23, SH20].

Empirical [RDKS22, SSN07, SSR20, SMR15]. Employing [VK22]. Enabled [KKT^{+}23]. encoded [BH02]. Encoder [AAL^{+}21]. Encoding [GJ20, RS16, Raj20]. Encryption [JM23, KB10, LY07b, PPS21, SP23, SCM^{+}23]. Encryption/Decryption [KB10]. Endoscopic [WLW^{+}07]. Endoscopy [PH23]. Endothelial [WZF18]. Energy [dV20, WL12, ZS12]. Engineering [AU07, HMM05, LK02, LTJY05, RR13, Zha02]. Enhanced [BR21, FTWR17, GCO^{+}06, JS21, JJ22, KS21a, KM21, Won09]. Enhancement [ALT03, BJ13, EPSF23, GGS19, GMYL10, Has21, HLM23, MT13, MCT23, Mon19, Nno17a, Nno17b, Nno20, Nno23, OJD20, RA22, RSG22, SMA^{+}18, SRK22b, PL02]. Ensemble [ABK^{+}23, GLX09, JERK23, SK22]. Entities [AU07]. Entropy [BCB16]. Entry [PK21]. Enumeration [OSN03]. Environment [CGG^{+}06, GSF^{+}03, Kan05, NMM10, Pou02, SV22, TCB06]. Environments [WGPC03, ZLG^{+}07]. Epicardial [HSEI03]. Epidermal [WSO^{+}03]. Epilepsy [ZZY^{+}22]. Equalization [MTH19, SRK22b, ZH21]. Equation [GLN17, dV20, Nno22, Nno23, WSZ09, ZSAS06]. Equation-Based [dV20]. Equations [FHR23, Nno17a]. Equilibrium [DPSA23]. Equivalence [Haw07]. Error [DPSA23]. Estimating [Zou08]. Estimation [AAS04, ASV13, AL16, ASM07, CA06, DAM08, EFY^{+}05, GS03, HKP^{+}19, JR20, KRS19, MNB13, MA21, Nno21, NBB05, PDB12, SR23a, YCVdS17, YPYW04, YCWZ09, ZLL^{+}07, KL02, TMB04]. Euclidean [GA01]. Evaluating [RS16]. Evaluation [AP23, HYC05, Kan02, LPL14, Liu10, MMP^{+}14, NMM10, Raj19, SVM18, SSR20, SMR15]. Evolution [BNE18]. Evolutionary [KSG05, SVM18]. Exact [KZCP07]. Excluding [WSW05]. Exit [PK21]. Expectation [WCZC10]. Expeditious [DSM18]. Experimental [BJ13]. Experiments [LK02]. Exploration [SB15]. Explore [SP21]. Exploring [LN12]. Exponential [HZG07, PB16, WSZ09]. Expression [BDD23, GBP23, SBDN22, WXS^{+}18, ZYNC06, dSP16]. Extending [MTH19]. External [KS21d]. Extra [PS22, Ruk20]. Extraction [ABK^{+}23, AK14, AK15, BYN22, CBU23, CS08, CS18, Dut19, EPSF23, FA06, FTWR17, FZYH08, GGTC22a, GGC18, JBC13, KH11, KVSM11, LWZ17, LZS01, NETM10, PBK23, PM17, Raj23, SG13, SSGM15, SGTG22, SCY12, SKC20, TYKO14, TG08, THW^{+}07, TM04, WSO^{+}03, YLPC08]. Extractor [ZZ18]. Extremely [RM21].

Face [AH03, Bha17, Bin19, CHZ^{+}03, DP23, DA04, FGT15, GLX09, HWH01, HRH^{+}10, JJ22, KDAB19, KZ09, LME23, LWCC11, MR06, MLG^{+}03, NMR11, PW05, PBK23, Ran12, SC16, SWC07, SL09, SWZM03, XTLY07, XLW12, Yan01, YF20, ZSCG07, Liu02a, WSL02]. Faces [ATA14, CML07]. Facial [BDD23, DR20, DA04, GBP23, GS03, MLG^{+}03, SBDN22, SL10, WXS^{+}18, ZYNC06, dSP16]. Factor [CYV17]. Factorization [MKH16]. Family [CHZ^{+}03, KH20]. Farsi [AKE19]. Fashion [YCL13]. Fast [AM07, Abd17, AAL^{+}21, CHY07, CZ01, CC06, FLH^{+}07, GPPA04, HON03, LC12, LLLY18, PSM04, PPS21, PB16, Sal22, SW05, TQJN01, TMB04, ZH21, KL02]. Faster [LS03]. Feature [AD20, ASR20, BYN22, BNE18, Bin19, CZS^{+}19, DR20, DLZ^{+}11, DA04, FPHL06, FZYH08, GGTC22a, GGTC22b, HK22, HON03, IMM22, KRS17, KVSM11, LPS15, MII17, MLG^{+}03, NETM10, NBRR10, PM17, Raj23, RDKS22, SP11, SAS22, SSN07, TMSS04, YHZ06, ZZ18, ZWM23, SS02]. Feature-Based [AD20]. Features [ATA14, AKE19, Bha17, CS08, CZW^{+}17, CMR08, FB02, GBP23, GJ22, GFZ14, HTCZ21, HLB01, HL13, Kha22, LZW06, Mon19, SH20, SAS22, SSR20, SC16, SCY12, VG09, WLZ10, YKA08, dSP16]. Feed [SK09]. Feed-Forward [SK09]. Feedback [CW05, DDV03, KM12, LY07a, LLW^{+}11, WCX02, WCFT03]. Feedback-Based [LLW^{+}11]. Feedback-Free [KM12]. Fetal [RK22]. Fiber [SAA^{+}07, YW14]. Fidelity [BMK19]. Fiducials [DAM08]. Field [DP22, LW06, SKL15, YCWZ09, YDZ11, YMG04]. Field-Based [SKL15]. Fields [CDD05, PKN23]. Film [BJ13]. Filter [AK23, BH20, BBG15, BJ16, Cha08, CDD05, MG21, NIT13, Nno20, PG21, RP23, RSG22, SVM18, SV20, SV21, Sal22, VSK18, YW14]. Filter-Based [CDD05, Nno20]. Filtering [DL13b, GCM23, LME08, NBB05, ZQJ09]. Filters [EPSF23, FMH12, Luk04, SSF23]. Finding [ADV02, AK09, KGH01, VG09]. Fine [VDBK21]. Fine-Tuned [VDBK21]. Finger [BB22, MRH23]. Finger-Vein [BB22]. Fingerprint [ALT03, AK23, BBB^{+}05, CMR08, DS19, GFZ14, Kon03, NA09, RM21, ZYZH05]. Fingerprint-Based [DS19, Kon03]. Fingerprints [LZF15, TK03]. Fingerspelling [SGM11]. Finite [SW05, BH02]. Firefly [ETB22, SMA^{+}18, SM23]. Fisher [JS18, XTLY07]. Fitting [BB21, FLW^{+}12, WL12]. FKNN [KKKH14]. Flexible [BS04]. Flow [FLW^{+}12, KKK22, dV20, YCWZ09, ZQJ09]. Flower [JSJ22]. Fluctuation [YTW11]. Fluid [GKS21, KGM09]. Fluid-Mix [KGM09]. Fly [KS20]. Fly-Based [KS20]. FM [LZF15]. FM-Based [LZF15]. FO [NSS23]. FO-DPSO [NSS23]. Fog [ZH21]. Food [JR20, KKG23]. Footprint [XPZ03]. Foraging [HLM23]. Force [SKL15]. Foreground [TYKO14, YDZ11]. Forensics [KRS19]. Forest [GJ22, ROAK14].

Forestation [Dut19]. Forests [GFZ14]. Forgery [CG21, Raj23, SAS22, SR23b]. Form [MS04, MB04]. Formation [LLMR07]. Forms [Haw07]. Formulation [Nno19a]. Forward [SK09]. Fourier [LPZ06, SD01]. Fourth [Nno23]. Fourth-Order [Nno23]. Foveal [AKS^{+}08]. FPGA [Nno17a, Nno17b]. FPGA-Based [Nno17a, Nno17b]. FPH [HL09]. Fractal [PSDM22, SP11]. Fractional [NK20, Nno19a, Nno20, Nno21, VSK18, YL16]. Fractional-Order [NK20, Nno22]. Fragile [AAA^{+}19, SSA13]. Fragmentation [KH11]. Fragmented [NBRR10]. Frame [ASV13, LHS05, YZ04]. Framelet [Nno23]. Framework [CD22, CBCT07, FZYH08, GSP^{+}08, HEB19, Hel18, LPL14, NRK20, RS22, SPC02, Won09, ZI21, ZLG^{+}07]. Free [DSM18, KM12, MS04, NLC^{+}09, VS22, MB04]. Free-Form [MS04, MB04]. Freeform [KS06]. Frequency [BRAS21]. Fringe [PS22]. Fruit [KS20, LF15, SRK22a]. Frustum [DDSP13]. Full [Pou02, Ref15]. Full-Colored [Ref15]. Full-View [Pou02]. Fully [MR04]. Function [BG14, BR22, CNvW^{+}14, MKS^{+}09, PPS21, SB15, TYO02]. Functional [ETB22, NMM10, ZZY^{+}22]. Functions [dSMP15]. Fundamentals [Way01]. Fundus [ESJaO23]. Fungiform [HL09]. Fused [SR23a]. Fusion [AMR^{+}10, AKE19, BS19, BF11, BRAS21, GM16, HLM23, JU13, JRR21a, KZ09, KP23, MS17, MG21, MR06, NBRR10, Nno21, Nno23, PK22, PW07, PCV10, RP23, RS22, RRR23, VJ23, XLW12]. Fusion-Based [Nno23, PK22]. Future [CG21, LGD23, SP21]. Fuzzy [AM07, DDV03, GWZ07, KS10, KS20, LPS02, MS17, MG21, MA02, NESW06, NIT13, Nno17a, Pal01, Ray14, SVM18, SLBK02, SRM09, SL02, WSL02].

Gabor [FMH12, ZYNC06, ZSCG07]. Gait [Cha08, NETM10]. Game [CG23]. GAN [AP23, RDPP21]. GAN-based [RDPP21]. Gate [LW06]. Gated [FLH^{+}07]. Gauss [YW14]. Gaussian [GMYL10, KS21b, Kit17]. GBDA [LY07a]. Gender [ATA14]. General [CYZ09]. Generalized [LLZ03, XTLY07, BH02]. Generated [AP23, HZG07]. Generating [BMH08, JM22]. Generation [BR04, CD22, DPC^{+}10, Käm08, KK15, LYT06, LXX^{+}01, MSC13, PKN23, SBYG04, ZW09]. Generic [Hel18, SMR15]. Genetic [HK22, JERK23, MM17, RS21, Ray14, SSS04, ZLK05]. Genetically [Kha22, SLBK02]. Geometric [AD20, BS04, FJ12, LLS05, NVN^{+}12, PDB12, RLKB01, SWR^{+}01, VR09, ZH21]. Geometrical [MS04, dSP16]. Geometries [Cai16]. Geometry [Liu10, KKB02]. Gesture [HZH15, JZH14, PFQ^{+}15, RA13, SAG^{+}22]. Glance [BBG15]. Glaucoma [ESJaO23]. GLCM [KKKH14]. GLCM-Based [KKKH14]. Glioma [GJ22]. Global [LSMC16, MS04, SC16, WL12, WYZR18, XWB06, ZL12]. Glomerulus [OKI^{+}03]. GMSC [SB15]. GPPBTF [XLW12]. GPU [AK09, BG14, HZZ11, KA13, LN12, LSSH12]. GPU- [LN12]. GPU-Accelerated [KA13]. GPU-Based [AK09, BG14, LSSH12]. Gradient [Nno21, PIG21, TS01, YL10]. Gradient-Anisotropic [Nno21]. Gradient-Oriented [TS01]. Grading [GS23, MKS^{+}09, SSP21]. Granularity [CZW^{+}17]. Graph [CNS^{+}15, LME08, MMSC11, TAJ23, TA16]. Graph-Based [CNS^{+}15, LME08]. Graph-Cut [MMSC11]. Graphic [BP23]. Graphics [AW05, Ano07b, BG01, CH08, DP22, LGS^{+}01, PXLS08, UHW02, YPYW04]. Graphs [LHW02]. Gray [DA10, GKS21, Hos07, PB16, WZZG04, YL12, YW12, BR21, MCT23]. Gray-Level [DA10, WZZG04, YL12, YW12]. Gray-Scale [PB16]. Grayscale [CDD05].

Grey [SG21]. Growing [DSA15, LWZ17, SB20, KO04]. GRU [PBK23]. GRU-Based [PBK23]. Guest [LLMR07]. Guide [FD06, Lie01]. Guided [AMR^{+}10, BBS^{+}21, Sal22, MA21]. Gumbel [Kit17]. Gurmukhi [JLS09, SGKA23]. GVTI [JM22]. GVTI-AE [JM22]. GWO [MCT23].

H.264 [ASM07]. H.264/AVC [ASM07]. Haar [CZW^{+}17]. Haar-Like [CZW^{+}17]. Hahn [GCB^{+}09]. Hair [LG03]. Hand [AK15, JZH14, KSG05, KZ06, RA13, SAG^{+}22, SDK18]. Hand-Held [AK15]. Handheld [MBN13]. Handling [Pal01]. Handwriting [GGTC22a]. Handwritten [AKE19, DK05, JS18, NESW06, PGW04, PM17, RDKS22, Ref15, SNL12]. Haptic [HWL^{+}06, ZSAS06]. Hardware [CH08, KKW06, Nno17a, Nno17b, VP22, YPYW04, ZSXL08]. Hardware-Accelerated [ZSXL08]. Hardware-Visibility [KKW06]. Harmonic [HS06, Haw07]. Hash [PPS21]. Hausdorff [WB02]. Hazing [Nno19a, Nno19b, Nno21]. Hazy [Nno20]. HD [LLLY18]. HDR [LSSH12, SS21]. Head [ANC17, HFI15, MR04, NVN^{+}12]. Heart [CBU23, RS21, SD01]. Held [AK15]. Hessenberg [SSFC09]. Heuristic [LQDZ15]. Hexagonal [VS22]. Hexrep [VS22]. HHO [KJJ^{+}23]. HHO-Based [KJJ^{+}23]. Hidden [GLX09, MA12, WS02]. Hiding [AAA^{+}19, GA05, LHS05, TYM12, VDKP05]. Hierarchical [BZ10, SPM13, SM23]. Hierarchies [LSES07]. High [BM11, DPR18, FTWR17, HZZ11, MR04, RM21, SG13]. High-Definition [HZZ11]. High-Quality [BM11, MR04]. High-Resolution [FTWR17]. Hindi [SU15]. Histogram [GGTC22a, HL13, MTH19, SU15, YDZ11, ZH21, SRK22b]. Histogram-Based [SU15]. Histograms [YHZ06]. Histopathology [GGS21]. Historical [BHM21]. HMB [GGTC22a]. HMM [AKE19]. HOG [SC16, SH20]. Homogeneous [Ska06, Ska08]. Homomorphic [Nno17a, SP23]. Horizon [IKY02]. Hough [CS08, SL02, WCXL05]. HRCT [MMSC11]. HRV [TSJ20]. Hue [II15, MTH19]. Human [APC17, AD20, Aro21, CBCT07, FJ12, FB02, HRH^{+}10, IMM22, KK15, KKK22, LYL12, MLG^{+}03, NETM10, RI05a, SBH08, SFC08, SL10, TYM12, TMSS04, VDPT19, WSO^{+}03, WXS07, Yan01, YCL13, ZLK05]. Human-Interpretable [FB02]. HVS [BBG15, RMG18]. Hybrid [BR21, BBS^{+}21, CNvW^{+}14, CBU23, DR20, DAM08, DPR18, GBP23, HK22, KBL11, KS10, Kha22, KP23, Nno17b, Ran12, RR05, SVM18, SV21, SW05, SG13, VSK18, YG20]. Hyper [ZZY^{+}22]. Hyper-Network [ZZY^{+}22]. Hyperdimensional [SPC02]. Hyperledger [SP23]. Hyperplasia [HL09]. Hyperspectral [AAL^{+}21, DSG23].

ICP [YCVdS17]. Identification

[BB20, FB02, HL09, KGM09, KDAB19, RA22, RDKS22, RM21, Sch03, SKL21]. Identifying [RK22]. IDH [GJ22]. IDH-Based [GJ22]. IDWT [DPR20]. iFACE [HWH01]. Illumination [LSMC16, Nno19b, XWB06, ZWM23]. Illustrative [Cai16]. Image [Abd17, AHMS22, ALT03, AAL^{+}21, Al 06, AK04, AMR^{+}10, AA11, Ano07b, ASR20, AL16, BH02, BM14, BVG22, BR22, BBB^{+}05, BJ16, BP23, BS19, BM11, BHM21, BRAS21, CNS^{+}15, CG21, CD22, CHT09, CZS^{+}19, CDD05, CW05, CCAS11, DM13, DLZ^{+}11, DP20, DA10, DS19, DMMS07, DL13a, DDV03, Dut19, EZ21, ETB22, FPHL06, Far17, FZYH08, GA22, GM16, GWZ07, GPPA04, GY06, GMYL10, GCO^{+}06, GKS21, HTCZ21, HSEI03, HS06, HHFA22, HZZ11, HZH15, Hel12, HHL^{+}08, HRH^{+}10, Hu05, JM23, JKW^{+}15, JR20, KS21b, KZCP07, KKG23, KB15, KD22, KAE21, KH20, KS21c, Kit17, KNN10, KRS19, KS20, KM21, KM22, KJJ^{+}23, KNP23, KP23, KS21d, LK23, LLS05, LW07, LGD23, LME08, LSZ01, LY07a, LY07b, LLTY07, LYL12, LSSH12, LLMR07, LPL14, LXX^{+}01, LLW^{+}11, LWY14, LQDZ15, LSKC05, MG21, MZE10]. Image [MCT23, Mon19, NVN^{+}12, NIT13, NRK20, NCM^{+}21, Nno17b, Nno19b, Nno20, Nno21, Nno22, OLHL08, PSM04, Pal01, Pal21, PCLH21, PG21, PC03a, PS22, PKN23, PCV10, PAG^{+}21, Por09, PK14, PS01, PS12, RS16, Raj20, RA22, RDPP21, RP23, RS22, RR21, RRR23, Ref15, RYLZ06, SNdL18, SSF23, SAG^{+}22, SV20, SV21, Sal22, SMA^{+}18, SP11, SSFC09, SPM13, SM14, Sha17, Sha22, SAS22, SMSK18, SSBD23, SLSZ09, SB20, SK09, SRM09, SSA13, SL09, SL11, SBYG04, SYLC05, SDY19, SS21, TB18, TYO02, TC14, TSAS22, TSZ^{+}17, VS22, VG09, VS13, VSK18, VJ23, WAI21, WAI22, WCX02, WCFT03, WF05, WC11, WCY11, WZPW11, WCB12, WL12, WYZR18, WD23, WXS^{+}18, XHB08, XL07, YG20, Yan01, YL12, YMA03, YHZ06, YL16, YW16, ZH21, Zha02, ZL12, ZZWG06, ZS12, ZLL^{+}07, EK02]. image [PL02, SS02, YZ02]. Image-Adaptive [YG20]. Image-Based [GY06, SBYG04, YMA03, KKG23]. Image-Driven [TC14]. Imaged [BT05]. Imagery [BFT11, KGSA22, SG13, SWR^{+}01]. Images [AK14, AK23, Aro21, Aro22, BMK19, BCB16, BH20, BYN22, BF11, CHY07, CBU23, DPR06, ESJaO23, FR16, FGT15, GKV23, GCM23, GLN17, GGS19, GMYL10, GBG13, GP16, GGS21, GSSM22, HKP^{+}19, Has21, Hel18, Hos07, HAO08, JM22, JSJ22, JU13, JH10, JKM21, KH11, KSJ21, Kan05, KT07, KSG05, KKT^{+}23, KH13, KMA22, LZZ07, LZZ^{+}13, LWLY08, LHT^{+}11, MS17, MR16, MMSC11, MPR09, MKS^{+}09, MBN13, MF22, NK20, NVN^{+}12, NLC^{+}09, NNNS18, Nno23, NBB05, OÖ03, PSDM22, PH23, PBC05, PB16, PBHM03, PB23, Raj19, RA22, RSG22, RM21, Sar04, SSP21, SRG22, SB22, SKL21, SKC20, SR23b, SRK22b, SMR15, TSY^{+}03, THW^{+}07, TM04, TSAS21, VDBK21, WSO^{+}03, WZZG04, WL12, Won09, WG01, WCZC10, YW12, ZLK05, ZS06, EK02]. Imaging [DSG23, HN02, II23, LLMR07, VRH^{+}23, XHB08]. Imbalanced [SK22]. Impaired [PCV10]. Impairment [YTW11]. Imperatives [GA23]. Implementation [AKS^{+}08, ASR20, CA06, GPPA04, KB10, KM12, Nno19a]. Implicit [FG06, TYO02, Zha22]. Implicits [BG14]. Impostor [CC06]. Improve [Has21, LBP05, MF22]. Improved [BM14, DDSP13, CD22, JJ22, KM22, LZF15, LZZ07, MT13, Nno19b, RS21, SV21, Sal22, SP11, SM14, SG21, SK09, VH07, WCB12]. Improvement [DF11, LLW^{+}11, MG21, SSN07]. Improvements [LN12]. Improving [CNS^{+}15, CW05]. Impulse [BH20, GCM23]. Imputation [FHR23]. Incomplete [GWZ07]. Incorporating [BS04]. Increasing [GP16]. Independency [RA13]. Independent [BR21, SDK18, WYL12]. Index [Ano01, Ano03a, Ano04a, Ano05a, Ano06, Ano07a, Ano08a, Ano09, Ano10, Ano11, Ano13, Ano14, Ano15, Ano16, Ano17, Ano18, Ano19, Ano20, Ano21, WZPW11, Ano02, Ano12]. Indexed [APC17]. Indexing [HTC22, KMSP17]. Indian [RTP^{+}23, RDKS22]. indica [GS23]. Indoor [TR10]. Industry [LK02]. Inertial [PCV10]. Inexpensive [YMA03]. Infected [PH23]. Inference [TSP23]. Informatics [LLMR07]. Information [BR04, Dut19, EBA15, Fen01, II15, KM21, KKK22, Lee02, LLMR07, PW07, SG13, VDKP05, WC11, ZSCG07, ZZY^{+}22]. Information-Theoretic [VDKP05]. Infra [Wei13]. Infra-Red [Wei13]. Infrared [BRAS21, JM22, WYZR18]. Initialization [GO04, LQDZ15]. Ink [GGTC22b]. Inpainting [CHT09, FJS11, TSP23, ZI21, ZS12]. Input [JC01, SBH08]. Inputs [ESJaO23]. inSAR [BBS^{+}21]. Insertion [KS21a]. Insight [NESW06]. Inspection [NLC^{+}09]. Inspiration [ABK^{+}23]. Inspired [BBG15, CM23, DLZ^{+}11, DTJ10, FJ12, RMG18, SLSZ09]. Instance [XL07]. Integer [Nno22]. Integer/Fractional [Nno22]. Integer/Fractional-Order [Nno22]. Integrated [AMB16, GJ22, JM23, KP23, RP23]. Integrating [KZ06, SC16, TSZ^{+}17]. Integration [BKKP22, KKK22, LPS15, MF22, SSN07, TCB06]. Intelligence [SVM18]. Intelligent [FD06, HSCW01, MRH23]. Intensity [NBB05, SS21, WF05]. Intentionality [FD06]. Inter [Abd17, SU15, YZ04]. Inter-Frame [YZ04]. Inter-Image [Abd17]. Inter-Point [SU15]. Interaction [MRBL05]. Interactive [CHY07, ES10, HSCW01, KA13, LGS^{+}01, NFP10, OKI^{+}03, TYKO14, WCC^{+}06, WG01, YCL13]. Interest [APC17, SSP21]. Interest-Based [SSP21]. Interface [GSF^{+}03, LF08, SBH08]. Interfaces [STB10]. Interference [GA05]. Intermediate [VG09]. Internal [KS21d]. Interpixel [WLW^{+}07]. Interpixel-Shifted [WLW^{+}07]. Interpolation [BT05, HS06, LC12, LLZ03, MLS^{+}04, MR12, dOGC^{+}19, Ruk20, SSF23, WZZG04, WCB12]. Interpretable [FB02]. interpretation [ABH^{+}02]. Intersection [SK15, Ska08]. Intersections [GR01]. Intra [SRK22a]. Intra-Class [SRK22a]. Intrinsic [MLS^{+}04]. Introduction [Ano05b]. Intuitive [SBH08]. Invariance [ZWM23]. Invariant [AAE11, Bha17, GO04, HHFA22, MZE10, PBK23, SAS22, VR09]. Invariants [FJ12, GCB^{+}09]. Invasive [GS23, GGS21]. Inverse [Haw07]. Inversion [HZG07].

Investigation [JN14]. Invisible

[AAA^{+}19, TYM12]. involving [KTK02]. IoT [LK23]. Iris [FMH12, KKKH14]. ISA [ETB22]. ISA-Radar [ETB22]. Iso [FA06]. Iso-Surface [FA06]. Isocontouring [LNW06]. Isothetic [SP20]. Iteration [DAM08]. Iterative [BM14, BKKP22, KAE21, LZZ07, SK22, SG11].

Jigsaw [GOO^{+}16]. Jitter [PC03b]. Joint [ND21, PG21, RP23]. JPEG [ASV13, ASJ20]. JPEG2000 [BMK19, BFT11, OLHL08]. Julia [WSZ09].

Kalman [Cha08]. Kannada

[AK14, SDK18, UPR11]. Karhunen

[NIT13]. Kernel

[Bin19, HZG07, KS20, LY07a].

Kernel-Based [KS20].

Kernel-Coefficient-Based [Bin19]. Kernels [HEB19]. Key [GA22, SP23]. Key-Based [GA22, SP23]. Keypoint [LYP^{+}08]. Keypoints [GGTC22b]. Kidney [FHR23, OKI^{+}03]. Kinect [HFI15]. Kinematic [HMM05]. KL [ZZ18]. KNN [KKKH14]. Knowledge [MMSC11, YL10]. Krawtchouk [HHFA22]. Kumaraswamy [SRK22b].

L. [GS23]. Labelling [WXS07]. Language [BB20]. Laplacian [PIG21, SWC07]. Lapse [SJ12]. Large [ESS03, HTC22, KHYV02, PC03b, WGPC03, YL12, ZLG^{+}07]. Large-Scale [KHYV02, PC03b, WGPC03]. Largest [AK09]. Latency [PC03b]. Layer [GRDM19]. Layered [OLHL08]. LBP [ROAK14, SMR15, XLW12]. LCD [HL13]. LDA [MR06]. LDA-Based [MR06]. Leaf [SGTG22, TA16]. Learning [ATA14, Ano07b, BDD23, BB22, CBU23, CCL^{+}07, DSG23, GBP23, GA23, GSSM22, HJ13, HTC22, HWL^{+}06, JERK23, JC01, JAG21, JJ22, KGSA22, KH20, KKT^{+}23, KMA22, LY07b, LPL14, MRH23, MF22, PK22, PD17, PSJ20, Raj23, RP23, RK22, RK23, SR23a, SGKA23, SG21, SGTG22, SWC07, SRY21, VS22, VRH^{+}23, XL07, ZWM23]. Learning-Based [CBU23, KMA22, LPL14, RP23, VS22]. Least [TA16]. Legendre [Hos07, SJ21]. Legible [Cai16]. Length [GJ20, Ska06]. Lesion [NR20]. Less [Nno17b]. Level [BJ13, BH10, DA10, Hos07, MR06, MH10, MR12, PZZ^{+}01, ROAK14, SLN04, SWR^{+}01, WZZG04, WCC^{+}06, WCY11, YW11, YL12, YW12, ZL12]. Level-Set [MR12]. Levels [JC01, SS21, WZF18]. License [PDH^{+}21]. Lifting [BVG22, BR22]. Light [AK15, MPR09, YMG04]. Lighting [KRS19, MLS^{+}04]. Lightweight [LK23]. Like [CZW^{+}17]. Limited [II23]. Line [AA11, AK14, BYNZ05, GR01, KVSM11, LBP05, LYL12, Raj19, WCXL05, YL12]. Linear [BBG15, BKKP22, CS08, JS18, LWCC11, WCC^{+}06]. Lines [JLS09, Ref15]. Link [ETB22]. Lion [HLM23]. Listless [SPM13, SM14]. Liveness [GFZ14]. Liver [SSR20, SKL21]. Local [BJ16, FMH12, GJ22, JJ22, KM21, LPS15, LBP05, LZW06, MS04, PIG21, SAS22, SC16, SWC07, WL12, WYZR18, ZSCG07, ZL12, dSMP15]. Localization [DAM08, PP21, SR23b, TR10]. Localized [KNP23]. Locally [DC05]. Locating [YW12]. Location [MLG^{+}03]. Locomotion [KK15]. Locust [CG23]. Loève [NIT13]. Log [Nno17b, SMA^{+}18]. Log-Based [Nno17b]. Logarithmic [MZE10]. Logic [KS10]. Logo [JBC13, LZ09]. look [BJ04]. Loops [LZZ07]. Losing [Mon19]. Loss [LHS05]. Low [AK14, CZS^{+}19, Hu05, MH10, PSDM22, SPM13]. Low-Dose [CZS^{+}19]. Low-Level [MH10]. Low-Resolution [PSDM22]. LSB [DA10]. Lumigraphs [MLS^{+}04]. Lung [KKT^{+}23, SM23, TSY^{+}03]. Lungs [MMSC11]. LWT [TB18].

Machine [Ano07b, GSSM22, HTC22, JERK23, JAG21, KGSA22, KH20, KKT^{+}23, NLC^{+}09, PSJ20, RK22, RK23, SGTG22, TA16]. Machines [YKA08]. Macrotextures [LBP05]. Magnetic [LW07, PW07]. Magnification [ANC17, JS21]. Magnitude [KKK22]. Mainfold [OCL^{+}01]. Makeup [Bha17]. Malignant [KMA22]. Malik [YW16]. Mammogram [BH20, BJ13, SJ21, SM21, SMR15]. Mammograms [MT13]. Mammographic [HK22, MNB13, VDBK21]. Management [CHZ^{+}03, KKG23]. Mangiferad [GS23]. Mango [GS23]. Manipulation [HWL^{+}06, LLW^{+}11, SWC07, UHW02]. Manta [HLM23]. Manuscripts [OJD20]. Map [JKM21, KS06, PPS21, PP21, TYKO14, TSAS22, Liu02a]. Mapped [MH10]. Mapping [HSEI03, JKW^{+}15, KNN10, LW06, LXX^{+}01]. Maps [ND21, SBYG04, ZSXL08, ABH^{+}02]. Marginal [XTLY07]. Marked [GGTC22a]. Marking [KVSM11]. Markov [AL16, CL01, GLX09, MA12, Wei14, YDZ11]. Matched [RS21]. Matching [AL10, BS19, CZS^{+}19, HZZ11, LYP^{+}08, NG07, PCV10, ZYNC06, ZYZH05, HU04, KTK02, MB04, WB02]. Material [KH11]. Materials [BT05]. Mathematical [KS10, Raj19]. Mathematically [Zha22]. Matrix [FZYH08, MKH16]. Matrix-Based [FZYH08]. Matter [BR21, GKS21]. Matting [LGD23]. max [WSL02]. Maximization [WCZC10]. Maximum [VG09]. Mayfly [CG23]. Mean [DM13, GCM23, HXCW13, Mon19, RI05a, ZH21]. Means [CZS^{+}19, KS20, MA02]. Measure [BCB16, DM13, GBG13, YTW11, WB02]. Measurement [BMK19, SSBD23]. Measures [TS02]. Media [DWZ^{+}05, KBL11, KGSA22, NNNS18]. Median [BH20, Luk04]. Medical [AP23, BS19, DPR06, GM16, JSJ22, KBL11, KD22, KT07, LLMR07, LWLY08, NRK20, PK14, RP23, RRR23, SP23, Sha17, SSP21, STB10, SB20, VJ23, WF05, WL12, YZ04, Zha22]. Melanoma [GA23]. Mellitus [NSS23]. Members [KH20]. Memetic [Mon19]. Memory [HON03]. Menstrual [RS21]. Mental [RK23]. Mesh [BZ10, BG01, LC12, LPZ06, MA12, MKH16, MH10, RFT04, TM04, VDPT19, ZW09]. Meshes [HON03, KVSM11, LLL01]. Meshing [PSM04]. Metabolic [CBCT07]. Metaheuristics [RS22]. Metaspread [Aro21]. Method [Bin19, BM11, DDSP13, CNvW^{+}14, DPSA23, EZ21, FPHL06, Far17, FMH12, GKV23, GGTC22a, HTC22, Has21, II15, KD22, KAE21, LPS15, LZZ^{+}13, MII17, MNB13, MKS^{+}09, NG07, Pal21, PM17, PK21, SV20, SK21, SK22, TSY^{+}03, WSZ09, WCY11, WD23, YL12, ZH21, ZLY09, ZS06]. Methodology [AM07]. Methods [ALT03, BSS07, DMMS07, FGT15, KB15, OJD20, RR05, Sal05, SW05, SK22, WCC^{+}06, ZL12]. Metric [DLZ^{+}11, LLTY07, LHT^{+}11, YTW11, RFT04]. Metrics [CNS^{+}15]. MicroRNA [KH13]. Microscope [MPR09]. Microscopic [GCO^{+}06, SRK22b]. Microscopy [KD22, SJ12]. Microsurgical [WCC^{+}06]. Microvessel [WZF18]. Mimic [SL10]. min [WSL02]. min-max [WSL02]. Mini [II15, SMA^{+}18]. Minimization [DPSA23, ZS12]. Minimization-Based [DPSA23]. Minimizing [PC03b]. Mining [KHYV02, LZZ07, Pou02, RLKB01, UHW02]. Minutiae [BBB^{+}05, FTWR17, LZF15, ZYZH05]. Missing [YL12]. Mix [KGM09]. Mixed [AW05, TSAS22, YW14]. mixture [MA02]. Mobile [MRBL05, Nno19a]. Mobile-Embedded [Nno19a]. MobileNetV2 [PH23]. MobileNetV2-BiLSTM [PH23]. Modal [RRR23]. Modalities [VK22]. Model [ABK^{+}23, CBCT07, DLZ^{+}11, GLX09, GBP23, GGC18, Hel12, HMM05, JM23, Kan02, KNP23, KS21d, LG03, MRH23, PH23, Sal05, SG21, SRM09, SKL21, TV03, TSAS21, VRH^{+}23, WYZR18, YF20, YL16, YW16, Zou08, dSP16, HU04, KTK02]. Model-Based [Sal05, TSAS21, LG03]. Modeled [Zha22]. Modeling [CYZ09, DSA15, HKP^{+}19, LWT02, LTJY05, LF15, LFZL06, MS04, PZZ^{+}01, SWR^{+}01, TCB06, THBR05, YCL13, MB04]. Models [AAS04, CWFW03, DA04, ESS03, GO04, LQDZ15, MA12, NFP10, PR04, TSY^{+}03, WGPC03, ZW09, MA02]. Modes [Por09]. Modified [EPSF23, JS21, JSJ22, ROAK14, SL11, VP22, VR09, WCXL05, WB02]. Moisture [PB23]. Moments [GCB^{+}09, HHFA22, Hos07, NVN^{+}12, SJ21, VR09]. Monitoring [TTL07]. Monocular [SFC08, ZLK05]. Monte [XWB06]. Morphable [YF20]. Morphing [LLZ03]. Morphologic [LLZ03]. Morphological [FJS11, GPPA04, dOGC^{+}19]. Morphology [KS10, KD22]. Mosaic [FPHL06, LYT06]. Motion [ADV02, ANC17, ASM07, BR04, BYNZ05, CA06, CZ01, EFY^{+}05, GS03, Kan02, NMM10, NBB05, PW07, RCH02, Sha22, SFC08, SWR^{+}01, TG08, TMSS04, XPZ03, ZWZ08, KL02, TMB04]. Motions [FRC20]. Mountain [VH07]. Mounted [HFI15, JSW12]. Move [CG21, Raj23, SAS22, SR23b]. Movement [SR23a]. Mover [Wei14]. Movies [RTP^{+}23]. Moving [DSM18, JSW12, WPGF09, Wei13]. MPEG [BR04, LHS05, PK01, WCRW01, YCWZ09]. MPEG-2 [YCWZ09]. MPEG-7 [PK01, WCRW01]. MR [GKS21, LW07, MR04, MF22, SRG22, SB22, WCZC10]. MRI [BR21, BS19, BH10, IIU22, LQDZ15, RSG22, TSAS21]. MSER [SKC20]. Mtar [EAK^{+}06]. Mudras [RDS23]. Multi [BF11, CZW^{+}17, FA06, LWCC11, MF22, Nno20, PZZ^{+}01, RS22, RRR23, SSGM15, SM21, SS21, TA16, WAI22, YPYW04, YDZ11, ZLG^{+}07, KL02]. Multi-Approach [BF11]. Multi-Atlas [MF22]. Multi-band [RS22]. Multi-Baseline [YPYW04]. Multi-Class [TA16]. Multi-Granularity [CZW^{+}17]. Multi-Histogram [YDZ11]. Multi-Linear [LWCC11]. Multi-Modal [RRR23]. Multi-Oriented [SSGM15]. Multi-Resolution [PZZ^{+}01, YPYW04]. Multi-Scale [FA06, Nno20, SS21, KL02]. Multi-Server [ZLG^{+}07]. Multi-Type [SM21]. Multi-Warping [WAI22]. Multibiometric [ROAK14]. Multiclass [KKKH14, SK22]. Multifocus [MG21]. Multiframe [LZZ07]. Multilevel [DPSA23, EZ21, EZ22, ZLY09]. Multilinear [FGT15]. Multimedia [Fen01, GA05, LS03, PK01, PC03b]. Multimodal [AAL^{+}21, CN21, HTC22, LME23]. Multiorientation [FJS11]. Multiorientation-Based [FJS11]. Multiple [AAE11, APC17, BJ04, ESJaO23, FJ12, FHR23, FR16, GFZ14, JC01, KBL11, MS17, SP23, SSN07, SGTG22, XL07]. Multiple-Instance [XL07]. Multiplier [IMM22, Nno17b]. Multiplier-Less [Nno17b]. Multiresolution [BG01, BSS07, CA06, LTJY05, LSES07, ZW09]. Multiscale [LYL12, Nno21, PK14, PBHM03, RS16, RRR23]. Multisensor [MRBL05]. Multispectral [JU13, PS22]. Multistructure [FJS11]. Multiwavelet [LWLY08, RM21]. Murals [AMB16]. Mutual [KH13, ZSCG07]. MVI [HTC22]. My [KH20].

Native [LFZL06]. Natural [ASR20, CM23, DK05, GKV23, KGSA22, Raj19, TTK04]. Navigation [HSCW01, Kan05, PCV10, TR10, HU04]. Ncut [RYLZ06]. Nearest [FHR23]. Needle [HWL^{+}06]. Negative [MKH16]. Neighbor [KGH01]. Neighborhood [DA10, HZZ11]. Neighbors [FHR23]. Nested [PXLS08]. Net [RDPP21, TSAS21]. Network [AP23, CG23, DWZ^{+}05, ETB22, FHR23, HLM23, IMM22, KGSG22, Kha22, KS20, KM22, PB23, RR21, SAG^{+}22, SR23a, SRK22a, SM23, SCM^{+}23, TAJ23, TSP23, VS22, VRH^{+}23, ZZY^{+}22, PL02, SS02, WSL02]. Networked [TCB06]. Networks [Aro21, CML07, ESS03, GS23, HEB19, KDAB19, LME23, ND21, NRK20, NCM^{+}21, SJK^{+}21, SBYG04, VDBK21]. Neural [AP23, Aro21, CG23, CML07, ETB22, FHR23, GS23, HEB19, HLM23, KGSG22, KDAB19, Kha22, KS20, KM22, LME23, ND21, NESW06, NRK20, NCM^{+}21, PB23, SAG^{+}22, SR23a, SJK^{+}21, SRK22a, SCM^{+}23, VS22, VDBK21, PL02, SS02, WSL02]. neural-network-based [PL02]. Neuro [SRM09]. Neuro-Wavelet [SRM09]. Newly [WCXL05]. Newton [KAE21, WSZ09]. Next [DPC^{+}10]. Neyname [DSA15]. NGSC [GJ20]. Nine [DPR20]. Nitric [WZF18]. NN [AKE19]. NN-Based [AKE19]. No [DM13]. No-Reference [DM13]. Nodule [TSY^{+}03]. Noise [ASV13, BH20, BGX06, DL13b, Far17, GA05, GCM23, GLN17, HHCS19, Kit17, LWLY08, MA02]. Noising [WD23]. Noisy [AL10, TM04]. Non [CH08, DG09, FLH^{+}07, GS23, LLZ03, MKH16, RMG18, Ruk20, STPG23, Sha22, SPK08]. Non-blind [RMG18]. Non-Extra [Ruk20]. Non-Invasive [GS23]. Non-Negative [MKH16]. Non-Parametric [DG09]. Non-Rigid [CH08, LLZ03, Sha22]. Non-Subsampled [STPG23]. Non-Symmetric [DG09]. Non-Uniform [FLH^{+}07, SPK08]. Nonlinear [CNvW^{+}14, CNS^{+}15, GMYL10, ZQJ09]. Nonlocal [CZS^{+}19]. Nonparametric [AL16]. Nonphotorealistic [Kan05]. Nonrigid [Sha17]. Nonsubsampled [JU13]. Nonuniform [BM11]. Norm [BJ16]. Normalization [DP20]. Normalized [ZLY09]. Notch [RSG22]. Novel [DPSA23, DPR20, GM16, GGTC22b, GGC18, JERK23, KAE21, LYP^{+}08, LHT^{+}11, LWCC11, MHN08, MSC13, NETM10, NR20, PSJ20, SSF23, SV21, SDK18, SB15, WCY11]. Nu [SSH17]. Nuclear [HSEI03]. Nuclei [KD22]. Number [II23, SP20]. Numerically [HZG07]. Nutritional [KKG23].

Object [DM13, DAW01, DG10, DSM18, JSW12, MII17, MA12, Mon20, NK11, NBRR10, PS22, SB15, TYKO14, TSY^{+}03, TTL07, TG08, WPGF09, Wei13, ZSS08, ZDT05, WS02, WB02]. Object-Based [PS22]. object-oriented [WS02]. Objective [GWZ07]. Objects [DTJ10, FG06, OCL^{+}01, SLSZ09, KO04, MB04]. Observation [OKI^{+}03]. Obstacles [DTJ10, IKY02, TMB04]. Occluded [DG10, KD22, WSW05]. Occlusion [II15, PBK23]. Occlusion-Invariant [PBK23]. Occlusions [LZZ^{+}13]. Occurrence [NA09]. Octree [WCC^{+}06]. Octree-Based [WCC^{+}06]. Octrees [DDSP13]. Old [RDPP21]. OMBO [VJ23]. OMBO-based [VJ23]. on-board [YZ02]. On-Line [LYL12]. on-road [HU04]. Online [JAPP10, LZW06, SNL12, TMSS04]. onto [HSEI03]. Opacity [LN12]. Open [LLL01, VDKP05]. Operations [GPPA04]. Operator [DL13b, DL13a, VS22]. Operators [KVSM11, dSMP15]. Opposition [DPSA23]. Optical [CMR08, HKP^{+}19, HC10, KKK22, dV20, PW07, WLW^{+}07, YCWZ09, ZQJ09]. Optimal [ASM07, BFT11, DPSA23, DDV03, LZ09, Luk04, PBK23, SMA^{+}18, SP23, ZW09]. Optimality [MM17]. Optimization [CD22, CCL^{+}07, CG23, GWZ07, HK22, HMM05, HLM23, JS21, MCT23, PCV10, PB23, SG21, SM23, VRH^{+}23, WGPC03]. Optimization-Based [HLM23, JS21, PB23]. Optimization-Tuned [CG23]. Optimized [BR22, DPR18, ETB22, Kha22, KS21d, LWZ17, SRK22a, SRY21]. Optimizer [DPSA23, GS23]. Optimizer-Based [GS23]. Order [NK20, Nno23, YL16, Nno22]. Ordered [BG01]. Ordering [KNP23, LME08]. Organization [DDV03]. organizing [Liu02a]. Orientation [CVdSdS16, CS08, KKK22]. Oriented [KM21, SSGM15, TS01, WS02]. ORT [GJ20]. Orthogonal [AL16, DAM08]. Orthonormal [KSJ21]. Out-of-Core [CWFW03]. Outdoor [Nak01, TQJN01, TR10]. Outdoor/Indoor [TR10]. Outline [AL10, Sar04]. Outlook [SG21]. Ovary [KKG23]. Over-Sampled [LN12]. Overcome [KKG23]. Overcoming [RI05b]. Overlap [VG09]. Overlapped [ZS06]. Overlapping [JLS09]. Oversampling [SK22]. Overview [Nak01]. Ownership [RR11, SP23]. Oxide [WZF18].

Packet [BR09]. Page [AS01]. Paintings [DC05]. Pairing [CCL^{+}07]. Pairwise [AL16, JS18]. Palm [LPS15, RA22]. Palmprint [FTWR17, KZ09, SRBZ01]. Palmprints [KSG05]. PAN [GS22, STPG23]. PAN-Sharpening [GS22, STPG23]. Panchromatic [JU13]. Panels [HL13]. Papillae [HL09]. Parallel [Cai16, DPR20, IIU22, PXLS08, ZSXL08]. Parallel-Split [ZSXL08]. Paralysis [GS03]. Parameter [BAR20]. Parameterization [BZ10, LZZ^{+}13, TS01]. Parameterized [SBH08, Sha22, HU04]. Parameters [DS12, TSJ20, Zou08]. Parametric [DG09, RP23]. Parenchymal [MMSC11]. Parkinson [JRR21b]. Partial [Nno17a, Nno22, Nno23]. Partially [DG10, KB10, ZS06]. Particle [NBB05]. Particles [NBRR10, SKL15]. Particles-Based [NBRR10]. Partitioning [SM14]. Pass [CWFW03]. Patch [BMS06, WYZR18]. Path [BYNZ05]. Paths [LSES07, SKL15]. Pattern [CL01, CBU23, GJ22, JJ22, PBK23, Ray14, SAS22, SMR15, TYM12, ZS06]. Patterns [BB22, FMH12, KM21, PF02, RA22, ZSCG07]. PC [BS04]. PC-Cluster [BS04]. PCA [MR06]. PCNN [RRR23]. PDE [SWR^{+}01]. Peak [VG09]. PECSI [Won09]. Pedestrian [CZW^{+}17, TR10]. Peeling [MR04]. People [Has21, TTL07]. Pepper [HHCS19]. Perceived [Has21]. Perception [CG21, LK02, MRBL05, TYM12]. Perception-Based [LK02]. Perceptual [GSF^{+}03]. Perceptually [EFY^{+}05, Won09]. Perceptually-Enhanced [Won09]. Performance [AP23, HYC05, Luk04, Mas19, OJD20, Raj19, Raj20, RPR23, Sch03, SK22, TS02]. Performing [BBG15]. Periocular [LME23]. Perona [YW16]. Persian [DSA15, PM17]. Person [LME23, RA22]. Personalized [AH03]. Perspective [FD08, FLW^{+}12]. Perspective- [FD08]. PET [BS19, KZCP07]. PET-MRI [BS19]. Phase [BBS^{+}21, LPS15, PG21, SRY21]. Phones [WZW07]. Photo [CHZ^{+}03, Nak01]. Photo-Realism [Nak01]. Photographic [BP23]. Photography [SMA^{+}18]. Photorealistic [BP23]. PHY [GJ20]. PHY-NGSC-Based [GJ20]. Pictures [ADV02]. Piecewise [DS12]. Pipelined [KB10]. Pipelines [PXLS08]. Pixel [DA10, RR21, Ruk20, SSA13]. Pixel-Wise [SSA13]. Pixels [Far17]. Places [PK21]. Planar [FD08, LSMC16]. Plane [AK09]. Planning [WGPC03]. Plant [THBR05, TA16]. Plate [PDH^{+}21]. Platform [HZZ11]. Platforms [JSW12]. Plenoptic [BYNZ05]. Plot [ES10]. Plot/Character [ES10]. Plot/Character-Based [ES10]. Point [AK09, BAR20, FD08, GSP^{+}08, GMLASVV18, Liu10, SU15, SCY12, TV03]. Point-Based [GSP^{+}08]. Point-Sampled [Liu10]. Points [APC17, BME07, DG09, KS06, MII17, OÖ03]. Poisson [ZSAS06]. Pollination [JSJ22]. Pollination-based [JSJ22]. Polycystic [KKG23]. Polygon [BZ10, MCS13, Raj19]. Polygonal [ESS03, OCL^{+}01, RR13, SSS04]. Polygonization [PB16]. Polyhedral [TV03]. Polynomial [DS12]. Polynomials [AL16]. Pooling [YW11]. Poor [PP21]. Porous [NNNS18]. Pose [DAM08]. position [TMB04]. Positioning [SNL12]. Positron [CFC05]. Possessing [SDY19]. Post [BFT11]. Post-Compression [BFT11]. Postal [SGKA23, SGM11]. Posture [RI05a, ZLK05]. Practical [LWT02, Won09]. Practicing [JAPP10]. Practitioner [Lie01]. Pre [MZE10, OLHL08]. Pre-Shifted [MZE10]. Pre-Specifiable [OLHL08]. Precise [CD22, DSM18, KTO04]. Prediction [Aro22, FHR23, PDH^{+}21, RTP^{+}23, RR21, VS13, YL12, YZ04]. Prediction-Based [RR21]. Preface [ATDT13, Ano08b, STSI15]. Prefix [PC03b]. Pregnancy [RK22]. Pregnant [RK23]. Preprocessing [KKW06, NIT13]. presence [MA02]. Presentation [Por09]. Preservation [BBG15, STPG23, SDY19, ZLL^{+}07]. Preserve [BM11]. Preserved [PG21]. Preserving [BAR20, HON03, KD22, MCT23, Mon19, SSF23, TSAS22, BH02]. Principal [AHMS22, SL10]. Print [BT05, LPS15, TYM12]. Print-And-Scan [TYM12]. Printed [JLS09, RDKS22]. Prior [CZS^{+}19, PCLH21, YL10]. Privacy [TSAS22]. Privacy-Preserving [TSAS22]. Probabilistic [CS08, Nno20, RP23, TTK04]. Probabilities [NA09]. Problem [AAE11, FD08, KGM09, HU04]. Problems [CML07, VDKP05]. Procedure [LY07b]. Procedures [AMR^{+}10]. Process [CL01, HJ13, KM22]. Processing [CDD05, DPR20, DMMS07, GCO^{+}06, HJ13, KB15, KH20, LSSH12, LLMR07, LQDZ15, MRBL05, MPR09, Pal01, PSJ20, SNdL18, SSBD23, SLSZ09, VS22, WCFT03, Yan01, YZ02]. Professionals [RK23]. Profile [XHB08]. Profiles [TS01]. Progenitor [SJ12]. Programmable [LW06]. Progressive [Liu10, MNB13]. Prohibited [WHC13]. Projection [AKS^{+}08, BM14, LSMC16]. Projections [GWZ07, II23]. Projective [Ska08, WSW05]. Propagation [DPR06]. Properties [BAR20, SP20]. Property [LPZ06]. Proportional [SBH08]. Prostate [MF22, SLN04]. Prostheses [DHA10]. Protection [RR11, SP21]. Prototyping [PZZ^{+}01]. Provable [MCS13]. Pruning [CG23]. Public [LPZ06, PK21]. publications [Zha02]. Pulse [NRK20]. Punched [SDK18]. Purpose [Dut19]. Push [DWZ^{+}05]. Puzzles [GOO^{+}16].

Quad [OCL^{+}01]. Quadratic [SSH17]. Quadric [SK15]. Quadtree [KGH01]. Quality [BBS^{+}21, BM11, DM13, DLZ^{+}11, GBG13, GP16, LLW^{+}11, LHT^{+}11, MR04, NMR11, OLHL08, PP21, WZPW11, YW11]. Quality-Guided [BBS^{+}21]. Quantification [WZF18]. Quantization [ASJ20, BMS06, Hu05, KJJ^{+}23, SRM09, SSA13, Yan08]. Quantization-Based [SSA13]. Quantizer [Abd17]. Quartic [DS12]. Quaternion [HHFA22, WZPW11]. Queries [KKW06]. Query [LS03, SSN07].

R [JRR21a]. Radar [ETB22, Käm08]. Radial [BG14]. Radiogenomic [GJ22]. Radiographic [GJ22]. Radon [Haw07, HZG07]. Random [GJ22, GFZ14, Kit17, PKN23, ROAK14, SK22, YDZ11]. Range [DMMS07, NBB05, PW05]. Ranking [FGT15]. Rat [VRH^{+}23]. Rate [BFT11, Hu05, LZ09, RS21, SPM13]. Rate-Distortion [BFT11]. Ratio [BM11]. Ray [GBG13, GSSM22, HLM23, LN12, LWLY08, SK15, TSY^{+}03, BCB16]. Ray-Quadric [SK15]. Rayset [GY06]. RBFNN [VS13]. Re [BSS07]. Re-Using [BSS07]. Reaction [ZSS08].

Reaction-Diffusion [ZSS08]. Reading [WZW07]. Real [ATA14, BMK19, BMS06, DDSP13, GSP^{+}08, HZH15, JZH14, JSW12, NMM10, PDH^{+}21, SNdL18, SD01, YPYW04, Zha22, YZ02]. Real-Rime [Zha22]. Real-Time [BMK19, BMS06, DDSP13, GSP^{+}08, HZH15, JZH14, JSW12, NMM10, PDH^{+}21, SNdL18, YPYW04, YZ02]. Realism [DF11, Nak01]. Realistic [CC06, LSMC16, THBR05]. Reality [ES10, GSP^{+}08, KMSP17, LFZL06]. Reasoning [JRR20]. Reassembly [GOO^{+}16]. Recognition [AU07, ATA14, AA11, AKE19, BDD23, BNE18, Bha17, BB20, CL01, DP23, DG10, ETB22, FMH12, FTWR17, FZYH08, GGTC22a, GLX09, GBP23, HHFA22, HZH15, IMM22, II15, JS18, JZH14, JJ22, KKKH14, KKK22, LZZ^{+}13, LWCC11, MRH23, MII17, MR06, MBN13, NETM10, ND21, Pal01, PGW04, PW05, PBK23, PM17, PDH^{+}21, PD17, RI05a, Ran12, RA13, Ray14, SAG^{+}22, SBDN22, SNL12, SH20, SLSZ09, SC16, SWZM03, SGM11, SP20, TSY^{+}03, TA16, UPR11, VP22, WXS^{+}18, XTLY07, XLW12, YKA08, ZYNC06, ZSCG07, dSP16, BJ04, Liu02a, MB04]. Recognizing [PP01]. Reconstructing [LZF15]. Reconstruction [CHT09, CCL^{+}07, DP22, FLH^{+}07, GWZ07, HSEI03, HN02, II23, KZCP07, KS21c, KS06, Pap07, PFQ^{+}15, SK15, SL09, TMSS04, ZLK05]. Recovery [PG21, PBC05, SSA13]. Rectification [FLW^{+}12]. Red [Wei13]. Reducing [GCM23]. Reduction [Far17, GLN17, GS22, PAD11, RDS23]. Redundancy [SR16]. Reference [BR09, DM13, LSZ01, NLC^{+}09]. Reference-Free [NLC^{+}09]. Reference-Watermarking [BR09]. Refinement [Nno19b, SG11]. Reformative [WCZC10]. Refractions [KTO04]. Region [Al 06, BJ16, DG09, Kit18, LWZ17, LG03, LME23, MT13, Nno19b, Pal21, SSP21, SB20, SKC20, WC11, YLPC08]. Region-Based [Al 06, MT13]. Region-Wise [BJ16]. Regional [BS19]. Regions [MKS^{+}09, YL12]. Registration [AK04, BS04, CYV17, CH08, HSEI03, JKW^{+}15, KAE21, MZE10, NVN^{+}12, PBHM03, Sha17, Sha22, TYO02, WF05, ZS06, HU04]. Regularity [HHL^{+}08, SSF23]. Regularity-Preserving [SSF23]. Regularization [WD23, BH02]. Regularizers [SWR^{+}01]. Rehabilitation [NMM10]. Reinforced [TT06]. related [Zha02]. relation [LPS02]. relations [LPS02]. Relationship [DG10, MII17, ZYZH05]. Relative [SNL12]. relaxation [KTK02]. Relevance [CGP03, DDV03, LY07a, WCX02, WCFT03]. Reliability [Kan02]. Reliable [Lie01]. Relief [LW06]. Reminiscent [RDPP21]. Remote [GSP^{+}08, LGS^{+}01, TK03, YZ02]. Removal [ASV13, BH20, BT05, HHCS19, JY22, LLS05, LWLY08, Ref15, ZH21, WS02]. Removing [BGX06]. Rendering [BS04, Cai16, DDSP13, CC06, ESS03, GY06, LSZ01, MMP^{+}14, PXLS08, SW05, TQJN01, THW^{+}07, WSW05, WCC^{+}06, YMG04, ZDT05]. Rendering-Based [MMP^{+}14]. Repainting [AMB16]. Replicas [CMR08]. Replication [SV22]. Repositories [PC03b]. Representation [AHMS22, BYNZ05, BRAS21, EAK^{+}06, GS22, HHFA22, LZS01, SL10, TSZ^{+}17, WCFT03, ZYZH05]. Representing [NG07]. Rerendering [LSSH12]. Research [CG21, NM20, PXLS08]. Reshaping [SBH08]. Residual [SJK^{+}21, SRY21, Yan08]. Resistant [LLS05]. resizing [PL02]. RESNET [SJK^{+}21]. Resolution [AK14, BM14, BR22, FTWR17, JS21, NFP10, PSDM22, PZZ^{+}01, SG13, SDY19, SRY21, YPYW04, PL02]. Resonance [JBC13, LW07]. Restoration [AK23, AMB16, BJ16, CZS^{+}19, DC05, KNP23, LZZ07, PS12, SL11, YL12, ZLL^{+}07, BH02]. Results [LBP05]. Retina [DHA10]. Retinal [EPSF23, HKP^{+}19, JH10]. Retinopathy [GGC18, NM20, SG21]. Retrieval [AAE11, ASR20, BKKP22, DS19, DDV03, Fen01, HTCZ21, KM21, KM22, LY07a, LZS01, MA12, PB23, TV03, VDPT19, WCX02, WCFT03, WCY11, WG01, YHZ06, EK02]. Reverse [BSS07, LTJY05, RR13]. Reversible [MM17, MR16]. Review [AU07, ALT03, CGG^{+}06, DSG23, GA23, GGS19, HYC05, IIU22, KGSA22, KKG23, MG23, PKN23, RPR23, SB20, SR23b, VK22]. Revisit [SWR^{+}01]. RFA [AMR^{+}10]. RGB [NK20]. Rhythmic [ADV02]. Rice [MG23]. Rider [CD22, JS21]. Rightful [RR11]. Rigid [CYV17, CH08, KAE21, LLZ03, Sha22]. Rime [Zha22]. Risk [RK22]. River [Dut19]. RLS [SSF23]. Road [SBYG04, HU04, TMB04]. Robot [TCB06]. Robust [AIBA14, AW05, AK14, BR09, CZ01, DG10, EAK^{+}06, II15, JKW^{+}15, KH11, LHS05, MR16, NBRR10, Pal21, PBHM03, RR11, RMG18, SGB20, SVM18, SP11, TTL07, WB02, WYL12]. ROI [LWZ17]. Roles [DP22]. Rose [Zou08]. Rotational [HN02]. Rough [SS02]. Rule [GM16, VJ23]. Ruled [Ref15]. Ruled-Lines [Ref15]. Run [GJ20, MSC13].

SA4 [RM21]. Sail [PB23]. Saliency [TYKO14]. Salp [KM22]. Salt [HHCS19]. Salt-and-Pepper [HHCS19]. Sampled [LN12, Liu10]. Sampling [SS21, XPZ03]. Sampling-Based [XPZ03]. SAR [BJ04, KSJ21]. SARS [GSSM22]. SARS-CoV-2 [GSSM22]. Satellite [BF11, Dut19, KGSA22, KS20, PSDM22, PAG^{+}21, SG13]. Sauvola [EPSF23]. Scalable [HHL^{+}08]. Scalar [MMP^{+}14]. Scale [DL13a, FA06, GCB^{+}09, KHYV02, Nno20, PC03b, PB16, SAS22, SS21, WGPC03, ZLG^{+}07, KL02]. Scale-Invariant [SAS22]. Scale-Space [DL13a]. Scales [LBP05]. scaling [SLBK02]. SCAM [YMG04]. Scan [SKL21, TYM12]. Scanner [STB10, YMA03]. Scans [MR04, WXS07]. Scattered [KS06]. SCCI [RR05]. SCCI-Hybrid [RR05]. Scene [FJ12, GKV23, LSZ01, LZS01, LWY14, SKC20]. Scenes [CM23, KKW06, KMSP17, Nak01, TQJN01]. Scheme [AAL^{+}21, BR09, BAR20, DA10, FR16, GLX09, GJ20, LME08, NETM10, PPS21, PS12, RMG18, SSA13, SYLC05, TSAS22, WCX02, YG20, PL02]. Schemes [Hu05]. Science [KKT^{+}23]. Science-Enabled [KKT^{+}23]. Scissors [HSCW01]. Scope [LGD23]. Scores [BBB^{+}05]. Screen [BJ13]. Script [JLS09, PD17, RDKS22, SGKA23, SDK18]. SDIK [GGTC22b]. Seam [HJ13]. Seamless [BMS06]. Seams [BMH08]. Search [CZ01, EZ21, EZ22, LS03, PC03a, PF02, PB23, RR21, ZS06, KL02]. Searching [LLTY07]. Secret [FR16, GA22]. Secure [SP23, TB18]. Secured [CN21]. Securing [LK23, SMSK18]. Security [JKM21, KBL11]. Seeded [LWZ17, SB20]. Segment [Bha17]. Segmentation [AS01, AK14, AL16, BCB16, BR21, BHM21, BH10, CHY07, CNvW^{+}14, CNS^{+}15, CW05, DPSA23, Dut19, GGS19, GKS21, JSJ22, JLS09, Kan02, KD22, LYL12, LPL14, LG03, MNB13, Mas19, MMSC11, MKH16, MT13, MF22, NSS23, NR20, PK22, PDB12, Pal21, PKN23, PR04, RPR23, RYLZ06, SU15, SB20, SJ12, SWR^{+}01, VG09, VH07, WPGF09, WC11, WL12, WCZC10, YDZ11, WSL02]. Segmented [AD20, NVN^{+}12]. Segments [HLB01]. Selection [BYN22, BNE18, DPSA23, HEB19, HK22, Kan02, SS02]. Selective [GCM23, GLX09, Nno22, PS12]. Self [DL13b, MA21, PBC05, Liu02a]. Self-Dual [DL13b]. Self-guided [MA21]. self-organizing [Liu02a]. Semantic [Al 06, CGP03, FJ12, HLB01, Hel18, PK22, ZWZ08]. Semi [AAA^{+}19, FRC20, HJ13, HTC22, STB10, ZS06]. Semi-Automated [FRC20]. Semi-Automatic [HJ13, HTC22, STB10]. Semi-Fragile [AAA^{+}19]. Semiconductor [LK02, NLC^{+}09]. Sensing [IIU22, YZ02]. Sensitivity [HC10]. Sensor [FD08, YZ02]. Sensors [HC10]. Sensory [VDBK21]. Sentiment [ABK^{+}23]. Sentinel [PB23]. Sentinel-1 [PB23]. Separability [DM13]. Separation [BF11, DSM18, Kan02]. Sequence [ASV13, DPR06, PC03a, SGM11, VS13, ZW09]. Sequences [KKK22, LWT02, NMR11, NK11, SFC08, TTK04, VDPT19, WPGF09, Wei13, YZ04]. Series [LYT06, PF02]. Server [DWZ^{+}05, ZLG^{+}07]. Set [AK09, BME07, BH10, CBCT07, HHFA22, MR12, Ray14, ROAK14, WSZ09, ZL12]. Sets [SLN04, SWR^{+}01, TV03]. SHA [PPS21]. SHA-256 [PPS21]. Shading [SW05]. Shadow [DSM18, JY22, MLS^{+}04, ZSXL08]. Shadow-Free [DSM18]. Shadows [TSY^{+}03]. Shape [BAR20, CVdSdS16, CYV17, DS12, EK02, EAK^{+}06, GMLASVV18, HL13, KZ06, LBP05, LZW06, LTJY05, LFZL06, NNNS18, dOGC^{+}19, SSH17, SB15, SG11, TS01, YCVdS17, ZZWG06, KTK02, SL02]. Shape-Based [CVdSdS16, dOGC^{+}19, YCVdS17, ZZWG06, EK02]. Shapes [AL10, BME07, NG07]. Sharing [FR16]. Sharp [SCY12]. Sharpening [GS22, STPG23]. Shearlet [BS19, STPG23]. Shift [HXCW13]. Shifted [MZE10, WLW^{+}07]. Shirt [HRH^{+}10]. Shock [CDD05]. Shopping [GSF^{+}03]. Shot [KRS17, LZS01]. Showface [AH03]. Shrinkage [KSJ21]. Sidewalk [ZWM23]. Sign [VP22]. Signal [PG21]. Signals [ZWZ08]. Signature [GMLASVV18, LZW06, SB15, SYLC05]. Significance [SP21]. Significant [SSS04].

Silhouette [APC17, Cha08, XHB08]. Silicon [CMR08]. Similar [JS18]. Similarity [Abd17, AK09, CNvW^{+}14, CNS^{+}15, LS03, LLTY07, PC03a]. Similarity-Based [Abd17, PC03a]. Simple [BKKP22, HON03, OCL^{+}01]. Simplification [CWFW03, HON03, TC14]. Simulating [MTVM01]. Simulation [DHA10, LF15, WCC^{+}06, ZSAS06, ZSS08]. Simulations [KGM09]. Simulator [DPC^{+}10, HWL^{+}06]. Sine [PPS21]. Single [BM14, CWFW03, Nno19b, Nno21, Pal21, PCLH21, SDY19, TC14, ZH21]. Single-Image [PCLH21]. Single-Pass [CWFW03]. Singular [AS08, BR09]. Singularity [HHL^{+}08]. Sinogram [CFC05]. Size [Abd17, MSC13]. Skeleton [CYZ09, ND21]. Skeleton-Based [ND21]. Skeleton-Driven [CYZ09]. Skeletonization [DK05]. Skew [AW05, MHN08]. Skin [CYZ09, NR20, SJK^{+}21, WSO^{+}03]. Skinning [CYZ09]. Sky [Mas19, Nno19b, Pal21]. Slices [CS18]. Slope [RP23]. Small [PK22, WYZR18, Wei13]. Smart [MRBL05, YZ02]. Smearing [BHM21]. Smooth [YL12, YW12]. Smoothing [BBG15, KP23, PS12]. Snakes [GO04, TTK04]. Sobolev [PIG21]. Social [KGSA22]. Soft [PK14]. Software [MH10]. Soil [PB23]. Solution [FD08, Haw07, HU04, KH20, Raj19, Zou08]. Solving [WSZ09]. Some [Sar04, SSBD23, WSZ09]. Sort [KS21a]. Sorting [GS23]. Sources [BF11]. Sourcing [Lee02]. Space [DL13a, HN02, LLW^{+}11, OKI^{+}03, RMG18, STPG23, Ska08]. Spaces [MRBL05]. Spam [ABK^{+}23]. Spanning [MSC13]. Sparse [AHMS22, BRAS21, GS22, KS21c, dV20, MKH16, STPG23, SWZM03, TSZ^{+}17, WD23]. Spatial [BRAS21, CFC05, DM13, DG10, Dut19, GA22, KMSP17, YW11, GGTC22b].

Spatial-Temporal [YW11]. Spatio

[CS18, IMM22, ND21, TSP23].

Spatio-Temporal [CS18, IMM22, TSP23]. Specifiable [OLHL08]. Specification [FD06]. Specified [BM11, SKL15]. Speckle [DL13b]. Speckling [Sal22]. SPECT [FLH^{+}07]. Spectral [GS22, HC10, STPG23]. Spectrum [BS19, JM22, LY07b]. Speech [BB20, DR20, JAPP10, SDK18, YKA08]. Speed [CM23, DPR18, Kon03]. Speeding [HZZ11, LYP^{+}08]. Sphere [BG14]. Spherical [KS06]. SPIHT [CCAS11]. Spinal [GKS21]. Spiral [MZE10]. Spline [DS12, LLL01, SSH17, SPK08, MB04]. Split [DPR20, ZSXL08]. Spoof [NA09]. Sporting [WHC13]. Spotting [GGTC22b]. Sprite [BR04]. Square [DAM08]. Squares [TA16]. Squirrel [PB23]. Stable [Haw07]. Stacking [JERK23]. Stage [DPR20, GS22, RRR23, SU15, SK22]. Standard [PK01]. State [THBR05]. Static [CMR08, GFZ14, JZH14, LYL12, NRK20]. Statistical [HHCS19, LW07, MSC13, PR04, TSJ20]. Statistics [KH11, LF08]. Steered [RP23]. Steganalysis [Sal05]. Steganographic [DA10]. Steganography [AAA^{+}19, FR16, GA22, RR21, Sal05, SK21]. stego [SMSK18]. Stencil [CC06]. step [KL02]. Stereo [MA21, SK15, TMB04]. Stereoscopic [CZ01]. Still [SWR^{+}01, Won09]. Stitching [LLLY18]. Stochastic [HJ13, JBC13]. Story [ES10]. Story-Based [ES10]. Storyboard [MSC13]. Straightness [PB16]. Strategies [AAE11, SV22]. Strategy [BNE18, GLX09, YCVdS17]. Streaming [DWZ^{+}05]. Stress [RK23]. Stretch [SRY21]. Stroke [SNL12]. Stroke-Based [SNL12]. Structural [ZZY^{+}22]. Structure [BME07, BG01, DF11, HN02, KH13, NBB05, WCFT03]. Structured [JC01, Yan08]. Structures [KGM09]. Studies [SAA^{+}07, SSBD23]. Study [EZ22, FG06, GSSM22, RDKS22, SSN07, WAI21]. Styles [BSS07]. Stylistic [SSGM15]. Stylometry [JN14]. Sub [WCX02]. Sub-Vector [WCX02]. Subdivision [BAR20, BSS07]. Subdivision-Based [BSS07]. Subsampled [STPG23]. Subsequence [PC03a]. Subspace [FGT15, Kan02, Liu02a]. Substitution [DA10]. Subtitles [RTP^{+}23]. Success [RTP^{+}23]. Suitable [PPS21]. Summarization [NMR11]. Sunscald [LF15]. Super [BM14, BR22, JS21, SDY19, SRY21]. Super-Resolution [SDY19, SRY21]. Superellipsoid [HKP^{+}19]. Superpixels [LYL12]. Supervised [AAL^{+}21, LPL14, PK21]. Support [DG09, JRR20, Lee02, TA16, YKA08]. Suppression [GA05]. Surface [BB21, DP22, FA06, HSEI03, KS06, MS04, Pap07, SLN04, SSBD23, STB10, TQJN01, TM04, WSO^{+}03, WLZ10, XHB08, YMG04, Zha22, MB04, WS02]. Surfaces [LSMC16, LLL01, Sha22, SPK08]. Surveillance [NMR11, NCM^{+}21, PPS21]. Survey [BDD23, BP23, CTI^{+}18, DP23, HXCW13, II23, JAG21, KB15, LGD23, Lie01, Mon20, SLSZ09, SRG22, SL09, WAI21, YCL13]. Sustenance [JR20]. SVD [TB18]. SVM [DP20, JRR21b, SH20]. Swarm [KM22, SM23, VRH^{+}23]. switching [KTK02]. Symbols [PP01]. Symmetric [DG09]. Symmetry [KH13, MCS13]. Syndrome [KKG23]. Synthesis [BMS06, CHT09, YPYW04, ZWZ08]. Synthetic [AP23, HWH01]. System [AMR^{+}10, AW05, AH03, AMB16, BDD23, CBCT07, DR20, DWZ^{+}05, ES10, FJ12, GCO^{+}06, HTCZ21, HFI15, HN02, JY22, JZH14, JAPP10, KKG23, Lee02, LGS^{+}01, MLG^{+}03, MBN13, NCM^{+}21, Nno17b, PXLS08, ROAK14, SBDN22, SAA^{+}07, SMSK18, SGKA23, SCM^{+}23, TCB06, TK03, XHB08, ZYZH05, ABH^{+}02, YZ02].

Systematic [BP23, KKG23]. Systems [CTI^{+}18, DSA15, HMM05, HYC05, Kon03, NESW06, THBR05].

T1 [MR04]. T1-Weighted [MR04]. T2FRF [AK23]. Tackling [AAE11]. Talking [HWH01]. Tampered [SSA13]. Tangible [STB10]. Target [WYZR18]. Task [LLW^{+}11]. Taxonomy [GY06]. Technique [AK14, AK15, BBS^{+}21, BHM21, DAM08, EPSF23, FRC20, KBL11, KT07, KNN10, KJJ^{+}23, MMSC11, MR04, MSC13, dOGC^{+}19, SGB20, SSP21, VH07]. Techniques [BB21, DP23, GGS19, GRDM19, JAG21, JRR20, LGD23, MMP^{+}14, MF22, PKN23, PSJ20, RS16, Raj19, Raj20, RA22, RPR23, RK22, RK23, WAI21, WGPC03, Yan01]. Technologies [Way01]. Technology [KB15, VK22]. Telegraph [GLN17]. Template [LNW06, SP21, SBYG04]. Template-Based [LNW06]. Templates [CN21, ZYNC06]. Temporal [CFC05, CS18, IMM22, ND21, SR16, TSP23, YTW11, YW11]. Tension [BAR20]. Tensor [CYV17, DF11, FGT15]. Tensors [CVdSdS16]. Ternary [GJ22]. Test [Bha17, MSC13]. Text [AW05, AK14, AK15, FLW^{+}12, GKV23, SMSK18, SU15, SDK18, SSGM15, SKC20, YW12]. Text/Graphics [AW05]. Textile [XHB08]. Textural [Kit17, Kit18]. Texture [AAE11, AG09, BMS06, CHT09, CZW^{+}17, HSEI03, HL09, KZ06, LBP05, LW06, LWCC11, PR04, SSR20, VR09, WCRW01, YHZ06]. Texture-Based [LWCC11]. Texture-Mapping [HSEI03]. TFT [HL13]. TFT-LCD [HL13]. Their [TS01]. Them [PP01]. Theorem [TYO02]. Theoretic [SP20, VDKP05]. Theoretical [FZYH08]. Theory [TAJ23]. Theory-Based [TAJ23]. Thermal [JM22, SBDN22]. Three [BAR20, GOO^{+}16, HSEI03, KGM09, WZZG04, WYL12, WXS07].

Three-Dimensional [GOO^{+}16, KGM09, WZZG04]. Three-Point [BAR20]. Threshold [DPSA23, PCLH21]. Thresholded [NRK20]. Thresholding [AW05, EZ21, EZ22, EPSF23, PAD11, PK14, ZLY09, KK04]. Thresholding-Based [PAD11]. Tiled [BS04]. Time [BMK19, BMS06, BG01, DDSP13, GSP^{+}08, HZH15, JZH14, JSW12, NMM10, PDH^{+}21, PF02, RCH02, SNdL18, SD01, SJ12, WD23, YPYW04, YZ02]. Time-Varied [RCH02]. TLBO [GKS21]. Toggle [DL13b, DL13a]. TOM [BG01]. Tomatoes [II15]. Tomography [CFC05, HKP^{+}19, II23, KKT^{+}23, NNNS18, WLW^{+}07]. Tonal [Nno20]. Tongue [HL09, YL10]. Tool [DHA10]. Tools [MRBL05, TCB06]. Topographic [PP21]. Topological [KVSM11, WXS07]. Total [KNP23]. Totally [BG01]. Touching [JLS09]. Tracing [BG14, KTO04, RR05]. Track [SB15]. Trackers [TS02]. Tracking [DAW01, DA04, HXCW13, HSCW01, JY22, JSW12, Mon20, NK11, NBRR10, PFQ^{+}15, SAA^{+}07, SFC08, SJ12, TTL07, TMSS04, TTK04, WYL12, YLPC08, YW14]. Trademark [NIT13, EK02]. Traffic [VP22]. Training [HWL^{+}06, NMM10, PK22]. Trajectory [CVdSdS16]. Transcoding [CGP03]. Transcribing [FRC20]. Transcription [SU15]. Transfer [KMA22, SR23a]. Transform [BR09, BKKP22, CS08, GM16, GA01, Haw07, HZG07, JBC13, JU13, KSJ21, KDAB19, KT07, MS17, NIT13, PK14, RS16, RSG22, RM21, SMA^{+}18, SSFC09, SS16, SAS22, SK21, SRY21, VG09, VR09, YG20, SL02, LPZ06, WCXL05]. Transform-Based [MS17]. Transformation [AK09, SR16, ZYNC06]. Transformations [DL13a, LS03]. Transformer [TSP23]. Transition [Lie01, SP21]. Translation [GCB^{+}09, UPR11]. Transmission [KNN10, OLHL08, SP23]. Tree [DG10, KNP23, MSC13, PS01, SPM13, WCFT03, Yan08, WS02]. Tree-Structure [WCFT03]. Tree-Structured [Yan08]. Trees [PS22]. Tri [JJ22, KM21, WYL12]. Tri-Directional [KM21]. Tri-Tracking [WYL12]. Triangle [HON03, YZ04]. Triangle-Based [YZ04]. Triangular [BZ10, DG10, PS01]. Triangulated [MG21]. Triangulation [BZ10]. Trigonometric [DS12, SSH17]. Trimmed [BH20]. Truncation [BFT11, BVG22]. TRUS [SLN04]. Tumor [RPR23, SRG22, SB22, TSAS21, WZF18]. Tuned [CG23, VDBK21]. Tuning [RI05a]. Turbulence [LZZ07, PIG21]. Twin [TA16]. Twitter [ABK^{+}23]. Two [DS12, GA01, GS22, KSJ21, NG07, SU15, SK22, WCY11, YW11, KL02]. Two-Dimensional [KSJ21]. Two-Level [WCY11, YW11]. Two-Stage [GS22, SU15, SK22]. two-step [KL02]. Type [SM21]. Types [AKE19, JM22].

U [TSAS21]. U-Net [TSAS21]. UAVs [Mas19]. Ulcers [NSS23]. Ultrasound [PBHM03, SV20, SD01]. Uncalibrated [LWT02]. Uncertain [Pap07]. Uncertainty [MMP^{+}14, Pal01]. Uncoded [CCAS11]. Understanding [FJ12, MRBL05]. Underwater [HLM23, Nno23]. Unified [YPYW04]. Uniform [FLH^{+}07, FMH12, SPK08]. Units [BB20]. Unmanned [SMA^{+}18]. Unstable [Haw07]. Unsupervised [AL16, BH10, LPL14, ZWM23]. Unwrapping [BBS^{+}21]. Upper [PFQ^{+}15]. Urban [PS22]. Use [WC11]. Useful [KB15]. User [BM11, DR20, GSF^{+}03, JC01, KHYV02, KZ09, LF08, MRBL05, STB10, SKL15]. User-Specified [BM11, SKL15]. Using [AAE11, AAS04, AHMS22, AP23, AK04, AKE19, AG09, Aro21, ASR20, AL16, BR04, BR21, BVG22, BR09, BJ16, BKKP22, BB22, BH10, BB20, BT05, BSS07, CVdSdS16, CA06, DDSP13, CNvW^{+}14, CNS^{+}15, CYV17, CBU23, CS18, CCAS11, CZ01, CC06, CMR08, CH08, CGP03, DPSA23, DM13, DP20, DA10, DAW01, DA04, DDV03, EFY^{+}05, EZ21, ESJaO23, EPSF23, FB02, GPPA04, GS23, GJ22, GKS21, GFZ14, GGS21, HTCZ21, HSEI03, HFI15, HZH15, HLM23, IMM22, JERK23, JJ22, JBC13, JRR20, JR20, JRR21b, KBL11, KSJ21, KS10, KGSG22, KGSA22, Kha22, KKW06, KKT^{+}23, Kit17, KS06, KH13, KNN10, KKKH14, KZ09, KA13, KM21, LPS15, LSSH12, LZZ^{+}13, LME23, LGS^{+}01, dV20, MRH23, MII17, MNB13, MM17, MG23, MLS^{+}04, MTH19, MSC13, MT13, MR12, NMR11]. Using [ND21, NNNS18, NIT13, NBB05, PK22, PSDM22, PH23, PBK23, PPS21, Pap07, PFQ^{+}15, PAG^{+}21, PK14, PB16, PB23, RTP^{+}23, RS16, RA22, Raj23, Ran12, RP23, RR21, RSG22, RK23, RS21, RRR23, RYLZ06, RM21, ROAK14, SSF23, SAG^{+}22, SBDN22, SK15, SMA^{+}18, SR23a, SH20, Sch03, SPM13, SS16, SAS22, SJK^{+}21, SSP21, SRG22, SG21, SJ21, SRM09, SSA13, SGTG22, SRK22a, Ska08, SKC20, SCM^{+}23, SP20, TSY^{+}03, TV03, TK03, TA16, VG09, VS13, VDBK21, WCFT03, WL12, WCZC10, YPYW04, YKA08, ZH21, Aro22, Bha17, CD22, DPR20, KKG23, NCM^{+}21, PDH^{+}21, PSJ20, RFT04, SP23, SGKA23, VRH^{+}23, VSK18, VJ23, WSL02, WB02]. Utilizing [AS01, MZE10].

V [JS21]. Validation [PDB12]. Value [BR09, MTH19]. Values [PP21]. Vanishing [LBP05]. Variability [RS21]. Variable [Yan08]. Variable-Branch [Yan08]. Variance [DM13]. Variation [FG06, KNP23, Sal22]. Variations [JERK23]. Varied [RCH02]. Various [BDD23, BB20, RA22]. Varying [SSP21]. Vascular [RA22]. VEA [GGC18]. Vector [Abd17, AS08, GWZ07, Hu05, KRS17, KJJ^{+}23, NR20, RTP^{+}23, SRM09, TA16, WCX02, Yan08, YKA08]. Vector-Based [NR20]. Vectorization [AA11, KTO04]. Vectors [Bha17, Kit17]. Vehicle [JY22, PDH^{+}21, SMA^{+}18]. Vein [BB22, MRH23]. Venation [THBR05]. Verification [Kon03, KZ06, LZW06, SRBZ01]. version [WB02]. Vertex [KVSM11]. Vertices [SSS04]. Very [PK22]. Vessel [CHY07, EPSF23, GGC18, HKP^{+}19, HSCW01]. Vesselness [CHY07]. Vessels [JH10]. Via [DF11, LS03, Nno19b, PCV10, STPG23, SSN07, SWR^{+}01, VH07, ZS06, ABK^{+}23, HU04, KMA22, Nno21, PCLH21]. Video [BR04, BKKP22, CVdSdS16, CA06, Cha08, CS18, CCAS11, CZ01, CGP03, DAW01, DSM18, EFY^{+}05, GJ20, HTC22, HLB01, HFI15, HLM23, HRH^{+}10, JS21, KM12, KKK22, LF08, LZS01, LWT02, LHS05, LLW^{+}11, LLLY18, MRBL05, MSC13, NMR11, NPZ01, NK11, NCM^{+}21, PPS21, SGB20, SK21, SR16, SGM11, TG08, TMSS04, TTK04, TSP23, WPGF09, WHC13, YTW11, YW11, ZI21]. Video-Based [Cha08, TMSS04, WHC13]. Videophone [YZ04]. Videos [AAA^{+}19, GBP23, Lie01, LLLY18]. View [DDSP13, ESS03, IKY02, OÖ03, Pou02, YPYW04, KKB02]. View-Dependent [ESS03, KKB02]. Viewpoint [TC14]. Views [FJ12, Käm08, OÖ03, WYL12]. Vijayanagara [AMB16]. Virtual [AK04, BMH08, CGG^{+}06, CC06, ES10, GSP^{+}08, GSF^{+}03, HSCW01, Kan05, KMSP17, LFZL06, NMM10, PZZ^{+}01, SBYG04, TCB06, WGPC03, ZLG^{+}07]. Visibility [KKW06, MTH19, ZDT05]. Visible [BRAS21, CBCT07, JM22]. Vision [BBG15, BB21, DP22, DTJ10, HZH15, JZH14, NLC^{+}09, PDH^{+}21, PCV10, SK15, TR10, TMB04]. Vision-Based [HZH15, JZH14, PDH^{+}21, SK15, TR10, TMB04]. Visual [BB20, DSA15, DHA10, FJ12, FB02, HTCZ21, HLB01, JN14, JC01, Pou02, SSN07, TYM12, WYL12, YLPC08, YKA08]. Visualization [AMR^{+}10, EBA15, FA06, GSP^{+}08, GCO^{+}06, KHYV02, KA13, Lee02, LLMR07, MMP^{+}14, MPR09, Por09, RDS23, SAA^{+}07, SPC02, Zha22]. Visualizations [Cai16]. Visually [PCV10]. Vitality [CMR08]. Vivid [VK22]. Vivo [WLW^{+}07]. VLSI [KB10]. Vocabulary [RA13]. VOI [THW^{+}07]. Volume [Ano03a, Ano04a, Ano05a, Ano06, Ano07a, Ano08a, Ano09, Ano10, Ano11, Ano12, Ano13, Ano14, Ano15, Ano16, Ano17, Ano18, Ano19, Ano20, Ano21, FA06, FG06, LN12, MMP^{+}14, Ska06, TMSS04, THW^{+}07, WSW05, WCC^{+}06, ZDT05, Ano02]. Volumetric [KA13, WSO^{+}03]. Voxels [WSW05]. VR [GCO^{+}06, OKI^{+}03].

Warping [FLW^{+}12, PSM04, WAI22]. Water [TQJN01]. Watermark [KBL11]. Watermarking [AAA^{+}19, AS08, BR09, EFY^{+}05, LZ09, LLS05, LPZ06, LSKC05, MM17, MR16, RR11, RMG18, SGB20, SSFC09, SSA13, SPK08, TB18, YG20]. Watershed [MT13]. Wave [dV20]. Wavelet [BVG22, BR22, BR09, CA06, CHT09, GM16, GGC18, HHL^{+}08, HAO08, JBC13, KRS17, KDAB19, KT07, KNP23, LSZ01, MR16, NA09, Nno21, PK14, Raj20, RSG22, RS21, RYLZ06, SV21, SRM09, SR16, VG09, VDBK21, ZYNC06]. Wavelet-Based [HAO08, KRS17, NA09, Raj20, VDBK21]. Wavelet-Domain [CA06, MR16]. Way [XWB06]. WCFS [AG09]. Weak [MCS13]. Web [KMSP17, Zha22]. Weber [JJ22, KM21]. Weight [AK15]. Weighted [DP20, Luk04, MR04, RS22, SV20, SMR15, TV03, VJ23]. Weighting [MCT23, WCX02, YCVdS17]. White [GKS21]. Who [KH20]. Wide [CYV17]. Wiener [SV21]. Wireless [PH23, SYLC05]. Wise [BJ16, SSA13]. Within [YCWZ09]. Without [Mon19, WZF18]. Wolf [MCT23, SG21]. Women [RK23]. Word [AK14, AKE19, GGTC22b]. Words [DSA15]. World [ATA14]. Worn [SSBD23]. Wrinkles [BMH08]. WSFS [AG09]. WWW [KHYV02].

X [BCB16, GBG13, GSSM22, LWLY08, TSY^{+}03]. X-Ray [GBG13, GSSM22, LWLY08, TSY^{+}03, BCB16].

YCbCr [RMG18].

Zernike [SC16]. ZTR [CCAS11].

References

Alvarez:2011:VLD

[AA11] Miguel Alvarez and María-Elena Algorri. Vectorization and line detection for automatic image recognition. International Journal of Image and Graphics (IJIG), 11(3): 439–470, July 2011. CODEN ???? ISSN 0219-4678.

Aditya:2019:ISF

[AAA^{+}19] B. P. Aditya, U. G. K. Avaneesh, K. Adithya, Akshay Murthy, R. Sandeep, and B. Kavyashree. Invisible semi-fragile watermarking and steganography of digital videos for content authentication and data hiding. International Journal of Image and Graphics (IJIG), 19(3):??, 2019. ISSN 0219-

                                                      4678. URL https://www.
                                                      worldscientific.com/doi/
                                                      10.1142/S0219467819500153.

Abbadeni:2011:TPI

[AAE11] Noureddine Abbadeni, Haikel S.

Alhichri, and Alaa B. Elmasry. Tackling the problem of invariant texture retrieval using multiple strategies. International Journal of Image and Graphics (IJIG), 11(1): 43–64, January 2011. CO-DEN ???? ISSN 0219-4678.

Akrour:2021:FHI

[AAL^{+}21] Leila Akrour, Soltane Ameur, Mourad Lahdir, Régis Fournier, and Amine Nait Ali. Fast hyperspectral image encoder based on supervised multimodal scheme. International Journal of Image and Graphics (IJIG), 21(01):??, January 2021. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467821500078.

Abd-Almageed:2004:ADM

[AAS04] Wael Abd-Almageed and Christopher E. Smith. Active deformable models using density estimation. International Journal of Image and Graphics (IJIG), 4(3):343–??, July 2004. CODEN ???? ISSN 0219-4678.

Abdelwahab:2017:IIS

[Abd17] Ahmed A. Abdelwahab.

Inter-image similarity-based fast adaptive block size vector quantizer for image coding. International Journal of Image and Graphics (IJIG), 17 (3):1750017, July 2017. CO-DEN ???? ISSN 0219-4678.

Ablameyko:2002:CSI

[ABH^{+}02] S. Ablameyko, V. Bereishik, M. Homenko, D. Lagunovsky, N. Paramonova, and O. Patsko. A complete system for interpretation of color maps. International Journal of Image and Graphics (IJIG), 2(3): 453–??, July 2002. CODEN ???? ISSN 0219-4678.

Ainapure:2023:DEM

[ABK^{+}23] Bharati S. Ainapure, Mythili Boopathi, Chandra Sekhar Kolli, , and C. Jackulin. Deep ensemble model for spam classification in Twitter via sentiment extraction: Bio-inspiration-based classification model. International Journal of Image and Graphics (IJIG), 23(04): ??, July 2023. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467823500341.

Arora:2020:GFB

[AD20] Tanvi Arora and Renu Dhir. Geometric feature-based classification of segmented human chromosomes. International Journal of Image and Graphics (IJIG), 20(01):??, January 2020. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467820500060.

Adams:2002:FBA

[ADV02] B. Adams, C. Dorai, and S. Venkatesh. Finding the beat: An analysis of the rhythmic elements of motion pictures. International Journal of Image and Graphics (IJIG), 2(2):215–??, April 2002. CODEN ???? ISSN 0219-4678.

Arivazhagan:2009:TCU

[AG09] S. Arivazhagan and L. Ganesan. Texture characterization using WSFS and WCFS. International Journal of Image and Graphics (IJIG), 9(1): 77–100, January 2009. CO-DEN ???? ISSN 0219-4678.

Arya:2003:PFA

[AH03] Ali Arya and Babak Hamidzadeh.

Personalized face animation in showface system. International Journal of Image and Graphics (IJIG), 3(2):345–??, April 2003. CODEN ???? ISSN 0219-4678.

Abedini:2022:IDU

[AHMS22] Maryam Abedini, Horriyeh Haddad, Marzieh Faridi Masouleh, and Asadollah Shahbahrami. Image denoising using sparse representation and principal component analysis. International Journal of Image and Graphics (IJIG), 22(04):??, July 2022. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467822500334.

AlNachar:2014:REB

[AIBA14] Rabih Al Nachar, Elie Inaty, Patrick J. Bonnin, and Yasser Alayli. A robust edge-based corner detector (EBCD). International Journal of Image and Graphics (IJIG), 14(4): 1450018, October 2014. CO-DEN ???? ISSN 0219-4678.

Alhichri:2004:AIR

[AK04] Haikel S. Alhichri and Mohamed Kamel. Automatic image registration using virtual circles. International Journal of Image and Graphics (IJIG), 4(2):281–??, April 2004. CODEN ???? ISSN 0219-4678.

Aiger:2009:GBA

[AK09] Dror Aiger and Klara Kedem. A GPU-based algorithm for approximately finding the largest common point set in the plane under similarity transformation. International Journal of Image and Graphics (IJIG), 9(2): 287–298, April 2009. CODEN ???? ISSN 0219-4678.

Angadi:2014:RST

[AK14] S. A. Angadi and M. M. Kodabagi. A robust segmentation technique for line, word and character extraction from Kannada text in low resolution display board images. International Journal of Image and Graphics (IJIG), 14(1– 2):1450003, 2014. CODEN ???? ISSN 0219-4678.

Angadi:2015:LWT

[AK15] S. A. Angadi and M. M.

Kodabagi. A light weight text extraction technique for hand-held device. International Journal of Image and Graphics (IJIG), 15(4): 1550017, October 2015. CO-DEN ???? ISSN 0219-4678.

Antony:2023:TFE

[AK23] Joycy K. Antony and K. Kanagalakshmi. T2FRF filter: an effective algorithm for the restoration of fingerprint images. International Journal of Image and Graphics (IJIG), 23(01):??, January 2023. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467823500043.

Arani:2019:HFW

[AKE19] Seyed Ali Asghar Abbaszadeh Arani, Ehsanollah Kabir, and Reza Ebrahimpour. Handwritten Farsi word recognition using NN-based fusion of HMM classifiers with different types of features. International Journal of Image and Graphics (IJIG), 19 (1):??, January 2019. CO-DEN ???? ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467819500013.

Ahlborn:2008:DIF

[AKS^{+}08] Benjamin A. Ahlborn, Oliver Kreylos, Sohail Shafii, Bernd Hamann, and Oliver G.

Staadt. Design and implementation of a foveal projection display. International Journal of Image and Graphics (IJIG), 8(2):243– 263, April 2008. CODEN ???? ISSN 0219-4678.

AlAghbari:2006:RBS

[Al 06] Zaher Al Aghbari. Regionbased semantic image classification. International Journal of Image and Graphics (IJIG), 6(3):357–??, July 2006. CODEN ???? ISSN 0219-4678.

Aouat:2010:MDN

[AL10] Saliha Aouat and Slimane Larabi. Matching descriptors of noisy outline shapes. International Journal of Image and Graphics (IJIG), 10(3): 299–325, July 2010. CODEN ???? ISSN 0219-4678.

Atiampo:2016:UIS

[AL16] Armand Kodjo Atiampo and Georges Laussane Loum. Unsupervised image segmentation with pairwise Markov chains based on nonparametric estimation of copula using orthogonal polynomials. International Journal of Image and Graphics (IJIG), 16(4): 1650020, October 2016. CO-DEN ???? ISSN 0219-4678.

Ailisto:2003:RFI

[ALT03] Heikki Ailisto, Mikko Lindholm, and Pauli Tikkanen. A review of fingerprint image enhancement methods. International Journal of Image and Graphics (IJIG), 3(3): 401–??, July 2003. CODEN ???? ISSN 0219-4678.

Abdelwahab:2007:FCD

[AM07] Ahmed A. Abdelwahab and Nora S. Muharram. A fast codebook design algorithm based on a fuzzy clustering methodology. International Journal of Image and Graphics (IJIG), 7(2):291– 302, April 2007. CODEN ???? ISSN 0219-4678.

Aswatha:2016:IRS

[AMB16] Shashaank M. Aswatha, Jayanta Mukherjee, and Partha Bhowmick. An integrated repainting system for digital restoration of Vijayanagara murals. International Journal of Image and Graphics (IJIG), 16(1): 1650005, January 2016. CO-DEN ???? ISSN 0219-4678.

Ali:2010:VFS

[AMR^{+}10] Wajid Ali, Tangui Morvan, Petter Risholm, Ole Jakob Elle, and Eigil Samset. A visualization and fusion system for image guided RFA procedures. International Journal of Image and Graphics (IJIG), 10(2):155–174, April 2010. CODEN ???? ISSN 0219-4678.

Al-Naji:2017:CCA

[ANC17] Ali Al-Naji and Javaan Chahl. Contactless cardiac activity detection based on head motion magnification. International Journal of Image and Graphics (IJIG), 17 (1):1750001, January 2017. CODEN ???? ISSN 0219-4678.

Anonymous:2001:AI

[Ano01]  Anonymous.  Author in-

dex. International Journal of Image and Graphics (IJIG), 1(4):735–??, October 2001. CODEN ???? ISSN 0219-4678.

Anonymous:2002:AIV

[Ano02] Anonymous. Author index volume 2 (2002). International Journal of Image and Graphics (IJIG), 2(4):669–??, October 2002. CODEN ???? ISSN 0219-4678.

Anonymous:2003:AIV

[Ano03a] Anonymous. Author index volume 3 (2003). International Journal of Image and Graphics (IJIG), 3(4):671–??, October 2003. CODEN ???? ISSN 0219-4678.

Anonymous:2003:Ea

[Ano03b] Anonymous. Editorial. International Journal of Image and Graphics (IJIG), 3(1):1– ??, January 2003. CODEN ???? ISSN 0219-4678.

Anonymous:2003:Eb

[Ano03c] Anonymous. Editorial. International Journal of Image and Graphics (IJIG), 3(2): 231–??, April 2003. CODEN ???? ISSN 0219-4678.

Anonymous:2003:Ec

[Ano03d] Anonymous. Editorial. International Journal of Image and Graphics (IJIG), 3(3): 399–??, July 2003. CODEN ???? ISSN 0219-4678.

Anonymous:2003:Ed

[Ano03e] Anonymous. Editorial. International Journal of Image and Graphics (IJIG), 3(4): 531–??, October 2003. CO-DEN ???? ISSN 0219-4678.

Anonymous:2004:AIV

[Ano04a] Anonymous. Author index (volume 4). International Journal of Image and Graphics (IJIG), 4(4):721–??, October 2004. CODEN ???? ISSN 0219-4678.

Anonymous:2004:Ea

[Ano04b] Anonymous. Editorial. International Journal of Image and Graphics (IJIG), 4(3): 341–??, July 2004. CODEN ???? ISSN 0219-4678.

Anonymous:2004:Eb

[Ano04c] Anonymous. Editorial. International Journal of Image and Graphics (IJIG), 4(4): 533–??, October 2004. CO-DEN ???? ISSN 0219-4678.

Anonymous:2005:AIV

[Ano05a] Anonymous. Author index (volume 5). International Journal of Image and Graphics (IJIG), 5(4):871–??, October 2005. CODEN ???? ISSN 0219-4678.

Anonymous:2005:I

[Ano05b] Anonymous. Introduction. International Journal of Image and Graphics (IJIG), 5 (1):1–??, January 2005. CO-DEN ???? ISSN 0219-4678.

Anonymous:2006:AIV

[Ano06] Anonymous. Author index (volume 6). International Journal of Image and Graphics (IJIG), 6(4):689–691, October 2006. CODEN ???? ISSN 0219-4678.

Anonymous:2007:AIV

[Ano07a] Anonymous. Author index (volume 7). International Journal of Image and Graphics (IJIG), 7(4):795–797, October 2007. CODEN ???? ISSN 0219-4678.

Anonymous:2007:MLI

[Ano07b] Anonymous. Machine learning in image and graphics. International Journal of Image and Graphics (IJIG), 7 (3):v, July 2007. CODEN ???? ISSN 0219-4678.

Anonymous:2008:AIV

[Ano08a] Anonymous. Author index (volume 8). International Journal of Image and Graphics (IJIG), 8(4):663–665, October 2008. CODEN ???? ISSN 0219-4678.

Anonymous:2008:P

[Ano08b] Anonymous. Preface. International Journal of Image and Graphics (IJIG), 8(2):vii, April 2008. CODEN ???? ISSN 0219-4678.

Anonymous:2009:AIV

[Ano09] Anonymous. Author index (volume 9). International Journal of Image and Graphics (IJIG), 9(4):629–631, October 2009. CODEN ???? ISSN 0219-4678.

Anonymous:2010:AIV

[Ano10]  Anonymous.  Author in-

dex (volume 10). International Journal of Image and Graphics (IJIG), 10(4):609– 611, October 2010. CODEN ???? ISSN 0219-4678.

Anonymous:2011:AIV

[Ano11]  Anonymous.  Author in-

dex (volume 11). International Journal of Image and Graphics (IJIG), 11(4):629– 631, October 2011. CODEN ???? ISSN 0219-4678.

Anonymous:2012:AIV

[Ano12]  Anonymous.  Author In-

dex (volume 12). International Journal of Image and Graphics (IJIG), 12(4): 1299001, October 2012. CO-DEN ???? ISSN 0219-4678.

Anonymous:2013:AIV

[Ano13]  Anonymous.  Author in-

dex (volume 13). International Journal of Image and Graphics (IJIG), 13(4): 1399001, October 2013. CO-DEN ???? ISSN 0219-4678.

Anonymous:2014:AIV

[Ano14]  Anonymous.  Author in-

dex (volume 14). International Journal of Image and Graphics (IJIG), 14(4): 1499001, October 2014. CO-DEN ???? ISSN 0219-4678.

Anonymous:2015:AIV

[Ano15]  Anonymous.  Author in-

dex (volume 15). International Journal of Image and Graphics (IJIG), 15(4): 1599001, October 2015. CO-DEN ???? ISSN 0219-4678.

Anonymous:2016:AIV

[Ano16]  Anonymous.  Author in-

dex (volume 16). International Journal of Image and Graphics (IJIG), 16(4): 1699001, October 2016. CO-DEN ???? ISSN 0219-4678.

Anonymous:2017:AIV

[Ano17]  Anonymous.  Author in-

dex (volume 17). International Journal of Image and Graphics (IJIG), 17(4): 1799001, October 2017. CO-DEN ???? ISSN 0219-4678.

Anonymous:2018:AIV

[Ano18]  Anonymous.  Author in-

dex (volume 18). International Journal of Image and Graphics (IJIG), 18(4):??, October 2018. ISSN 0219-

                                        4678. URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0219467818990012.

Anonymous:2019:AIV

[Ano19]  Anonymous.  Author in-

dex (volume 19). International Journal of Image and Graphics (IJIG), 19(4):??, October 2019. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467819990018.

Anonymous:2020:AIV

[Ano20] Anonymous. Author index (volume 20). International Journal of Image and Graphics (IJIG), 20(04):??, October 2020. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467820990016.

Anonymous:2021:AIV

[Ano21] Anonymous. Author index (volume 21). International Journal of Image and Graphics (IJIG), 21(04):??, October 2021. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467821990011.

Adhikari:2023:PEC

[AP23] Ramesh Adhikari and Suresh Pokharel. Performance evaluation of convolutional neural network using synthetic medical data augmentation generated by GAN. International Journal of Image and Graphics (IJIG), 23(01):??, January 2023. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S021946782350002X.

Alcantara:2017:HAC

[APC17] Marlon F. Alcantara, Helio Pedrini, and Yu Cao. Human action classification based on silhouette indexed interest points for multiple domains. International Journal of Image and Graphics (IJIG), 17(3):1750018, July 2017. CODEN ???? ISSN 0219-4678.

Arora:2021:CHM

[Aro21]  Tanvi Arora.  Classifica-

tion of human metaspread images using convolutional neural networks. International Journal of Image and Graphics (IJIG), 21(03):??, July 2021. ISSN 0219-

                                        4678. URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0219467821500339.

Arora:2022:CBP

[Aro22] Tanvi Arora. CNN-based prediction of COVID-19 using chest CT images. International Journal of Image and Graphics (IJIG), 22(04): ??, July 2022. ISSN 0219-

                                        4678. URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0219467822500395.

Amin:2001:PSC

[AS01] A. Amin and R. Shiu. Page segmentation and classification utilizing bottom-up approach. International Journal of Image and Graphics (IJIG), 1(2):345–??, April 2001. CODEN ???? ISSN 0219-4678.

Agarwal:2008:DWS

[AS08] Rashmi Agarwal and M. S.

Santhanam. Digital watermarking in the singular vector domain. International Journal of Image and Graphics (IJIG), 8(3):351–368, July 2008. CODEN ???? ISSN 0219-4678.

Araujo:2020:ECQ

[ASJ20] Leonardo C. Araujo, Joao P. H. Sansao, and Mario C. S. Junior. Effects of color quantization on JPEG compression. International Journal of Image and Graphics (IJIG), 20(03): ??, July 2020. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467820500266.

Ayed:2007:TOB

[ASM07] Mohamed Ali Ben Ayed, Amine Samet, and Nouri Masmoudi. Toward an optimal block motion estimation algorithm for H.264/AVC. International Journal of Image and Graphics (IJIG), 7 (2):303–320, April 2007. CO-DEN ???? ISSN 0219-4678.

Arunlal:2020:DIC

[ASR20] S. L. Arunlal, N. Santhi, and K. Ramar. Design and implementation of content-based natural image retrieval approach using feature distance. International Journal of Image and Graphics (IJIG), 20(02): ??, April 2020. ISSN 0219-

                                        4678. URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S021946782050014X.

Anagnostopoulos:2013:EFS

[ASV13] Vasileios I. Anagnostopoulos, Emmanuel S. Sardis, and Theodora A. Varvarigou. Estimation of frame sequence noise with removal of JPEG artifacts. International Journal of Image and Graphics (IJIG), 13(1):1350003:1– 1350003:31, January 2013. CODEN ???? ISSN 0219-4678.

Ali:2014:LDF

[ATA14] Haider Ali, Umair Ullah Tariq, and Muhammad Abid. Learning discriminating features for gender recognition of real world faces. International Journal of Image and Graphics (IJIG), 14(3): 1450011, July 2014. CODEN ???? ISSN 0219-4678.

Agrawal:2013:P

[ATDT13] Anupam Agrawal, R. C. Tripathi, Ellen Yi-Luen Do, and M. D. Tiwar. Preface. International Journal of Image and Graphics (IJIG), 13(2): ??, April 2013. CODEN ???? ISSN 0219-4678.

Ablameyko:2007:RED

[AU07] Sergey V. Ablameyko and Seiichi Uchida. Recognition of engineering drawing entities: Review of approaches. International Journal of Image and Graphics (IJIG), 7 (4):709–733, October 2007. CODEN ???? ISSN 0219-4678.

Amin:2005:RST

[AW05] Adnan Amin and Sue Wu.

A robust system for thresholding and skew detection in mixed text/graphics documents. International Journal of Image and Graphics (IJIG), 5(2):247–??, April 2005. CODEN ???? ISSN 0219-4678.

Bibi:2020:SPP

[BAR20] Khalida Bibi, Ghazala Akram, and Kashif Rehan. Shape preserving properties with constraints on the tension parameter of binary threepoint approximating subdivision scheme. International Journal of Image and Graphics (IJIG), 20(01):??, January 2020. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467820500059.

Brahme:2020:EVV

[BB20] Aparna Brahme and Umesh Bhadade. Effect of various visual speech units on language identification using visual speech recognition. International Journal of Image and Graphics (IJIG), 20(04): ??, October 2020. ISSN 0219-

4678. URL https://www.

                                        worldscientific.com/doi/
                                        10.1142/S0219467820500291.

Brown:2021:CSF

[BB21] Kyle Brown and Nikolaos Bourbakis. Curve and surface fitting techniques in computer vision. International Journal of Image and Graphics (IJIG), 21(04): ??, October 2021. CO-DEN ???? ISSN 0219-

                                        4678. URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0219467821500418.

Boucetta:2022:BAU

[BB22] Aldjia Boucetta and Leila Boussaad. Biometric authentication using finger-vein patterns with deep-learning and discriminant correlation analysis. International Journal of Image and Graphics (IJIG), 22(01):??, January 2022. ISSN 0219-

                                        4678. URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0219467822500139.

Bhowmick:2005:DMS

[BBB^{+}05] Partha Bhowmick, Arijit Bishnu, Bhargab Bikram Bhattacharya, Malay Kumar Kundu, C. A. Murthy, and Tinku Acharya. Determination of minutiae scores for fingerprint image applications. International Journal of Image and Graphics (IJIG), 5(3):537–571, July 2005. CODEN ???? ISSN 0219-4678. URL http:/

                                        /ejournals.wspc.com.sg/

ijig/05/preserved-docs/
0503/S0219467805001896.
pdf.

Bhattacharjee:2015:CBH

[BBG15] Debjyoti Bhattacharjee, Ashish Bakshi, and Kuntal Ghosh. Comparison between an HVS inspired linear filter and the bilateral filter in performing “vision at a glance” through smoothing with edge preservation. International Journal of Image and Graphics (IJIG), 15(4):1550015, October 2015. CODEN ???? ISSN 0219-4678.

Bentahar:2021:HTB

[BBS^{+}21] Tarek Bentahar, Atef Bentahar, Riad Saidi, Hichem Mayache, and Karim Ferroudji. Hybrid technique of the branch-cut and the quality-guided for inSAR phase unwrapping. International Journal of Image and Graphics (IJIG), 21(03): ??, July 2021. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467821500406.

Bandyopadhyay:2016:ASB

[BCB16] Oishila Bandyopadhyay, Bhabatosh Chanda, and Bhargab B. Bhattacharya. Automatic segmentation of bones in X-ray images based on entropy measure. International Journal of Image and Graphics (IJIG), 16(1): 1650001, January 2016. CO-DEN ???? ISSN 0219-4678.

Banerjee:2023:SVD

[BDD23] Rudranath Banerjee, Sourav De, and Shouvik Dey. A survey on various deep learning algorithms for an efficient facial expression recognition system. International Journal of Image and Graphics (IJIG), 23(03):??, May 2023. ISSN 0219-

                                        4678. URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0219467822400058.

Bennamoun:2004:E

[Ben04] Mohammed Bennamoun. Editorial. International Journal of Image and Graphics (IJIG), 4(1):1–??, January 2004. CODEN ???? ISSN 0219-4678.

Boulila:2011:MAS

[BF11] Wadii Boulila and Imed Riadh Farah. Multi-approach satellite images fusion based on blind sources separation. International Journal of Image and Graphics (IJIG), 11 (1):117–136, January 2011. CODEN ???? ISSN 0219-4678.

Balster:2011:PCR

[BFT11] Eric J. Balster, Benjamin T. Fortener, and William F. Turri. Post-compression rate-distortion development for embedded block coding with optimal truncation in JPEG2000 imagery. International Journal of Image and Graphics (IJIG), 11(4):611– 627, October 2011. CODEN ???? ISSN 0219-4678.

Bouvier:2001:TTO

[BG01] E. Bouvier and E. Gobbetti.

TOM: Totally ordered mesh a multiresolution structure for time critical graphics applications. International Journal of Image and Graphics (IJIG), 1(1):115–??, January 2001. CODEN ???? ISSN 0219-4678.

Batagelo:2014:GBS

[BG14] Harlen Costa Batagelo and João Paulo Gois. GPU-based sphere tracing for radial basis function implicits. International Journal of Image and Graphics (IJIG), 14(1– 2):1450004, 2014. CODEN ???? ISSN 0219-4678.

Bi:2006:NNR

[BGX06] Dong-Liang Bi, Wei Guo, and Ai-Dong Xu. A new noise removing algorithm. International Journal of Image and Graphics (IJIG), 6(4): 677–687, October 2006. CO-DEN ???? ISSN 0219-4678.

Bao:2002:IRB

[BH02] Paul Bao and Sung-Wai Hong. Image restoration based on generalized finite automata encoded edge preserving regularization. International Journal of Image and Graphics (IJIG), 2(3): 425–??, July 2002. CODEN ???? ISSN 0219-4678.

Bourouis:2010:SMB

[BH10] Sami Bourouis and Kamel Hamrouni. 3D segmentation of MRI brain using level set and unsupervised classification. International Journal of Image and Graphics (IJIG), 10(1):135–154, January 2010. CODEN ???? ISSN 0219-4678.

Bania:2020:ATM

[BH20] Rubul Kumar Bania and Anindya Halder. Adaptive trimmed median filter for impulse noise detection and removal with an application to mammogram images. International Journal of Image and Graphics (IJIG), 20(04): ??, October 2020. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467820500321.

Bhat:2017:MIF

[Bha17] Aruna Bhat. Makeup invariant face recognition using features from accelerated segment test and eigen vectors. International Journal of Image and Graphics (IJIG), 17(1):1750005, January 2017. CODEN ???? ISSN 0219-4678.

Boudraa:2021:ECS

[BHM21] Omar Boudraa, Walid Khaled Hidouci, and Dominique Michelucci. An efficient cooperative smearing technique for degraded historical document image segmentation. International Journal of Image and Graphics (IJIG), 21(02): ??, April 2021. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467821500121.

BinMortuza:2019:KCB

[Bin19] Fahad Bin Mortuza. Kernelcoefficient-based feature method for face detection. International Journal of Image and Graphics (IJIG), 19 (2):??, 2019. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467819500098.

Bhanu:2004:MLA

[BJ04] Bir Bhanu and Grinnell Jones III. Multiple look angle SAR recognition. International Journal of Image and Graphics (IJIG), 4(1):85–??, January 2004. CODEN ???? ISSN 0219-4678.

Bhale:2013:ESF

[BJ13] Aparna Narendra Bhale and Manish Ratnakar Joshi. Enhancement of screen film mammogram up to a level of digital mammogram: Experimental analysis. International Journal of Image and Graphics (IJIG), 13(2): ??, April 2013. CODEN ???? ISSN 0219-4678.

Bini:2016:IRU

[BJ16] A. A. Bini and P. Jidesh. Image restoration using adaptive region-wise p-norm filter with local constraints. International Journal of Image and Graphics (IJIG), 16(2): 1650008, April 2016. CODEN ???? ISSN 0219-4678.

Bommisetty:2022:CBV

[BKKP22] Reddy Mounika Bommisetty, Ashish Khare, Manish Khare, and P. Palanisamy. Contentbased video retrieval using integration of curvelet transform and simple linear iterative clustering. International Journal of Image and Graphics (IJIG), 22(02): ??, April 2022. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467822500188.

Bonyadi:2011:NHQ

[BM11] Mohammad Reza Bonyadi and Mohsen Ebrahimi Moghaddam. A nonuniform highquality image compression method to preserve userspecified compression ratio. International Journal of Image and Graphics (IJIG), 11 (3):355–375, July 2011. CO-DEN ???? ISSN 0219-4678.

Bareja:2014:IIB

[BM14] Milan N. Bareja and Chintan K. Modi. An improved iterative back projection based single image super resolution approach. International Journal of Image and Graphics (IJIG), 14(4): 1450015, October 2014. CO-DEN ???? ISSN 0219-4678.

Bougleux:2007:SDS

[BME07] Sébastien Bougleux, Mahmoud Melkemi, and Abderrahim Elmoataz. Structure detection from a 3D set of points with anisotropic alphashapes. International Journal of Image and Graphics (IJIG), 7(4):689–708, October 2007. CODEN ???? ISSN 0219-4678.

Baciu:2008:GSW

[BMH08] George Baciu, Liang Ma, and Jinlian Hu. Generating seams and wrinkles for virtual clothing. International Journal of Image and Graphics (IJIG), 8 (2):281–297, April 2008. CO-DEN ???? ISSN 0219-4678.

Baig:2019:RTF

[BMK19] M. Amir Baig, Athar A.

Moinuddin, and Ekram Khan. Real-time fidelity measurement of JPEG2000 coded images. International Journal of Image and Graphics (IJIG), 19(1):??, January 2019. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467819500050.

Bao:2006:RTS

[BMS06] Paul Bao, Xiaohu Ma, and Wan-Chi Siu. Real-time seamless texture synthesis based on patch quantization clustering. International Journal of Image and Graphics (IJIG), 6(3):475–??, July 2006. CODEN ???? ISSN 0219-4678.

Benchaou:2018:FSB

[BNE18] Soukaina Benchaou, M’Barek Nasri, and Ouafae El Melhaoui. Feature selection based on evolution strategy for character recognition. International Journal of Image and Graphics (IJIG), 18(3): ??, July 2018. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467818500146.

Birajdar:2023:SSP

[BP23] Gajanan K. Birajdar and Mukesh D. Patil. A systematic survey on photorealistic computer graphic and photographic image discrimination. International Journal of Image and Graphics (IJIG), 23(04):??, July 2023. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467823500377.

Babu:2004:SGM

[BR04] R. Venkatesh Babu and K. R. Ramakrishnan. Sprite generation from MPEG video using motion information. International Journal of Image and Graphics (IJIG), 4(2):263–??, April 2004. CODEN ???? ISSN 0219-4678.

Bhatnagar:2009:RRW

[BR09] Gaurav Bhatnagar and Balasubramanian Raman. Robust reference-watermarking scheme using wavelet packet transform and bidiagonalsingular value decomposition. International Journal of Image and Graphics (IJIG), 9 (3):449–477, July 2009. CO-DEN ???? ISSN 0219-4678.

Basheera:2021:GMS

[BR21] Shaik Basheera and M. Satya Sai Ram. Gray matter segmentation of brain MRI using hybrid enhanced independent component analysis. International Journal of Image and Graphics (IJIG), 21(03): ??, July 2021. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467821500297.

Bhasha:2022:AIS

[BR22] A. Valli Bhasha and B. D. Venkatramana Reddy. Automated image super resolution with the aid of activation function optimized deep CNN and adaptive wavelet lifting approach. International Journal of Image and Graphics (IJIG), 22(05):??, October 2022. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467822500462.

Budhiraja:2021:IVI

[BRAS21] Sumit Budhiraja, Iftisam Rummy, Sunil Agrawal, and Balwinder Singh Sohi. Infrared and visible image fusion based on sparse representation and spatial frequency in DTCWT domain. International Journal of Image and Graphics (IJIG), 21(02): ??, April 2021. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467821500170.

Brown:2004:IGR

[BS04] Michael S. Brown and W. Brent Seales. Incorporating geometric registration with PC-cluster rendering for flexible tiled displays. International Journal of Image and Graphics (IJIG), 4(4): 683–??, October 2004. CO-DEN ???? ISSN 0219-4678.

Biswas:2019:CPM

[BS19] Biswajit Biswas and Biplab Kanti Sen. Color PET-MRI medical image fusion combining matching regional spectrum in shearlet domain. International Journal of Image and Graphics (IJIG), 19(1):??, January 2019. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467819500049.

Brunn:2007:CRU

[BSS07] Meru Brunn, Mario Costa Sousa, and Faramarz F. Samavati. Capturing and reusing artistic styles with reverse subdivision-based multiresolution methods. International Journal of Image and Graphics (IJIG), 7(4): 593–615, October 2007. CO-DEN ???? ISSN 0219-4678.

Brown:2005:DRC

[BT05] Michael S. Brown and Yau-Chat Tsoi. Distortion removal for camera-imaged print materials using boundary interpolation. International Journal of Image and Graphics (IJIG), 5(2):311–??, April 2005. CODEN ???? ISSN 0219-4678.

Bhardwaj:2022:IAU

[BVG22] Anuj Bhardwaj, Vivek Singh Verma, and Sandesh Gupta. Image authentication using block truncation coding in lifting wavelet domain. International Journal of Image and Graphics (IJIG), 22(01): ??, January 2022. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467822500115.

BenSalah:2022:FES

[BYN22] Marwa Ben Salah, Ameni Yengui, and Mahmoud Neji. Feature extraction and selection in archaeological images for automatic annotation. International Journal of Image and Graphics (IJIG), 22(01): ??, January 2022. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467822500061.

Bing:2005:EPR

[BYNZ05] Cheng Bing, Wang Ying, Zheng Nanning, and Bian Zhengzhong. An efficient 3D plenoptic representation for approximating a path of motion to a curved line. International Journal of Image and Graphics (IJIG), 5(2):397–??, April 2005. CODEN ???? ISSN 0219-4678.

Bai:2010:HPT

[BZ10] Xiaoliang Bai and Shusheng Zhang. Hierarchical parameterization of triangular mesh with a boundary polygon triangulation. International Journal of Image and Graphics (IJIG), 10(3):449– 466, July 2010. CODEN ???? ISSN 0219-4678.

Cai:2006:DIW

[CA06] Weiting Cai and Malek Adjouadi. Design and implementation of wavelet-domain video compression using multiresolution motion estimation and compensation. International Journal of Image and Graphics (IJIG), 6(4): 533–549, October 2006. CO-DEN ???? ISSN 0219-4678.

Cai:2016:PRL

[Cai16] Haipeng Cai. Parallel rendering for legible illustrative visualizations of dense geometries on commodity CPUs. International Journal of Image and Graphics (IJIG), 16 (1):1650002, January 2016. CODEN ???? ISSN 0219-4678.

Cui:2007:DMF

[CBCT07] Yunfeng Cui, Jing Bai, Yingmao Chen, and Jiahe Tian.

A digital model framework of metabolic system based on visible human data set. International Journal of Image and Graphics (IJIG), 7(1): 143–157, January 2007. CO-DEN ???? ISSN 0219-4678.

Chamundeshwari:2023:HPE

[CBU23] Chamundeshwari, Nagashetteppa Biradar, and Udaykumar. Hybrid pattern extraction with deep learning-based heart disease diagnosis using echocardiogram images. International Journal of Image and Graphics (IJIG), 23(02): ??, March 2023. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467823500249.

Chow:2006:FRR

[CC06] S. K. Chow and K. L. Chan.

Fast and realistic rendering of deformable virtual characters using impostor and stencil buffer. International Journal of Image and Graphics (IJIG), 6(4):599–624, October 2006. CODEN ???? ISSN 0219-4678.

Chia:2011:BUI

[CCAS11] Wai Chong Chia, Li Wern Chew, Li-Minn Ang, and Kah Phooi Seng. Binaryuncoded image and video compression using SPIHT– ZTR coding. International Journal of Image and Graphics (IJIG), 11(3):415–437, July 2011. CODEN ???? ISSN 0219-4678.

Cheng:2007:OBP

[CCL^{+}07] Jun Cheng, Ronald Chung, Edmund Y. Lam, Kenneth S. M. Fung, and Yangsheng Xu. Optimization of bit-pairing codification with learning for 3D reconstruction. International Journal of Image and Graphics (IJIG), 7 (3):445–462, July 2007. CO-DEN ???? ISSN 0219-4678.

Chaudhari:2022:IFU

[CD22] Chaitrali Prasanna Chaudhari and Satish Devane. Improved framework using rider optimization algorithm for precise image caption generation. International Journal of Image and Graphics (IJIG), 22(02):??, April 2022. ISSN 0219-

                                        4678. URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0219467822500218.

Cheriet:2005:SFB

[CDD05] Mohamed Cheriet, Jean-Christophe Demers, and Sylvain Deblois. Shock filterbased diffusion fields — application to grayscale character image processing. International Journal of Image and Graphics (IJIG), 5(2):209–??, April 2005. CODEN ???? ISSN 0219-4678.

Chen:2005:TSC

[CFC05] Zhe Chen, David Dagan Feng, and Weidong Cai. Temporal and spatial compression of dynamic positron emission tomography in sinogram domain. International Journal of Image and Graphics (IJIG), 5(4):839–??, October 2005. CODEN ???? ISSN 0219-4678.

Chalamalasetty:2021:RPT

[CG21] Sai Pratheek Chalamalasetty and Srinivasa Rao Giduturi. Research perception towards copy-move image forgery detection: Challenges and future directions. International Journal of Image and Graphics (IJIG), 21(04):??, October 2021. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467821500546.

Chole:2023:LMO

[CG23] Vikrant Chole and Vijay Gadicha. Locust mayfly optimization-tuned neural network for AI-based pruning in chess game. International Journal of Image and Graphics (IJIG), 23(04): ??, July 2023. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467823500286.

Cao:2006:DRD

[CGG^{+}06] Weiqun Cao, Hendrik Gaertner, Hannes Guddat, Andreas M. Straube, Stefan Conrad, Ernst Kruijff, and Dirk Langenberg. Design review in a distributed collaborative virtual environment. International Journal of Image and Graphics (IJIG), 6 (1):45–??, January 2006. CO-DEN ???? ISSN 0219-4678.

Cucchiara:2003:SVT

[CGP03] Rita Cucchiara, Costantino Grana, and Andrea Prati. Semantic video transcoding using classes of relevance. International Journal of Image and Graphics (IJIG), 3(1): 145–??, January 2003. CO-DEN ???? ISSN 0219-4678.

Courty:2008:ANR

[CH08] Nicolas Courty and Pierre Hellier. Accelerating 3D nonrigid registration using graphics hardware. International Journal of Image and Graphics (IJIG), 8(1):81–98, January 2008. CODEN ???? ISSN 0219-4678.

Chan:2008:VBG

[Cha08] K. L. Chan. Video-based gait analysis by silhouette Chamfer distance and Kalman filter. International Journal of Image and Graphics (IJIG), 8 (3):383–418, July 2008. CO-DEN ???? ISSN 0219-4678.

Chen:2009:IRB

[CHT09] Huawei Chen, Ichiro Hagiwara, and A. Kiet Tieu. Image reconstruction based on combination of wavelet decomposition, inpainting and texture synthesis. International Journal of Image and Graphics (IJIG), 9(1):51–65, January 2009. CODEN ???? ISSN 0219-4678.

Cai:2007:CVC

[CHY07] Wenli Cai, Gordon J. Harris, and Hiroyuki Yoshida. Computation of vesselness in CTA images for fast and interactive vessel segmentation. International Journal of Image and Graphics (IJIG), 7(1): 159–176, January 2007. CO-DEN ???? ISSN 0219-4678.

Chen:2003:FAF

[CHZ^{+}03] Longbin Chen, Baogang Hu, Lei Zhang, Mingjing Li, and Hongjiang Zhang. Face annotation for family photo album management. International Journal of Image and Graphics (IJIG), 3(1):81–??, January 2003. CODEN ???? ISSN 0219-4678.

Cai:2001:MPP

[CL01] J. Cai and Z.-Q. Liu. Markov process in pattern recognition. International Journal of Image and Graphics (IJIG), 1 (2):287–??, April 2001. CO-DEN ???? ISSN 0219-4678.

Canejo:2023:EDN

[CM23] Marcos José Canêjo and Carlos Alexandre Barros De Mello. Edge detection in natural scenes inspired by the speed drawing challenge. International Journal of Image and Graphics (IJIG), 23(01): ??, January 2023. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467823500092.

Curran:2007:APD

[CML07] Kevin Curran, Neil Mc-Caughley, and Xuelong Li. Addressing the problems of detecting faces with neural networks. International Journal of Image and Graphics (IJIG), 7(4):617–640, October 2007. CODEN ???? ISSN 0219-4678.

Coli:2008:FSR

[CMR08] Pietro Coli, Gian Luca Marcialis, and Fabio Roli. Fingerprint silicon replicas: Static and dynamic features for vitality detection using an optical capture device. International Journal of Image and Graphics (IJIG), 8(4): 495–512, October 2008. CO-DEN ???? ISSN 0219-4678.

Choudhary:2021:MBB

[CN21] Swati K. Choudhary and Ameya K. Naik. Multimodal biometric-based authentication with secured templates. International Journal of Image and Graphics (IJIG), 21 (02):??, April 2021. ISSN 0219-4678. URL https://

www.worldscientific.com/
doi/10.1142/S0219467821500182.

Carvalho:2015:IGB

[CNS^{+}15] L. E. Carvalho, S. L. Mantelli Neto, A. C. Sobieranski, E. Comunello, and A. von Wangenheim. Improving graph-based image segmentation using nonlinear color similarity metrics. International Journal of Image and Graphics (IJIG), 15(4): 1550018, October 2015. CO-DEN ???? ISSN 0219-4678.

Carvalho:2014:HCS

[CNvW^{+}14] L. E. Carvalho, S. L. Mantelli Neto, A. von Wangenheim, A. C. Sobieranski, L. Coser, and E. Comunello. Hybrid color segmentation method using a customized nonlinear similarity function. International Journal of Image and Graphics (IJIG), 14(1– 2):1450005, 2014. CODEN ???? ISSN 0219-4678.

Chang:2008:LFE

[CS08] Kuan-Tsung Chang and Tian-Yuan Shih. Linear features extraction with an orientation constrained probabilistic Hough transform. International Journal of Image and Graphics (IJIG), 8(1): 157–168, January 2008. CO-DEN ???? ISSN 0219-4678.

Chen:2018:VCE

[CS18] Liang-Hua Chen and Chih-Wen Su. Video caption extraction using spatiotemporal slices. International Journal of Image and Graphics (IJIG), 18(2):1850009, April 2018. CODEN ???? ISSN 0219-4678.

Choudhury:2018:SBC

[CTI^{+}18] Bismita Choudhury, Patrick Then, Biju Issac, Valliappan Raman, and Manas Kumar Haldar. A survey on biometrics and cancelable biometrics systems. International Journal of Image and Graphics (IJIG), 18(1):1850006, January 2018. CODEN ???? ISSN 0219-4678.

Caetano:2016:VDU

[CVdSdS16] Felipe Andrade Caetano, Marcelo Bernardes Vieira, and Rodrigo Luis de Souza da Silva. A video descriptor using orientation tensors and shape-based trajectory clustering. International Journal of Image and Graphics (IJIG), 16(4):1650017, October 2016. CODEN ???? ISSN 0219-4678.

Chi:2005:DIB

[CW05] Zheru Chi and Qing Wang. Document image binarization with feedback for improving character segmentation. International Journal of Image and Graphics (IJIG), 5(2): 281–??, April 2005. CODEN ???? ISSN 0219-4678.

Cai:2003:SPA

[CWFW03] Kangying Cai, Wencheng Wang, Guangzheng Fei, and Enhua Wu. A single-pass approach to adaptive simplification of out-of-core models. International Journal of Image and Graphics (IJIG), 3 (2):291–??, April 2003. CO-DEN ???? ISSN 0219-4678.

Cejnog:2017:WAR

[CYV17] Luciano W. X. Cejnog, Fernando A. A. Yamada, and Marcelo Bernardes Vieira. Wide angle rigid registration using a comparative tensor shape factor. International Journal of Image and Graphics (IJIG), 17(1):1750006, January 2017. CODEN ???? ISSN 0219-4678.

Chang:2009:CSD

[CYZ09] Jian Chang, Xiaosong Yang, and Jian J. Zhang. Continuous skeleton-driven skinning — a general approach for modeling skin deformation. International Journal of Image and Graphics (IJIG), 9(4):591–608, October 2009. CODEN ???? ISSN 0219-4678.

Chiang:2001:SVC

[CZ01] T. Chiang and Y.-Q. Zhang.

Stereoscopic video coding using a fast and robust affine motion search. International Journal of Image and Graphics (IJIG), 1(2):231–??, April 2001. CODEN ???? ISSN 0219-4678.

Cheng:2019:LDC

[CZS^{+}19] Lu Cheng, Yuan-Ke Zhang, Yun Song, Chen Li, and Dao-Shun Guo. Low-dose CT image restoration based on adaptive prior feature matching and nonlocal means. International Journal of Image and Graphics (IJIG), 19 (3):??, 2019. ISSN 0219-

                                        4678. URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0219467819500177.

Cheng:2017:HLM

[CZW^{+}17] Ruzhong Cheng, Yongjun Zhang, Guoping Wang, Yong Zhao, and Rahmatulloev Khusravsho. Haar-like multigranularity texture features for pedestrian detection. International Journal of Image and Graphics (IJIG), 17(4): 1750023, October 2017. CO-DEN ???? ISSN 0219-4678.

Dornaika:2004:FFF

[DA04] F. Dornaika and J. Ahlberg. Face and facial feature tracking using deformable models. International Journal of Image and Graphics (IJIG), 4 (3):499–??, July 2004. CO-DEN ???? ISSN 0219-4678.

Dharwadkar:2010:SSG

[DA10] Nagaraj V. Dharwadkar and B. B. Amberker. Steganographic scheme for gray-level image using pixel neighborhood and LSB substitution. International Journal of Image and Graphics (IJIG), 10 (4):589–607, October 2010. CODEN ???? ISSN 0219-4678.

Didier:2008:HCP

[DAM08] Jean-Yves Didier, Fakhr-Eddine Ababsa, and Malik Mallem. Hybrid camera pose estimation combining square fiducials localization technique and orthogonal iteration algorithm. International Journal of Image and Graphics (IJIG), 8(1): 169–188, January 2008. CO-DEN ???? ISSN 0219-4678.

Dimitrova:2001:VCU

[DAW01] N. Dimitrova, L. Agnihotri, and G. Wei. Video classification using object tracking. International Journal of Image and Graphics (IJIG), 1 (3):487–??, July 2001. CO-DEN ???? ISSN 0219-4678.

Drago:2005:LAC

[DC05] Frédéric Drago and Norishige Chiba. Locally adaptive chromatic restoration of digitally acquired paintings. International Journal of Image and Graphics (IJIG), 5(3):617– 637, July 2005. CODEN ???? ISSN 0219-4678. URL http:

//ejournals.wspc.com.sg/
ijig/05/preserved-docs/
0503/S0219467805001914.
pdf.

Carvalho:2013:IVF

[DDSP13] Paulo Roberto De Carvalho, Jr., Maikon Cismoski Dos Santos, William Robson Schwartz, and Helio Pedrini. An improved view frustum culling method using octrees for 3D real-time rendering. International Journal of Image and Graphics (IJIG), 13 (3):??, July 2013. CODEN ???? ISSN 0219-4678.

Doulamis:2003:ECB

[DDV03] Anastasios Doulamis, Nikolaos Doulamis, and Theodora Varvarigou. Efficient contentbased image retrieval using fuzzy organization and optimal relevance feedback. International Journal of Image and Graphics (IJIG), 3(1): 171–??, January 2003. CO-DEN ???? ISSN 0219-4678.

Drew:2011:ICR

[DF11] Mark S. Drew and Graham D. Finlayson. Improvement of colorization realism via the structure tensor. International Journal of Image and Graphics (IJIG), 11(4): 589–609, October 2011. CO-DEN ???? ISSN 0219-4678.

Dinesh:2009:NPA

[DG09] R. Dinesh and D. S. Guru. Non-parametric adaptive approach for the detection of dominant points on boundary curves based on nonsymmetric region of support. International Journal of Image and Graphics (IJIG), 9 (4):541–557, October 2009. CODEN ???? ISSN 0219-4678.

Dinesh:2010:CTS

[DG10] R. Dinesh and D. S. Guru. Concept of triangular spatial relationship and B-tree for partially occluded object recognition: an efficient and robust approach. International Journal of Image and Graphics (IJIG), 10(3):423– 448, July 2010. CODEN ???? ISSN 0219-4678.

Durette:2010:SRT

[DHA10] Barthélémy Durette, Jeanny Hérault, and David Alleysson. Simulation of the retina: a tool for visual prostheses. International Journal of Image and Graphics (IJIG), 10(4):513–529, October 2010. CODEN ???? ISSN 0219-4678.

Dawoud:2005:NSN

[DK05] Amer Dawoud and Mohamed Kamel. Natural skeletonization: New approach for the skeletonization of handwritten characters. International Journal of Image and Graphics (IJIG), 5(2):267–??, April 2005. CODEN ???? ISSN 0219-4678.

Dorini:2013:SST

[DL13a] Leyza Baldo Dorini and Neucimar Jerônimo Leite. A scale-space toggle operator for image transformations. International Journal of Image and Graphics (IJIG), 13 (4):1350022, October 2013. CODEN ???? ISSN 0219-4678.

Dorini:2013:SDF

[DL13b] Leyza Baldo Dorini and Neucimar Jerônimo Leite. A self-dual filtering toggle operator for speckle noise filtering. International Journal of Image and Graphics (IJIG), 13(3):??, July 2013. CODEN ???? ISSN 0219-4678.

Deng:2011:IQM

[DLZ^{+}11] Cheng Deng, Jie Li, Yifan Zhang, Dongyu Huang, and Lingling An. An image quality metric based on biologically inspired feature model. International Journal of Image and Graphics (IJIG), 11 (2):265–279, April 2011. CO-DEN ???? ISSN 0219-4678.

De:2013:SDO

[DM13] Kanjar De and V. Masilamani. A spatial domain object separability based noreference image quality measure using mean and variance. International Journal of Image and Graphics (IJIG), 13 (2):??, April 2013. CODEN ???? ISSN 0219-4678.

Djebali:2007:CBM

[DMMS07] M. Djebali, M. Melkemi, K. Melkemi, and N. Sapidis. Coiflet based methods for range image processing. International Journal of Image and Graphics (IJIG), 7(2): 321–351, April 2007. CODEN ???? ISSN 0219-4678.

Oliveira:2019:NTB

[dOGC^{+}19] Walter Alexandre A. de Oliveira, Denise Guliato, Douglas Coelho Braga de Oliveira, Rodrigo Luis de Souza da Silva, and Gilson Antonio Giraldi. New technique for binary morphological shape-based interpolation. International Journal of Image and Graphics (IJIG), 19(2):??, 2019. CO-DEN ???? ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467819500074.

Dhariwal:2020:INW

[DP20] Sumit Dhariwal and Sellappan Palaniappan. Image normalization and weighted classification using an efficient approach for SVM classifiers. International Journal of Image and Graphics (IJIG), 20(04):??, October 2020. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467820500357.

Dhar:2022:SRR

[DP22] Soumi Dhar and Shyamosree Pal. Surface reconstruction: Roles in the field of computer vision and computer graphics. International Journal of Image and Graphics (IJIG), 22(01):??, January 2022. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467822500085.

Devi:2023:DSF

[DP23] Bhagyashri Devi and M. Mary Synthuja Jain Preetha. A descriptive survey on face emotion recognition techniques. International Journal of Image and Graphics (IJIG), 23(01):??, January 2023. ISSN 0219-

                                        4678. URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0219467823500080.

DeVisser:2010:DNG

[DPC^{+}10] Hans De Visser, Josh Passenger, David Conlan, Christoph Russ, David Hellier, Mario Cheng, Oscar Acosta, Sébastien Ourselin, and Olivier Salvado. Developing a next generation colonoscopy simulator. International Journal of Image and Graphics (IJIG), 10(2):203–217, April 2010. CODEN ???? ISSN 0219-4678.

Djemal:2006:AAC

[DPR06] Khalifa Djemal, William Puech, and Bruno Rossetto. Automatic active contours propagation in a sequence of medical images. International Journal of Image and Graphics (IJIG), 6(2):267–??, April 2006. CODEN ???? ISSN 0219-4678.

Divakara:2018:HSA

[DPR18] S. S. Divakara, Sudarshan Patilkulkarni, and Cyril Prasanna Raj. High speed area optimized hybrid DA architecture for 2D-DTCWT. International Journal of Image and Graphics (IJIG), 18(1):1850004, January 2018. CODEN ???? ISSN 0219-4678.

Divakara:2020:NDI

[DPR20] S. S. Divakara, Sudarshan Patilkulkarni, and Cyril Prasanna Raj. Novel DWT/IDWT architecture for 3D with nine stage 2D parallel processing using split distributed arithmetic. International Journal of Image and Graphics (IJIG), 20 (03):??, July 2020. CO-DEN ???? ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467820500175.

Das:2023:NSE

[DPSA23] Gyanesh Das, Rutuparna Panda, Leena Samantaray, and Sanjay Agrawal. A novel segmentation error minimization-based method for multilevel optimal threshold selection using opposition equilibrium optimizer. International Journal of Image and Graphics (IJIG), 23(02): ??, March 2023. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467823500213.

Debnath:2020:UAS

[DR20] Saswati Debnath and Pinki Roy. User authentication system based on speech and cascade hybrid facial feature. International Journal of Image and Graphics (IJIG), 20(03): ??, July 2020. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467820500229.

Dube:2012:PQT

[DS12] Mridula Dube and Reenu Sharma. Piecewise quartic trigonometric polynomial Bspline curves with two shape parameters. International Journal of Image and Graphics (IJIG), 12(4):1250028, October 2012. CODEN ???? ISSN 0219-4678.

Dixit:2019:FBD

[DS19] Umesh D. Dixit and M. S. Shirdhonkar. Fingerprintbased document image retrieval. International Journal of Image and Graphics (IJIG), 19(2):??, 2019. ISSN 0219-4678. URL https://

                                        www.worldscientific.com/
                                        doi/10.1142/S0219467819500086.

Dehshibi:2015:GPW

[DSA15] Mohammad Mahdi Dehshibi, Ali Shirmohammadi, and Andrew Adamatzky. On growing Persian words with Lsystems: Visual modeling of Neyname. International Journal of Image and Graphics (IJIG), 15(3):1550011, July 2015. CODEN ???? ISSN 0219-4678.

Dahiya:2023:RDL

[DSG23] Neelam Dahiya, Sartajvir Singh, , and Sheifali Gupta. A review on deep learning classifier for hyperspectral imaging. International Journal of Image and Graphics (IJIG), 23(04):??, July 2023. ISSN 0219-

                                        4678. URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0219467823500365.

Domadiya:2018:SFE

[DSM18] Prashant Domadiya, Pratik Shah, and Suman K. Mitra. Shadow-free, expeditious and precise, moving object separation from video. International Journal of Image and Graphics (IJIG), 18(1): 1850005, January 2018. CO-DEN ???? ISSN 0219-4678.

daSilva:2015:EDB

[dSMP15] Ricardo Dutra da Silva, Rosane Minghim, and Helio Pedrini. 3D edge detection based on Boolean functions and local operators. International Journal of Image and Graphics (IJIG), 15(1): 1550003, January 2015. CO-DEN ???? ISSN 0219-4678.

daSilva:2016:GFA

[dSP16] Flávio Altinier Maximiano da Silva and Helio Pedrini. Geometrical features and active appearance model applied to facial expression recognition. International Journal of Image and Graphics (IJIG), 16(4):1650019, October 2016. CODEN ???? ISSN 0219-4678.

Dramas:2010:AVB

[DTJ10] Florian Dramas, Simon J.

Thorpe, and Christophe Jouffrais. Artificial vision for the blind: a bio-inspired algorithm for objects and obstacles detection. International Journal of Image and Graphics (IJIG), 10(4):531– 544, October 2010. CODEN ???? ISSN 0219-4678.

Dutta:2019:ISB

[Dut19] P. K. Dutta. Image segmentation based approach for the purpose of developing satellite image spatial information extraction for forestation and river bed analysis. International Journal of Image and Graphics (IJIG), 19(1): ??, January 2019. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467819500025.

MacIel:2020:SOF

[dV20] Luiz Maurílio da Silvad Maciel and Marcelo Bernardes Vieira. Sparse optical flow computation using wave equation-based energy. International Journal of Image and Graphics (IJIG), 20 (04):??, October 2020. CO-DEN ???? ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467820500278.

Deng:2005:PBS

[DWZ^{+}05] Yuhui Deng, Frank Wang, Jiangling Zhang, Dan Feng, Fang Wang, and Hong Jiang. Push the bottleneck of streaming media system from streaming media server to network. International Journal of Image and Graphics (IJIG), 5(4):859–??, October 2005. CODEN ???? ISSN 0219-4678.

ElRube:2006:MRS

[EAK^{+}06] Ibrahim El Rubé, Naif Alajlan, Mohamed S. Kamel, Maher Ahmed, and George H. Freeman. Mtar: a robust 2D shape representation. International Journal of Image and Graphics (IJIG), 6(3): 421–??, July 2006. CODEN ???? ISSN 0219-4678.

Engelbrecht:2015:DVI

[EBA15] Louis Engelbrecht, Adele Botha, and Ronell Alberts. Designing the visualization of information. International Journal of Image and Graphics (IJIG), 15(2):1540005, April 2015. CODEN ???? ISSN 0219-4678.

Echizen:2005:PAV

[EFY^{+}05] Isao Echizen, Yasuhiro Fujii, Takaaki Yamada, Satoru Tezuka, and Hiroshi Yoshiura. Perceptually adaptive video watermarking using motion estimation. International Journal of Image and Graphics (IJIG), 5(1):89–??, January 2005. CODEN ???? ISSN 0219-4678.

ElBadawy:2002:SBI

[EK02] Ossama El Badawy and Mohamed Kamel. Shape-based image retrieval applied to trademark images. International Journal of Image and Graphics (IJIG), 2(3):375–??, July 2002. CODEN ???? ISSN 0219-4678.

Erwin:2023:RBV

[EPSF23] Erwin, Hadrians Kesuma Putra, Bambang Suprihatin, and Fathoni. Retinal blood vessel extraction using a new enhancement technique of modified convolution filters and Sauvola thresholding. International Journal of Image and Graphics (IJIG), 23(01): ??, January 2023. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467823500067.

El-Sattar:2010:NPC

[ES10] Hussein Karam Hussein Abd El-Sattar. A new plot/ character-based interactive system for story-based virtual reality applications. International Journal of Image and Graphics (IJIG), 10(1):113– 133, January 2010. CODEN ???? ISSN 0219-4678.

Elmoufidi:2023:CMI

[ESJaO23] Abdelali Elmoufidi, Ayoub Skouta, Said Jai-andaloussi, and Ouail Ouchetto. CNN with multiple inputs for automatic glaucoma assessment using fundus images. International Journal of Image and Graphics (IJIG), 23(01):??, January 2023. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467823500122.

El-Sana:2003:VDR

[ESS03] Jihad El-Sana and Neta Sokolovsky. View-dependent rendering for large polygonal models over networks. International Journal of Image and Graphics (IJIG), 3(2): 265–??, April 2003. CODEN ???? ISSN 0219-4678.

Elyounsi:2022:FAO

[ETB22] Asma Elyounsi, Hatem Tlijani, and Mohamed Salim Bouhlel. Firefly algorithm optimized functional link artificial neural network for ISAradar image recognition. International Journal of Image and Graphics (IJIG), 22(05): ??, October 2022. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467822500449.

Ehsaeyan:2021:MIT

[EZ21] Ehsan Ehsaeyan and Alireza Zolghadrasli. A multilevel image thresholding method using the Darwinian cuckoo search algorithm. International Journal of Image and Graphics (IJIG), 21(04):??, October 2021. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467821500522.

Ehsaeyan:2022:SDC

[EZ22] Ehsan Ehsaeyan and Alireza Zolghadrasli. A study on Darwinian crow search algorithm for multilevel thresholding. International Journal of Image and Graphics (IJIG), 22(01):??, January 2022. ISSN 0219-

4678. URL https://www.

                                        worldscientific.com/doi/
                                        10.1142/S0219467822500127.

Fang:2006:MSI

[FA06] Shiaofen Fang and Marwan Adada. Multi-scale isosurface extraction for volume visualization. International Journal of Image and Graphics (IJIG), 6(2):173–??, April 2006. CODEN ???? ISSN 0219-4678.

Farshi:2017:INR

[Far17] Taymaz Rahkar Farshi. Image noise reduction method based on compatibility with adjacent pixels. International Journal of Image and Graphics (IJIG), 17(3):1750014, July 2017. CODEN ???? ISSN 0219-4678.

Fischer:2002:AID

[FB02] S. Fischer and H. Bunke. Automatic identification of diatoms using visual humaninterpretable features. International Journal of Image and Graphics (IJIG), 2(1): 67–??, January 2002. CO-DEN ???? ISSN 0219-4678.

Favier:2006:DAI

[FD06] Pierre-Alexandre Favier and Pierre De Loor. From decision to action: Intentionality, a guide for the specification of intelligent agent’s behavior. International Journal of Image and Graphics (IJIG), 6 (1):87–??, January 2006. CO-DEN ???? ISSN 0219-4678.

Fabrizio:2008:ASP

[FD08] Jonathan Fabrizio and Jean Devars. An analytical solution to the perspective-n-point problem for common planar camera and for catadioptric sensor. International Journal of Image and Graphics (IJIG), 8(1):135–155, January 2008. CODEN ???? ISSN 0219-4678.

Feng:2001:CBR

[Fen01] D. Feng. Content-based retrieval of multimedia information. International Journal of Image and Graphics (IJIG), 1(1):83–??, January 2001. CODEN ???? ISSN 0219-4678.

Faudot:2006:SVV

[FG06] Dominique Faudot and Gilles Gesquiere. Study of volume variation of implicit objects. International Journal of Image and Graphics (IJIG), 6 (4):551–568, October 2006. CODEN ???? ISSN 0219-4678.

Filisbino:2015:CRM

[FGT15] Tiene A. Filisbino, Gilson A.

Giraldi, and Carlos E. Thomaz. Comparing ranking methods for tensor components in multilinear and concurrent subspace analysis with applications in face images. International Journal of Image and Graphics (IJIG), 15(1):1550006, January 2015.

CODEN ???? ISSN 0219-4678.

Fathima:2023:MIC

[FHR23] M. Dhilsath Fathima, R. Hariharan, and S. P. Raja. Multiple imputation by chained equations — K-nearest neighbors and deep neural network architecture for kidney disease prediction. International Journal of Image and Graphics (IJIG), 23(02):??, March 2023. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467823500146.

Fan:2012:GIC

[FJ12] N. Fan and Cheng Jin. Geometric invariants construction for semantic scene understanding from multiple views inspired by the human visual system. International Journal of Image and Graphics (IJIG), 12(2):1250012, April 2012. CODEN ???? ISSN 0219-4678.

Florinabel:2011:MBM

[FJS11] D. Jemi Florinabel, S. Ebenezer Juliet, and V. Sadasivam. Multiorientation-based multistructure morphological inpainting. International Journal of Image and Graphics (IJIG), 11(2):177–193, April 2011. CODEN ???? ISSN 0219-4678.

Fan:2007:FAR

[FLH^{+}07] Yi Fan, Hongbing Lu, Chongyang Hao, Zhengrong Liang, and Zhiming Zhou. Fast analytical reconstruction of gated cardiac SPECT with non-uniform attenuation compensation. International Journal of Image and Graphics (IJIG), 7(1):87–104, January 2007. CODEN ???? ISSN 0219-4678.

Fu:2012:DRA

[FLW^{+}12] Bin Fu, Wenxin Li, Minghui Wu, Rongfeng Li, and Zhuoqun Xu. A document rectification approach dealing with both perspective distortion and warping based on text flow curve fitting. International Journal of Image and Graphics (IJIG), 12(1): 1250002, January 2012. CO-DEN ???? ISSN 0219-4678.

Feddaoui:2012:IRM

[FMH12] Nadia Feddaoui, Hela Mahersia, and Kamel Hamrouni. Iris recognition method based on Gabor filters and uniform local binary patterns. International Journal of Image and Graphics (IJIG), 12(2): 1250014, April 2012. CODEN ???? ISSN 0219-4678.

Fang:2006:NMF

[FPHL06] Xianyong Fang, Zhigeng Pan, Gaoqi He, and Li Li. A new method of feature based image mosaic. International Journal of Image and Graphics (IJIG), 6(3):497–??, July 2006. CODEN ???? ISSN 0219-4678.

Fathimal:2016:SSS

[FR16] P. Mohamed Fathimal and P. Arockia Jansi Rani. K out of N secret sharing scheme for multiple color images with steganography and authentication. International Journal of Image and Graphics (IJIG), 16(2):1650010, April 2016. CODEN ???? ISSN 0219-4678.

Fuchsberger:2020:SAT

[FRC20] Alexander Fuchsberger, Brian Ricks, and Zhicheng Chen. A semi-automated technique for transcribing accurate crowd motions. International Journal of Image and Graphics (IJIG), 20(02):??, April 2020. ISSN 0219-

                                        4678. URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0219467820500126.

Fei:2017:EME

[FTWR17] Lunke Fei, Shaohua Teng, Jigang Wu, and Imad Rida. Enhanced minutiae extraction for high-resolution palmprint recognition. International Journal of Image and Graphics (IJIG), 17(4): 1750020, October 2017. CO-DEN ???? ISSN 0219-4678.

Feng:2008:TFM

[FZYH08] Guiyu Feng, David Zhang, Jian Yang, and Dewen Hu. A theoretical framework for matrix-based feature extraction algorithms with its application to image recognition.

International Journal of Image and Graphics (IJIG), 8 (1):1–23, January 2008. CO-DEN ???? ISSN 0219-4678.

Gavrilova:2001:TAC

[GA01] M. L. Gavrilova and M. H.

Alsuwaiyel. Two algorithms for computing the Euclidean distance transform. International Journal of Image and Graphics (IJIG), 1(4):635–??, October 2001. CODEN ???? ISSN 0219-4678.

Gang:2005:CNI

[GA05] Litao Gang and Ali N.

Akansu. Cover noise interference suppression in multimedia data hiding. International Journal of Image and Graphics (IJIG), 5(1):191–??, January 2005. CODEN ???? ISSN 0219-4678.

Gajabe:2022:SKB

[GA22] Rajashree Gajabe and Syed Taqi Ali. Secret key-based image steganography in spatial domain. International Journal of Image and Graphics (IJIG), 22(02):??, April 2022. CODEN ???? ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467822500140.

Gayatri:2023:CID

[GA23] Erapaneni Gayatri and S. L.

Aarthy. Challenges and imperatives of deep learning approaches for detection of melanoma: a review. International Journal of Image and Graphics (IJIG), 23(03): ??, May 2023. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467822400125.

Gupta:2013:QMC

[GBG13] Rajani Gupta, Prashant Bansod, and R. S. Gamad. Quality measure of the compressed echo, X-ray and CT images. International Journal of Image and Graphics (IJIG), 13(1):1350006:1– 1350006:29, January 2013. CODEN ???? ISSN 0219-4678.

Gavade:2023:HFD

[GBP23] Priyanka A. Gavade, Vandana S. Bhat, , and Jagadeesh Pujari. Hybrid features and deep learning model for facial expression recognition from videos. International Journal of Image and Graphics (IJIG), 23(05):??, September 2023. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467823500456.

Goh:2009:TSI

[GCB^{+}09] Hock-Ann Goh, Chee-Way Chong, Rosli Besar, Fazly Salleh Abas, and Kok-Swee Sim. Translation and scale invariants of Hahn moments. International Journal of Image and Graphics (IJIG), 9(2):271–285, April 2009. CODEN ???? ISSN 0219-4678.

Gantenapalli:2023:SMF

[GCM23] Srinivasa Rao Gantenapalli, Praveen Babu Choppala, , and James Stephen Meka. Selective mean filtering for reducing impulse noise in digital color images. International Journal of Image and Graphics (IJIG), 23(05):??, September 2023. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467823500493.

Guan:2006:VEC

[GCO^{+}06] Y. Q. Guan, Y. Y. Cai, M. Opas, Z. W. Xiong, and Y. T. Lee. A VR enhanced collaborative system for 3D confocal microscopic image processing and visualization. International Journal of Image and Graphics (IJIG), 6 (2):231–??, April 2006. CO-DEN ???? ISSN 0219-4678.

Guo:2014:FLD

[GFZ14] Yanyan Guo, Xiangdong Fei, and Qijun Zhao. Fingerprint liveness detection using multiple static features and random forests. International Journal of Image and Graphics (IJIG), 14(4): 1450021, October 2014. CO-DEN ???? ISSN 0219-4678.

Ghosh:2018:VVE

[GGC18] Swarup Kr Ghosh, Anupam Ghosh, and Amlan Chakrabarti. VEA: Vessel extraction algorithm by active contour model and a novel wavelet analyzer for diabetic retinopathy detection. International Journal of Image and Graphics (IJIG), 18(2): 1850008, April 2018. CODEN ???? ISSN 0219-4678.

Gill:2019:RES

[GGS19] Jasmeen Gill, Akshay Girdhar, and Tejwant Singh. A review of enhancement and segmentation techniques for digital images. International Journal of Image and Graphics (IJIG), 19(3):??, 2019. CODEN ???? ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S021946781950013X.

Gupta:2021:DCB

[GGS21] Isha Gupta, Sheifali Gupta, and Swati Singh. Different CNN-based architectures for detection of invasive ductal carcinoma in breast using histopathology images. International Journal of Image and Graphics (IJIG), 21(05):??, December 2021. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467821400039.

Gagaoua:2022:HMB

[GGTC22a] Meriem Gagaoua, Hamza Ghilas, Abdelkamel Tari, and Mohamed Cheriet. Histogram of marked background (HMB) feature extraction method for Arabic handwriting recognition. International Journal of Image and Graphics (IJIG), 22(02): ??, April 2022. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467822500152.

Ghilas:2022:SDI

[GGTC22b] Hamza Ghilas, Meriem Gagaoua, Abdelkamel Tari, and Mohamed Cheriet. Spatial Distribution of Ink at Keypoints (SDIK): a novel feature for word spotting in Arabic documents. International Journal of Image and Graphics (IJIG), 22(04): ??, July 2022. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467822500358.

Girishwaingankar:2020:PNB

[GJ20] Poorva Girishwaingankar and Sangeeta Milind Joshi. The PHY-NGSC-based ORT run length encoding scheme for video compression. International Journal of Image and Graphics (IJIG), 20 (02):??, April 2020. CO-DEN ???? ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467820500072.

Gore:2022:IBR

[GJ22] Sonal Gore and Jayant Jagtap. IDH-based radiogenomic characterization of glioma using local ternary pattern descriptor integrated with radiographic features and random forest classifier. International Journal of Image and Graphics (IJIG), 22(03): ??, May 2022. ISSN 0219-

                                        4678. URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0219467821400131.

Gudise:2021:MBI

[GKS21] Sandhya Gudise, Giri Babu Kande, and T. Satya Savithri. MR brain image segmentation to detect white matter, gray matter, and cerebro spinal fluid using TLBO algorithm. International Journal of Image and Graphics (IJIG), 21(03):??, July 2021. ISSN 0219-4678. URL https:

                                        //www.worldscientific.com/
                                        doi/10.1142/S0219467821500340.

Gaddour:2023:NMA

[GKV23] Houda Gaddour, Slim Kanoun, and Nicole Vincent. A new method for Arabic text detection in natural scene images. International Journal of Image and Graphics (IJIG), 23(01):??, January 2023. ISSN 0219-

                                        4678. URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0219467823500109.

Ghislain:2017:ATD

[GLN17] Pandry Koffi Ghislain, Georges Lausanne Loum, and Ouattara Nouho. Adaptation of telegraph diffusion equation for noise reduction on images. International Journal of Image and Graphics (IJIG), 17(2): 1750010, April 2017. CODEN ???? ISSN 0219-4678.

Gao:2009:FRS

[GLX09] Xinbo Gao, Jinxiu Li, and Bing Xiao. A face recognition scheme based on embedded hidden Markov model and selective ensemble strategy. International Journal of Image and Graphics (IJIG), 9 (3):355–367, July 2009. CO-DEN ???? ISSN 0219-4678.

Gambhir:2016:NFR

[GM16] Deepak Gambhir and Meenu Manchanda. A novel fusion rule for medical image fusion in complex wavelet transform domain. International Journal of Image and Graphics (IJIG), 16(4):1650022, October 2016. CODEN ???? ISSN 0219-4678.

Giangreco-Maidana:2018:CPS

[GMLASVV18] Alejandro J. Giangreco-Maidana, Horacio Legal-Ayala, Christian E. Schaerer, and Waldemar Villamayor-Venialbo. Contour-point signature shape descriptor for point correspondence. International Journal of Image and Graphics (IJIG), 18(2): 1850007, April 2018. CODEN ???? ISSN 0219-4678.

Gowda:2010:ANI

[GMYL10] Rahul Gowda, Shalin M.

Mehta, Yue Yang, and Baoxin Li. Adaptive nonlinear image enhancement of Gaussian degraded images. International Journal of Image and Graphics (IJIG), 10(3):365– 393, July 2010. CODEN ???? ISSN 0219-4678.

Giraldi:2004:ISI

[GO04] Gilson A. Giraldi and Antonio A. F. Oliveira. Invariant snakes and initialization of deformable models. International Journal of Image and Graphics (IJIG), 4(3): 363–??, July 2004. CODEN ???? ISSN 0219-4678.

Grim:2016:ART

[GOO^{+}16] Anna Grim, Timothy O’Connor, Peter J. Olver, Chehrzad Shakiban, Ryan Slechta, and Robert Thompson. Automatic reassembly of threedimensional jigsaw puzzles. International Journal of Image and Graphics (IJIG), 16 (2):1650009, April 2016. CO-DEN ???? ISSN 0219-4678.

Gupta:2016:CAA

[GP16] Pooja Gupta and Kuldip Pahwa. Clock algorithm analysis for increasing quality of digital images. International Journal of Image and Graphics (IJIG), 16(3): 1650016, July 2016. CODEN ???? ISSN 0219-4678.

Gatos:2004:FIM

[GPPA04] Basilios Gatos, Stavros J. Perantonis, Nikos Papamarkos, and Ioannis Andreadis. Fast implementation of morphological operations using binary image block decomposition. International Journal of Image and Graphics (IJIG), 4(2):183–??, April 2004. CODEN ???? ISSN 0219-4678.

Gavrilova:2001:CLI

[GR01] M. Gavrilova and J. Rokne.

Computing line intersections. International Journal of Image and Graphics (IJIG), 1 (2):217–??, April 2001. CO-DEN ???? ISSN 0219-4678.

Guillen-Reyes:2019:BLD

[GRDM19] Fernando O. Guillén-Reyes and Francisco J. Domínguez-Mota. Boundary layer detection techniques applied to edge detection. International Journal of Image and Graphics (IJIG), 19 (2):??, 2019. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467819500104.

Gunaratne:2003:EAF

[GS03] Pujitha Gunaratne and Yukio Sato. Estimation of asymmetry in facial actions for the analysis of motion dysfunction due to paralysis. International Journal of Image and Graphics (IJIG), 3(4):639–??, October 2003. CODEN ???? ISSN 0219-4678.

Gogineni:2022:TSP

[GS22] Rajesh Gogineni and Dhara J.

Sangani. A two-stage PANsharpening algorithm based on sparse representation for spectral distortion reduction. International Journal of Image and Graphics (IJIG), 22(01):??, January 2022. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467822500073.

Gill:2023:NIG

[GS23] Jasmeen Gill and Ravinder Pal Singh. Noninvasive grading and sorting of mango (Mangiferad indica L.) using antlion optimizerbased artificial neural networks. International Journal of Image and Graphics (IJIG), 23(05):??, September 2023. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467823500407.

Geng:2003:PUI

[GSF^{+}03] Weidong Geng, Wolfgang Strauss, Monika Fleischmann, Vladimir Elistratov, and Marina Kolesnik. Perceptual user interface in virtual shopping environment. International Journal of Image and Graphics (IJIG), 3(2):365–??, April 2003. CODEN ???? ISSN 0219-4678.

Ge:2008:PBA

[GSP^{+}08] Jinghua Ge, Daniel J. Sandin, Tom Peterka, Robert Kooima, Javier I. Girado, and Andrew Johnson. A pointbased asynchronous remote visualization framework for real-time virtual reality. International Journal of Image and Graphics (IJIG), 8(2):

189–207, April 2008. CODEN ???? ISSN 0219-4678.

Gupta:2022:SSC

[GSSM22] Anuj Kumar Gupta, Manvinder Sharma, Ankit Sharma, and Vikas Menon. A study on SARS-CoV-2 (COVID-19) and machine learning based approach to detect COVID-19 through X-ray images. International Journal of Image and Graphics (IJIG), 22(03): ??, May 2022. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467821400106.

Gao:2007:FVO

[GWZ07] Xin Gao, Yuanmei Wang, and Cishen Zhang. Fuzzy vector objective optimization algorithm for image reconstruction from incomplete projections. International Journal of Image and Graphics (IJIG), 7(1):55–69, January 2007. CODEN ???? ISSN 0219-4678.

Gong:2006:RTI

[GY06] Minglun Gong and Yee-Hong Yang. Rayset: a taxonomy for image-based rendering. International Journal of Image and Graphics (IJIG), 6(3):313–??, July 2006. CO-DEN ???? ISSN 0219-4678.

Hussain:2008:WBE

[HAO08] Muhammad Hussain, Turghunjan Abdukirim, and Yoshihiro Okada. Wavelet-based edge detection in digital images. International Journal of Image and Graphics (IJIG), 8(4):513–533, October 2008. CODEN ???? ISSN 0219-4678.

Hassan:2021:CEM

[Has21] Mohd Fikree Hassan. Color enhancement method to improve the colors of the images perceived by the elderly people. International Journal of Image and Graphics (IJIG), 21(01):??, January 2021. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467821500042.

Hawkins:2007:ESU

[Haw07] William G. Hawkins. On the equivalence of stable and unstable forms of the inverse circular harmonic transform solution for the Radon transform. International Journal of Image and Graphics (IJIG), 7(1):17–33, January 2007. CODEN ???? ISSN 0219-4678.

Heidary:2010:SSD

[HC10] Kaveh Heidary and H. John Caulfield. Spectral sensitivity design for optical sensors. International Journal of Image and Graphics (IJIG), 10 (1):21–39, January 2010. CO-DEN ???? ISSN 0219-4678.

Hamouda:2019:FAS

[HEB19] Maissa Hamouda, Karim Saheb Ettabaa, and Med Salim Bouhlel. Framework for automatic selection of kernels based on convolutional neural networks and CkMeans clustering algorithm. International Journal of Image and Graphics (IJIG), 19 (4):??, October 2019. CO-DEN ???? ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467819500190.

Helmy:2012:CMC

[Hel12] Tarek Helmy. A computational model for contextbased image categorization and description. International Journal of Image and Graphics (IJIG), 12(1): 1250001, January 2012. CO-DEN ???? ISSN 0219-4678.

Helmy:2018:GFS

[Hel18] Tarek Helmy. A generic framework for semantic annotation of images. International Journal of Image and Graphics (IJIG), 18(3):??, July 2018. CO-DEN ???? ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467818500134.

Hashimoto:2015:VCS

[HFI15] Hideyuki Hashimoto, Yuki Fujibayashi, and Hiroki Imamura. 3D video communication system by using Kinect and head mounted displays. International Journal of Image and Graphics (IJIG), 15 (2):1540003, April 2015. CO-DEN ???? ISSN 0219-4678.

Halder:2019:SSP

[HHCS19] Amiya Halder, Sayan Halder, Samrat Chakraborty, and Apurba Sarkar. A statistical salt-and-pepper noise removal algorithm. International Journal of Image and Graphics (IJIG), 19(1):??, January 2019. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467819500062.

Hassan:2022:NSI

[HHFA22] Gaber Hassan, Khalid M. Hosny, R. M. Farouk, and Ahmed M. Alzohairy. New set of invariant quaternion Krawtchouk moments for color image representation and recognition. International Journal of Image and Graphics (IJIG), 22(04):??, July 2022. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467822500371.

Ho:2008:RSI

[HHL^{+}08] Charlotte Yuk-Fan Ho, Tai-Chiu Hsung, Daniel Pak-Kong Lun, Bingo Wing-Kuen Ling, Peter Kwong-Shun Tam, and Wan-Chi Siu. Regularity scalable image coding based on wavelet singularity detection. International Journal of Image and Graphics (IJIG), 8(1): 109–134, January 2008. CO-DEN ???? ISSN 0219-4678.

Haddad:2013:SAC

[HJ13] Bashar Haddad and Amin Jarrah. Semi-automatic cracks correction based on seam processing, stochastic analysis and learning process. International Journal of Image and Graphics (IJIG), 13 (4):1350020, October 2013. CODEN ???? ISSN 0219-4678.

Hans:2022:HBB

[HK22] Rahul Hans and Harjot Kaur.

Hybrid biogeography-based optimization and genetic algorithm for feature selection in mammographic breast density classification. International Journal of Image and Graphics (IJIG), 22(03): ??, May 2022. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467821400076.

Han:2019:RBV

[HKP^{+}19] Myounghee Han, Yongjoo Kim, Jang Ryul Park, Benjamin J. Vakoc, Wang-Yuhl Oh, and Sukyoung Ryu. Retinal blood vessel caliber estimation for optical coherence tomography angiography images based on 3D superellipsoid modeling. International Journal of Image and Graphics (IJIG), 19 (2):??, 2019. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467819500116.

Huang:2009:FPH

[HL09] Bo Huang and Naimin Li. Fungiform papillae hyperplasia (FPH) identification by tongue texture analysis. International Journal of Image and Graphics (IJIG), 9(3): 479–494, July 2009. CODEN ???? ISSN 0219-4678.

Huang:2013:ADC

[HL13] Wei Huang and Hongtao Lu. Automatic defect classification of TFT-LCD panels with shape, histogram and color features. International Journal of Image and Graphics (IJIG), 13(3):??, July 2013. CODEN ???? ISSN 0219-4678.

Hanjalic:2001:RAV

[HLB01] A. Hanjalic, R. L. Lagendijk, and J. Biemond. Recent advances in video content analysis: From visual features to semantic video segments. International Journal of Image and Graphics (IJIG), 1(1): 63–??, January 2001. CO-DEN ???? ISSN 0219-4678.

Honnutagi:2023:UVE

[HLM23] Pooja Honnutagi, Y. S. Laitha, , and V. D. Mytri. Underwater video enhancement using manta ray foraging lion optimizationbased fusion convolutional neural network. International Journal of Image and Graphics (IJIG), 23(04):??, July 2023. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467823500316.

Hicks:2005:AMC

[HMM05] B. J. Hicks, G. Mullineux, and A. J. Medland. Automatic model creation for kinematic analysis and optimization of engineering systems. International Journal of Image and Graphics (IJIG), 5(3):481–499, July 2005. CODEN ???? ISSN 0219-4678. URL http:/

/ejournals.wspc.com.sg/
ijig/05/preserved-docs/
0503/S0219467805001860.
pdf.

Hild:2002:RSS

[HN02] M. Hild and K. Nishijima.

Reconstruction of 3D space structure with a rotational imaging system. International Journal of Image and Graphics (IJIG), 2(2):269–??, April 2002. CODEN ???? ISSN 0219-4678.

Hussain:2003:FSF

[HON03] Muhammad Hussain, Yoshihiro Okada, and Koichi Niijima. Fast, simple, feature preserving and memory efficient simplification of triangle meshes. International Journal of Image and Graphics (IJIG), 3(4):653–??, October 2003. CODEN ???? ISSN 0219-4678.

Hosny:2007:ECL

[Hos07] Khalid M. Hosny. Efficient computation of Legendre moments for gray level images. International Journal of Image and Graphics (IJIG), 7 (4):735–747, October 2007. CODEN ???? ISSN 0219-4678.

Hossain:2010:VIC

[HRH^{+}10] Md. Shafaeat Hossain, Khandaker Abir Rahman, Md. Hasanuzzaman, M. A. Bhuyian, and H. Ueno. Video image clustering based on human face and shirt color. International Journal of Image and Graphics (IJIG), 10(1):1–19, January 2010. CODEN ???? ISSN 0219-4678.

Hardy:2006:HII

[HS06] Alexandre Hardy and Willi-Hans Steeb. Harmonic interpolation for image compression. International Journal of Image and Graphics (IJIG), 6(4):657–675, October 2006. CODEN ???? ISSN 0219-4678.

Heng:2001:INV

[HSCW01] P.-A. Heng, H. Sun, K.-W. Chen, and T.-T. Wong. Interactive navigation of virtual vessel tracking with 3D intelligent scissors. International Journal of Image and Graphics (IJIG), 1(2):273–??, April 2001. CODEN ???? ISSN 0219-4678.

Haraguchi:2003:TDR

[HSEI03] Ryo Haraguchi, Naozo Sugimoto, Shigeru Eiho, and Yoshio Ishida. Three dimensional reconstruction of coronary arteries by using registration and texture-mapping onto epicardial surface on nuclear 3D image. International Journal of Image and Graphics (IJIG), 3(4):567–??, October 2003. CODEN ???? ISSN 0219-4678.

Hamroun:2022:MVI

[HTC22] Mohamed Hamroun, Karim Tamine, and Benoît Crespin. Multimodal video indexing (MVI): a new method based on machine learning and semi-automatic annotation on large video collections. International Journal of Image and Graphics (IJIG), 22(02): ??, April 2022. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S021946782250022X.

Hamroun:2021:NCB

[HTCZ21] Mohamed Hamroun, Karim Tamine, Frederic Claux, and Mourad Zribi. A new contentbased image retrieval system using deep visual features. International Journal of Image and Graphics (IJIG), 21(04): ??, October 2021. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467821500558.

Hu:2004:SCR

[HU04] Zhencheng Hu and Keiichi Uchimura. Solution of camera registration problem via 3D–2D parameterized model matching for on-road navigation. International Journal of Image and Graphics (IJIG), 4 (1):3–??, January 2004. CO-DEN ???? ISSN 0219-4678.

Hu:2005:LBR

[Hu05] Yu-Chen Hu. Low bit-rate image compression schemes based on vector quantization. International Journal of Image and Graphics (IJIG), 5(4):745–??, October 2005. CODEN ???? ISSN 0219-4678.

Hong:2001:IST

[HWH01] P. Hong, Z. Wen, and T. S. Huang. iFACE: a 3D synthetic talking face. International Journal of Image and Graphics (IJIG), 1(1):19–??, January 2001. CODEN ???? ISSN 0219-4678.

Heng:2006:HNM

[HWL^{+}06] Pheng-Ann Heng, Tien-Tsin Wong, Ka-Man Leung, Yim-Pan Chui, and Hanqiu Sun. A haptic needle manipulation simulator for Chinese acupuncture learning and training. International Journal of Image and Graphics (IJIG), 6(2):205–??, April 2006. CODEN ???? ISSN 0219-4678.

He:2013:RAM

[HXCW13] Liwen He, Yong Xu, Yan Chen, and Jiajun Wen. Recent advance on mean shift tracking: a survey. International Journal of Image and Graphics (IJIG), 13(3): ??, July 2013. CODEN ???? ISSN 0219-4678.

Hong:2005:RPE

[HYC05] Jin-Hyuk Hong, Eun-Kyung Yun, and Sung-Bae Cho. A review of performance evaluation for biometrics systems. International Journal of Image and Graphics (IJIG), 5(3):501–536, July 2005. CODEN ???? ISSN 0219-4678. URL http:/

/ejournals.wspc.com.sg/
ijig/05/preserved-docs/
0503/S0219467805001872.
pdf.

Huang:2007:AIE

[HZG07] Qiu Huang, Gengsheng L.

Zeng, and Grant T. Gullberg. An analytical inversion of the 180^{∘} exponential Radon transform with a numerically generated kernel. International Journal of Image and Graphics (IJIG), 7 (1):71–85, January 2007. CO-DEN ???? ISSN 0219-4678.

Heickal:2015:CVB

[HZH15] Hasnain Heickal, Tao Zhang, and Md. Hasanuzzaman. Computer vision-based realtime 3D gesture recognition using depth image. International Journal of Image and Graphics (IJIG), 15(1): 1550004, January 2015. CO-DEN ???? ISSN 0219-4678.

He:2011:SBN

[HZZ11] Liqiang He, Guangyong Zhang, and Yanyan Zhang. Speeding up best neighborhood matching algorithm for high-definition image on GPU Platform. International Journal of Image and Graphics (IJIG), 11(3):315– 337, July 2011. CODEN ???? ISSN 0219-4678.

Iwasaki:2015:RRM

[II15] Fumiya Iwasaki and Hiroki Imamura. A robust recognition method for occlusion of mini tomatoes based on hue information and the curvature. International Journal of Image and Graphics (IJIG), 15(2):1540004, April 2015. CODEN ???? ISSN 0219-4678.

Islam:2023:CSD

[II23] Md. Shafiqul Islam and Rafiqul Islam. A critical survey on developed reconstruction algorithms for computed tomography imaging from a limited number of projections. International Journal of Image and Graphics (IJIG), 23(04): ??, July 2023. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467823500262.

Islam:2022:CSP

[IIU22] Rafiqul Islam, Md Shafiqul Islam, and Muhammad Shahin Uddin. Compressed sensing in parallel MRI: a review. International Journal of Image and Graphics (IJIG), 22(04): ??, July 2022. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467822500383.

Iwata:2002:DOH

[IKY02] A. Iwata, K. Kato, and K. Yamamoto. The detection of obstacles by the horizon view camera. International Journal of Image and Graphics (IJIG), 2(2):331–??, April 2002. CODEN ???? ISSN 0219-4678.

Indhumathi:2022:HAR

[IMM22] C. Indhumathi, V. Murugan, and G. Muthulakshmii. Human action recognition using spatio-temporal multiplier network and attentive correlated temporal feature. International Journal of Image and Graphics (IJIG), 22(05):??, October 2022. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467822500516.

Jameel:2021:MLT

[JAG21] Samer Kais Jameel, Sezgin Aydin, and Nebras H. Ghaeb. Machine learning techniques for corneal diseases diagnosis: a survey. International Journal of Image and Graphics (IJIG), 21(02): ??, April 2021. ISSN 0219-

                                        4678. URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0219467821500169.

Jovanova:2010:OAS

[JAPP10] Blagica Jovanova, Ivica Arsov, Marius Preda, and Françoise Preteux. Online animation system for practicing cued speech. International Journal of Image and Graphics (IJIG), 10(4):497– 512, October 2010. CODEN ???? ISSN 0219-4678.

Jha:2013:LEU

[JBC13] Rajib Kumar Jha, Prabir Kumar Biswas, and B. N. Chatterji. Logo extraction using combined discrete wavelet transform and dynamic stochastic resonance. International Journal of Image and Graphics (IJIG), 13(1):1350004:1–1350004:21, January 2013. CODEN ???? ISSN 0219-4678.

Jaimes:2001:LSV

[JC01] A. Jaimes and S.-F. Chang. Learning structured visual detectors from user input at multiple levels. International Journal of Image and Graphics (IJIG), 1(3):415–??, July 2001. CODEN ???? ISSN 0219-4678.

Jahnavi:2023:NES

[JERK23] Yeturu Jahnavi, Poongothai Elango, S. P. Raja, and P. Nagendra Kumar. A novel ensemble stacking classification of genetic variations using machine learning algorithms. International Journal of Image and Graphics (IJIG), 23(02):??, March 2023. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467823500158.

Jlassi:2010:DBV

[JH10] Hejer Jlassi and Kamel Hamrouni. Detection of blood vessels in retinal images. International Journal of Image and Graphics (IJIG), 10(1): 57–72, January 2010. CO-DEN ???? ISSN 0219-4678.

Jatain:2022:EFR

[JJ22] Rashmi Jatain and Manisha Jailia. Enhanced face recognition using adaptive local tri Weber pattern with improved deep learning architecture. International Journal of Image and Graphics (IJIG), 22(05):??, October 2022. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467822500528.

Joshi:2021:SDI

[JKM21] Anand B. Joshi, Dhanesh Kumar, and D. C. Mishra. Security of digital images based on 3D Arnold cat map and elliptic curve. International Journal of Image and Graphics (IJIG), 21(01):??, January 2021. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467821500066.

Javadi:2015:ARI

[JKW^{+}15] Mohammad Saleh Javadi, Zulaikha Kadim, Hon Hock Woon, Khairunnisa Mohamed Johari, and Norshuhada Samudin. An automatic robust image registration algorithm for aerial mapping. International Journal of Image and Graphics (IJIG), 15(2):1540002, April 2015. CODEN ???? ISSN 0219-4678.

Jindal:2009:STC

[JLS09] Manish Kumar Jindal, Gurpreet Singh Lehal, and Rajendra Kumar Sharma. On segmentation of touching characters and overlapping lines in degraded printed Gurmukhi script. International Journal of Image and Graphics (IJIG), 9(3):321–353, July 2009. CODEN ???? ISSN 0219-4678.

Jameel:2022:GSI

[JM22] Samer Kais Jameel and Jafar Majidpour. Generating spectrum images from different types — visible, thermal, and infrared based on autoencoder architecture (GVTI-AE). International Journal of Image and Graphics (IJIG), 22(01):??, January 2022. ISSN 0219-

4678. URL https://www.

worldscientific.com/doi/
10.1142/S021946782250005X.

Jameel:2023:BAI

[JM23] Samer Kais Jameel and Jafar Majidpour. BCS-AE: Integrated image compressionencryption model based on AE and Block-CS. International Journal of Image and Graphics (IJIG), 23(05):??, September 2023. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S021946782350047X.

Jacobsen:2014:IED

[JN14] C. Robert Jacobsen and Morten Nielsen. Investigation of the effects of data collection on visual stylometry. International Journal of Image and Graphics (IJIG), 14(4): 1450019, October 2014. CO-DEN ???? ISSN 0219-4678.

Jiji:2020:FSE

[JR20] G. Wiselin Jiji and A. Rajesh. Food sustenance estimation using food image. International Journal of Image and Graphics (IJIG), 20(04): ??, October 2020. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467820500345.

Jiji:2020:DST

[JRR20] G. Wiselin Jiji, A. Rajesh, and P. Johnson Durai Raj. Decision support techniques for dermatology using casebased reasoning. International Journal of Image and Graphics (IJIG), 20(03):??, July 2020. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467820500242.

Jiji:2021:CRF

[JRR21a] G. Wiselin Jiji, A. Rajesh, and P. Johnson Durai Raj. CBI + R: a fusion approach to assist dermatological diagnoses. International Journal of Image and Graphics (IJIG), 21(01):??, January 2021. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467821500054.

Jiji:2021:DPD

[JRR21b] G. Wiselin Jiji, A. Rajesh, and P. Johnson Durai Raj. Diagnosis of Parkinson’s disease using SVM classifier. International Journal of Image and Graphics (IJIG), 21 (02):??, April 2021. CO-DEN ???? ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S021946782150011X.

Jangid:2018:HDS

[JS18] Mahesh Jangid and Sumit Srivastava. Handwritten Devanagari similar character recognition by Fisher linear discriminant and pairwise classification. International Journal of Image and Graphics (IJIG), 18(4):??, October 2018. ISSN 0219-

4678. URL https://www.

worldscientific.com/doi/
10.1142/S0219467818500225.

Jagdale:2021:MRO

[JS21] Rohita H. Jagdale and Sanjeevani K. Shah. Modified rider optimization-based V channel magnification for enhanced video super resolution. International Journal of Image and Graphics (IJIG), 21(01):??, January 2021. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467821500030.

Jayaraman:2022:MFP

[JSJ22] Kumaran @ Kumar Jayaraman, Koganti Srilakshmi, and Sasikala Jayaraman. Modified flower pollination-based segmentation of medical images. International Journal of Image and Graphics (IJIG), 22(02): ??, April 2022. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467822500267.

Jiang:2012:RTO

[JSW12] Ming-Xin Jiang, Zhi-Jing Shao, and Hong-Yu Wang. Real-time object tracking algorithm with cameras mounted on moving platforms. International Journal of Image and Graphics (IJIG), 12(3):1250020, July 2012. CODEN ???? ISSN 0219-4678.

Jiji:2013:FMP

[JU13] C. V. Jiji and Ravi Krishnan Unni. Fusion of multispectral and panchromatic images based on the nonsubsampled contourlet transform. International Journal of Image and Graphics (IJIG), 13(3): ??, July 2013. CODEN ???? ISSN 0219-4678.

Jadav:2022:DSD

[JY22] Kalpesh R. Jadav and Arvind R. Yadav. Dynamic shadow detection and removal for vehicle tracking system. International Journal of Image and Graphics (IJIG), 22(05):??, October 2022. ISSN 0219-

                                        4678. URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0219467822500504.

Jasim:2014:RTC

[JZH14] Mahmood Jasim, Tao Zhang, and Md. Hasanuzzaman. A real-time computer visionbased static and dynamic hand gesture recognition system. International Journal of Image and Graphics (IJIG), 14(1–2):1450006, 2014. CO-DEN ???? ISSN 0219-4678.

Kumar:2013:GAI

[KA13] Piyush Kumar and Anupam Agrawal. GPU-accelerated interactive visualization of 3D volumetric data using CUDA. International Journal of Image and Graphics (IJIG), 13(2):??, April 2013.

CODEN ???? ISSN 0219-4678.

Khosravi:2021:NIR

[KAE21] Javanshir Khosravi, Mohammad Shams Esfand Abadi, and Reza Ebrahimpour. A novel iterative rigid image registration algorithm based on the Newton method. International Journal of Image and Graphics (IJIG), 21(02): ??, April 2021. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467821500133.

Kampke:2008:AGR

[Käm08] Thomas Kämpke. Automatic generation of 3D radar display views. International Journal of Image and Graphics (IJIG), 8(4):551–572, October 2008. CODEN ???? ISSN 0219-4678.

Kanatani:2002:MSS

[Kan02] K. Kanatani. Motion segmentation by subspace separation: Model selection and reliability evaluation. International Journal of Image and Graphics (IJIG), 2(2):179–??, April 2002. CODEN ???? ISSN 0219-4678.

Kang:2005:NVE

[Kan05] Hyung W. Kang. Nonphotorealistic virtual environment navigation from images. International Journal of Image and Graphics (IJIG), 5(2): 433–??, April 2005. CODEN ???? ISSN 0219-4678.

Karthigaikumar:2010:PPV

[KB10] P. Karthigaikumar and K. Baskaran. Partially pipelined VLSI implementation of Blowfish encryption/decryption algorithm. International Journal of Image and Graphics (IJIG), 10(3):327–341, July 2010. CODEN ???? ISSN 0219-4678.

Keefer:2015:SDI

[KB15] Robert Keefer and Nikolaos Bourbakis. A survey on document image processing methods useful for assistive technology for the blind. International Journal of Image and Graphics (IJIG), 15(1): 1550005, January 2015. CO-DEN ???? ISSN 0219-4678.

Kallel:2011:SMM

[KBL11] Mohamed Kallel, Mohamed-Salim Bouhlel, and Jean-Christophe Lapayre. Security of the medical media using a hybrid and multiple watermark technique. International Journal of Image and Graphics (IJIG), 11(1):103– 115, January 2011. CODEN ???? ISSN 0219-4678.

Khan:2022:MPS

[KD22] Rafflesia Khan and Rameswar Debnath. Morphology preserving segmentation method for occluded cell nuclei from medical microscopy image. International Journal of Image and Graphics (IJIG), 22 (02):??, April 2022. ISSN 0219-4678. URL https://

www.worldscientific.com/
doi/10.1142/S021946782250019X.

Khadilkar:2019:FIB

[KDAB19] Samrat P. Khadilkar, Sunil R.

Das, Mansour H. Assaf, and Satyendra N. Biswas. Face identification based on discrete wavelet transform and neural networks. International Journal of Image and Graphics (IJIG), 19(4):??, October 2019. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467819500220.

Kumar:2001:EQD

[KGH01] B. P. Kumar, P. Gupta, and C. J. Hwang. An efficient quadtree datastructure for neighbor finding algorithm. International Journal of Image and Graphics (IJIG), 1(4):619–??, October 2001. CODEN ???? ISSN 0219-4678.

Kamath:2009:ICS

[KGM09] Chandrika Kamath, Abel Gezahegne, and Paul Miller. Identification of coherent structures in three-dimensional simulations of a fluid-mix problem. International Journal of Image and Graphics (IJIG), 9(3):389–410, July 2009. CODEN ???? ISSN 0219-4678.

Kaur:2022:RND

[KGSA22] Swapandeep Kaur, Sheifali Gupta, Swati Singh, and Tanvi Arora. A review on natural disaster detection in social media and satellite imagery using machine learning and deep learning. International Journal of Image and Graphics (IJIG), 22 (05):??, October 2022. CO-DEN ???? ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467822500401.

Kaur:2022:DAD

[KGSG22] Swapandeep Kaur, Sheifali Gupta, Swati Singh, and Isha Gupta. Detection of Alzheimer’s disease using deep convolutional neural network. International Journal of Image and Graphics (IJIG), 22(03):??, May 2022. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S021946782140012X.

Kamath:2011:RES

[KH11] Chandrika Kamath and Omar A. Hurricane. Robust extraction of statistics from images of material fragmentation. International Journal of Image and Graphics (IJIG), 11(3):377–401, July 2011. CODEN ???? ISSN 0219-4678.

Kuijper:2013:CBM

[KH13] Arjan Kuijper and Ilkka Havukkala. Comparing bitmapped MicroRNA structure images using mutual symmetry. International Journal of Image and Graphics (IJIG), 13(3):??, July 2013. CODEN ???? ISSN 0219-4678.

Kiley:2020:WMF

[KH20] Matthew R. Kiley and Md Shafaeat Hossain. Who are my family members? A solution based on image processing and machine learning. International Journal of Image and Graphics (IJIG), 20(04):??, October 2020. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467820500333.

Khadilkar:2022:CCD

[Kha22] Samrat Pundalik Khadilkar.

Colon cancer detection using hybrid features and genetically optimized neural network classifier. International Journal of Image and Graphics (IJIG), 22(02): ??, April 2022. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467822500243.

Kitts:2002:LSM

[KHYV02] B. Kitts, K. Hetherington-Young, and M. Vrieze. Largescale mining, discovery and visualization of WWW user clickpaths. International Journal of Image and Graphics (IJIG), 2(1):21–??, January 2002. CODEN ???? ISSN 0219-4678.

Kittisuwan:2017:TID

[Kit17] P. Kittisuwan. Textural image denoising using Gumbel random vectors in Gaussian noise. International Journal of Image and Graphics (IJIG), 17(1):1750003, January 2017. CODEN ???? ISSN 0219-4678.

Kittisuwan:2018:TRD

[Kit18] Pichid Kittisuwan. Textural region denoising: Application in agriculture. International Journal of Image and Graphics (IJIG), 18(4): ??, October 2018. ISSN 0219-

                                        4678. URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0219467818500249.

Kumar:2023:HBV

[KJJ^{+}23] T. Satish Kumar, S. Jothilakshmi, Batholomew C. James, M. Prakash, N. Arulkumar, and C. Rekha. HHO-based vector quantization technique for biomedical image compression in cloud computing. International Journal of Image and Graphics (IJIG), 23 (03):??, May 2023. ISSN 0219-4678. URL https://

                                        www.worldscientific.com/
                                        doi/10.1142/S0219467822400083.

Kanatani:2004:ATC

[KK04] Kenichi Kanatani and Yasushi Kanazawa. Automatic thresholding for correspondence detection. International Journal of Image and Graphics (IJIG), 4(1):21–??, January 2004. CODEN ???? ISSN 0219-4678.

Kim:2015:DDA

[KK15] Yejin Kim and Myunggyu Kim. Data-driven approach for human locomotion generation. International Journal of Image and Graphics (IJIG), 15(2):1540001, April 2015. CODEN ???? ISSN 0219-4678.

Kate:2002:TAA

[KKB02] Rohit Jaivant Kate, Prem Kalra, and Subhashis Banerjee. Towards an automatic approach for view-dependent geometry. International Journal of Image and Graphics (IJIG), 2(3):413–??, July 2002. CODEN ???? ISSN 0219-4678.

Kaur:2023:FIB

[KKG23] Rajdeep Kaur, Rakesh Kumar, , and Meenu Gupta. Food image-based nutritional management system to overcome polycystic ovary syndrome using DeepLearning: A systematic review. International Journal of Image and Graphics (IJIG), 23(05):??, September 2023. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467823500432.

Kushwaha:2022:HAR

[KKK22] Arati Kushwaha, Ashish Khare, and Manish Khare. Human activity recognition algorithm in video sequences based on integration of magnitude and orientation information of optical flow. International Journal of Image and Graphics (IJIG), 22(01): ??, January 2022. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467822500097.

Kulkarni:2014:GBM

[KKKH14] S. B. Kulkarni, Raghavendrarao B. Kulkarni, U. P. Kulkarni, and Ravindra S. Hegadi. GLCM-based multiclass iris recognition using FKNN and KNN. International Journal of Image and Graphics (IJIG), 14(3): 1450010, July 2014. CODEN ???? ISSN 0219-4678.

Kiran:2023:MLD

[KKT^{+}23] S. Vishwa Kiran, Inderjeet Kaur, K. Thangaraj, V. Saveetha, R. Kingsy Grace, and N. Arulkumar. Machine learning with data science-enabled lung cancer diagnosis and classification using computed tomography images. International Journal of Image and Graphics (IJIG), 23(03): ??, May 2023. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467822400022.

Kim:2006:VPC

[KKW06] Jaeho Kim, Hyungseok Kim, and Kwangyun Wohn. Visibility preprocessing for complex 3D scenes using hardwarevisibility queries. International Journal of Image and Graphics (IJIG), 6(1):1–??, January 2006. CODEN ???? ISSN 0219-4678.

Ke:2002:MST

[KL02] Shih-Hao Ke and Tsu-Tian Lee. A multi-scale twostep fast search algorithm for block motion estimation. International Journal of Image and Graphics (IJIG), 2(4): 633–??, October 2002. CO-DEN ???? ISSN 0219-4678.

Kodavalla:2012:DVC

[KM12] Vijay Kumar Kodavalla and P. G. Krishna Mohan. Distributed video coding: Feedback-free architecture and implementation. International Journal of Image and Graphics (IJIG), 12(2): 1250010, April 2012. CODEN ???? ISSN 0219-4678.

Kumar:2021:ECB

[KM21] Gangavarapu Venkata Satya Kumar and Pillutla Gopala Krishna Mohan. Enhanced content-based image retrieval using information oriented angle-based local tri-directional Weber patterns. International Journal of Image and Graphics (IJIG), 21(04):??, October 2021. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467821500467.

Kumar:2022:ICB

[KM22] Gangavarapu Venkata Satya Kumar and P. G. Krishna Mohan. Improved content based image retrieval process based on deep convolutional neural network and salp swarm algorithm. International Journal of Image and Graphics (IJIG), 22(05): ??, October 2022. ISSN 0219-

                                        4678. URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0219467822500474.

Kumar:2022:DLB

[KMA22] V. Akash Kumar, Vijaya Mishra, and Monika Arora. Deep learning-based classification of malignant and benign cells in dermatoscopic images via transfer learning approach. International Journal of Image and Graphics (IJIG), 22(05):??, October 2022. ISSN 0219-

                                        4678. URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0219467822500413.

Kontakis:2017:SIC

[KMSP17] Konstantinos Kontakis, Athanasios G. Malamos, Malvina Steiakaki, and Spyros Panagiotakis. Spatial indexing of complex virtual reality scenes in the Web. International Journal of Image and Graphics (IJIG), 17(2): 1750009, April 2017. CODEN ???? ISSN 0219-4678.

Kulkarni:2010:ITU

[KNN10] Shivali D. Kulkarni, Ameya K.

Naik, and Nitin S. Nagori. 2D image transmission using bandwidth efficient mapping technique. International Journal of Image and Graphics (IJIG), 10(4):559– 573, October 2010. CODEN ???? ISSN 0219-4678.

Kumar:2023:WTO

[KNP23] K. Praveen Kumar, C. Venkata Narasimhulu, and K. Satya Prasad. 2D wavelet tree ordering based localized total variation model for efficient image restoration. International Journal of Image and Graphics (IJIG), 23(03): ??, May 2023. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467822400095.

Kanatani:2004:ADC

[KO04] Kenichi Kanatani and Naoya Ohta. Automatic detection of circular objects by ellipse growing. International Journal of Image and Graphics (IJIG), 4(1):35–??, January 2004. CODEN ???? ISSN 0219-4678.

Konvalinka:2003:VSF

[Kon03] Ira Konvalinka. Verification speed in fingerprintbased biometric systems. International Journal of Image and Graphics (IJIG), 3(3): 447–??, July 2003. CODEN ???? ISSN 0219-4678.

Kumawat:2023:IDH

[KP23] Anchal Kumawat and Sucheta Panda. An integrated double hybrid fusion approach for image smoothing. International Journal of Image and Graphics (IJIG), 23(01):??, January 2023. ISSN 0219-

                                        4678. URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0219467823500031.

Kavitha:2017:WBF

[KRS17] J. Kavitha, P. Arockia Jansi Rani, and S. Sowmyayani. Wavelet-based feature vector for shot boundary detection. International Journal of Image and Graphics (IJIG), 17 (1):1750002, January 2017. CODEN ???? ISSN 0219-4678.

Kumar:2019:IFB

[KRS19] Manoj Kumar, Anuj Rani, and Sangeet Srivastava. Image forensics based on lighting estimation. International Journal of Image and Graphics (IJIG), 19 (3):??, 2019. ISSN 0219-

                                        4678. URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0219467819500141.

Knopf:2006:FSR

[KS06] George K. Knopf and Archana P. Sangole. Freeform surface reconstruction from scattered points using a deformable spherical map. International Journal of Image and Graphics (IJIG), 6(3):341–??, July 2006. CODEN ???? ISSN 0219-4678.

Kaur:2010:HED

[KS10] Amandeep Kaur and Chandan Singh. A hybrid edge detector using fuzzy logic and mathematical morphology. International Journal of Image and Graphics (IJIG), 10(2):251–272, April 2010. CODEN ???? ISSN 0219-4678.

Kumar:2020:FFB

[KS20] Sandeep Kumar and L. Suresh.

Fruit fly-based artificial neural network classifier with kernel-based fuzzy c-means clustering for satellite image classification. International Journal of Image and Graphics (IJIG), 20(02):??, April 2020. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467820500163.

Kalaivani:2021:EBI

[KS21a] A. Kalaivani and K. Swetha.

An enhanced bidirectional insertion sort over classical insertion sort. International Journal of Image and Graphics (IJIG), 21(02): ??, April 2021. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467821500248.

Kang:2021:GIB

[KS21b] Henry Kang and Ioannis Stamoulis. Gaussian image binarization. International Journal of Image and Graphics (IJIG), 21(04):??, October 2021. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467821500479.

Kiran:2021:CSI

[KS21c] S. Shashi Kiran and K. V. Suresh. Challenges in sparse image reconstruction. International Journal of Image and Graphics (IJIG), 21 (03):??, July 2021. CO-DEN ???? ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467821500261.

Kuzhali:2021:AID

[KS21d] S. Elavaar Kuzhali and D. S. Suresh. Automated image denoising model: Contribution towards optimized internal and external basis. International Journal of Image and Graphics (IJIG), 21(04): ??, October 2021. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467821500510.

Kharma:2005:PCC

[KSG05] Nawwaf Kharma, Ching Y. Suen, and Pei F. Guo. Palmprints: a cooperative coevolutionary algorithm for clustering hand images. International Journal of Image and Graphics (IJIG), 5(3):595–616, July 2005. CODEN ???? ISSN 0219-4678. URL http:/

/ejournals.wspc.com.sg/

ijig/05/preserved-docs/
0503/S0219467805001902.
pdf.

Kamath:2021:DSI

[KSJ21] Priya R. Kamath, Kedarnath Senapati, and P. Jidesh. Despeckling of SAR images using shrinkage of two-dimensional discrete orthonormal S-transform. International Journal of Image and Graphics (IJIG), 21(02): ??, April 2021. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467821500236.

Khare:2007:DCW

[KT07] Ashish Khare and Uma Shanker Tiwary. Daubechies complex wavelet transform based technique for denoising of medical images. International Journal of Image and Graphics (IJIG), 7(4):663–687, October 2007. CODEN ???? ISSN 0219-4678.

Kameyama:2002:CRM

[KTK02] Keisuke Kameyama, Kazuo Toraichi, and Yukio Kosugi. Constructive relaxation matching involving dynamical model switching and its application to shape matching. International Journal of Image and Graphics (IJIG), 2(4):655–??, October 2002. CODEN ???? ISSN 0219-

Koyama:2004:VPR

[KTO04] Kazuhiro Koyama, Yoshiaki Tomizawa, and Minoru Okada. Vectorization and precise refractions in beam tracing. International Journal of Image and Graphics (IJIG), 4(2):325–??, April 2004. CODEN ???? ISSN 0219-4678.

Kudelski:2011:FLE

[KVSM11] Dimitri Kudelski, Sophie Viseur, Giovanni Scrofani, and Jean-Luc Mari. Feature line extraction on meshes through vertex marking and 2D topological operators. International Journal of Image and Graphics (IJIG), 11(4): 531–548, October 2011. CO-DEN ???? ISSN 0219-4678.

Kumar:2006:IST

[KZ06] Ajay Kumar and David Zhang. Integrating shape and texture for hand verification. International Journal of Image and Graphics (IJIG), 6(1):101–??, January 2006. CODEN ???? ISSN 0219-4678.

Kumar:2009:UAU

[KZ09] Ajay Kumar and David Zhang. User authentication using fusion of face and palmprint. International Journal of Image and Graphics (IJIG), 9(2):251–270, April 2009. CODEN ???? ISSN 0219-4678.

Kao:2007:EAA

[KZCP07] Chien-Min Kao, Yu Zou, Seungryong Cho, and Xiaochuan Pan. An exact analytic approach to 3D PET image reconstruction. International Journal of Image and Graphics (IJIG), 7(1):35–54, January 2007. CODEN ???? ISSN 0219-4678.

Lelandais:2005:STL

[LBP05] S. Lelandais, L. Boutte, and J. Plantier. Shape from texture: Local scales and vanishing line computation to improve results for macrotextures. International Journal of Image and Graphics (IJIG), 5(2):329–??, April 2005. CODEN ???? ISSN 0219-4678.

Lai:2012:FMI

[LC12] Shuhua Lai and Fuhua (Frank)

Cheng. Fast mesh interpolation and mesh decomposition with applications. International Journal of Image and Graphics (IJIG), 12(1): 1250006, January 2012. CO-DEN ???? ISSN 0219-4678.

Lee:2002:AIV

[Lee02] J. Lee. ABSolute: An information visualization system for decision support in sourcing. International Journal of Image and Graphics (IJIG), 2 (1):1–??, January 2002. CO-DEN ???? ISSN 0219-4678.

Leung:2008:UID

[LF08] Man-Kang Leung and Chi-Wing Fu. A user interface design for acquiring statistics from video. International Journal of Image and Graphics (IJIG), 8(3):327–349, July 2008. CODEN ???? ISSN 0219-4678.

Liu:2015:CMS

[LF15] Shiguang Liu and Dongfang Fan. Computer modeling and simulation of fruit sunscald. International Journal of Image and Graphics (IJIG), 15 (3):1550013, July 2015. CO-DEN ???? ISSN 0219-4678.

Lu:2006:BNS

[LFZL06] Baifang Lu, Zhaowei Fan, Jianmin Zheng, and Lin Li. Bio-native shape modeling and virtual reality for bio education. International Journal of Image and Graphics (IJIG), 6(2):251–??, April 2006. CODEN ???? ISSN 0219-4678.

Liu:2003:MBA

[LG03] Zhi-Qiang Liu and Jessica Y. Guo. A model-based approach to hair region segmentation. International Journal of Image and Graphics (IJIG), 3(3):481–??, July 2003. CODEN ???? ISSN 0219-4678.

Lepcha:2023:IMC

[LGD23] Dawa Chyophel Lepcha, Bhawna Goyal, and Ayush Dogra. Image matting: a comprehensive survey on techniques, comparative analysis, applications and future scope. International Journal of Image and Graphics (IJIG), 23(01):??, January 2023. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467823500110.

Luo:2001:RCD

[LGS^{+}01] Y. Luo, R. Galli, D. Sánchez, A. Bennasar, J. Fornés, J. C. Serra, J. M. Huéscar, and J. Gayà. A remote cooperative design system using interactive 3D graphics. International Journal of Image and Graphics (IJIG), 1(1): 153–??, January 2001. CO-DEN ???? ISSN 0219-4678.

Liu:2005:DBV

[LHS05] Hongmei Liu, Jiwu Huang, and Yun Q. Shi. DWTbased video data hiding robust to MPEG compression and frame loss. International Journal of Image and Graphics (IJIG), 5(1):111–??, January 2005. CODEN ???? ISSN 0219-4678.

Lu:2011:NCI

[LHT^{+}11] Wen Lu, Lihuo He, Wenjian Tang, Fei Gao, and Weilong Hou. A novel compressed images quality metric. International Journal of Image and Graphics (IJIG), 11(2):281– 292, April 2011. CODEN ???? ISSN 0219-4678.

Luo:2002:EG

[LHW02] B. Luo, E. Hancock, and R. Wilson. Eigenspaces for graphs. International Journal of Image and Graphics (IJIG), 2(2):247–??, April 2002. CODEN ???? ISSN 0219-4678.

Lienhart:2001:RTD

[Lie01] R. Lienhart. Reliable transition detection in videos: a survey and practitioner’s guide. International Journal of Image and Graphics (IJIG), 1(3):469–??, July 2001. CODEN ???? ISSN 0219-4678.

Liu:2002:ASS

[Liu02a] Zhi-Qiang Liu. Adaptive subspace self-organizing map and its applications in face recognition. International Journal of Image and Graphics (IJIG), 2(4):519–??, October 2002. CODEN ???? ISSN 0219-4678.

Liu:2002:E

[Liu02b] Zhi-Qiang Liu. Editorial. International Journal of Image and Graphics (IJIG), 2(4): 501–??, October 2002. CO-DEN ???? ISSN 0219-4678.

Liu:2010:EPP

[Liu10] Yong-Jin Liu. On the evaluation of progressive pointsampled geometry. International Journal of Image and Graphics (IJIG), 10(1): 73–91, January 2010. CO-DEN ???? ISSN 0219-4678.

Last:2002:PBA

[LK02] M. Last and A. Kandel.

Perception-based analysis of engineering experiments in the semiconductor industry. International Journal of Image and Graphics (IJIG), 2(1):107–??, January 2002. CODEN ???? ISSN 0219-4678.

Lata:2023:DDL

[LK23] Navdeep Lata and Raman Kumar. DSIT: a dynamic lightweight cryptography algorithm for securing image in IoT communication. International Journal of Image and Graphics (IJIG), 23(04): ??, July 2023. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467823500353.

Li:2001:DCS

[LLL01] G. Li, X. Li, and H. Li. Discrete clothoid spline surfaces on open meshes. International Journal of Image and Graphics (IJIG), 1(4):575–??, October 2001. CODEN ???? ISSN 0219-4678.

Lu:2018:FVS

[LLLY18] Yan Lu, Bin Liu, Weihai Li, and Nenghai Yu. Fast video stitching for aerially captured HD videos. International Journal of Image and Graphics (IJIG), 18(3): ??, July 2018. ISSN 0219-

4678. URL https://www.

worldscientific.com/doi/
10.1142/S0219467818500183.

Liang:2007:GEM

[LLMR07] Jerome Zhengrong Liang, Hongbing Lu, Dimitris N. Metaxas, and Joseph M. Reinhardt. Guest editorial: Medical imaging informatics — an information processing from image formation to visualization. International Journal of Image and Graphics (IJIG), 7(1):1–15, January 2007. CODEN ???? ISSN 0219-4678.

Lee:2005:IWR

[LLS05] Choong-Hoon Lee, Heung-Kyu Lee, and Youngho Suh. Image watermarking resistant to combined geometric and removal attacks. International Journal of Image and Graphics (IJIG), 5(1):37–??, January 2005. CODEN ???? ISSN 0219-4678.

Li:2007:SMI

[LLTY07] Xuelong Li, Jing Li, Dacheng Tao, and Yuan Yuan. A similarity metric in image searching. International Journal of Image and Graphics (IJIG), 7 (2):211–225, April 2007. CO-DEN ???? ISSN 0219-4678.

Liu:2011:IQF

[LLW^{+}11] Haoting Liu, Jie Li, Zheng Wang, Jian Cheng, Hanqing Lu, and Yan Zhao. Image quality feedback-based adaptive video definition improvement for the space manipulation task. International Journal of Image and Graphics (IJIG), 11(2):153–175, April 2011. CODEN ???? ISSN 0219-4678.

Liu:2003:NRB

[LLZ03] Wenyu Liu, Hua Li, and Guangxi Zhu. Non-rigid body interpolation based on generalized morphologic morphing. International Journal of Image and Graphics (IJIG), 3 (2):325–??, April 2003. CO-DEN ???? ISSN 0219-4678.

Lezoray:2008:GBO

[LME08] O. Lezoray, C. Meurie, and A. Elmoataz. Graph-based ordering scheme for color image filtering. International Journal of Image and Graphics (IJIG), 8(3):473–493, July 2008. CODEN ???? ISSN 0219-4678.

Lohith:2023:MBP

[LME23] M. S. Lohith, Yoga Suhas Kuruba Manjunath, and M. N. Eshwarappa. Multimodal biometric person authentication using face, ear and periocular region based on convolution neural networks. International Journal of Image and Graphics (IJIG), 23(02): ??, March 2023. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467823500195.

Lee:2012:EGC

[LN12] Jong Kwan Lee and Timothy S. Newman. Exploring GPU- and cluster-based improvements for over-sampled volume ray casting opacity correction. International Journal of Image and Graphics (IJIG), 12(4):1250025, October 2012. CODEN ???? ISSN 0219-4678.

Lakshmipathy:2006:TBI

[LNW06] Jagannathan Lakshmipathy, Wieslaw L. Nowinski, and Eric A. Wernert. Templatebased isocontouring. International Journal of Image and Graphics (IJIG), 6(2):187–??, April 2006. CODEN ???? ISSN 0219-4678.

Lin:2014:LBF

[LPL14] Jian Lin, Bo Peng, and Tianrui Li. A learning-based framework for supervised and unsupervised image segmentation evaluation. International Journal of Image and Graphics (IJIG), 14(3): 1450014, July 2014. CODEN ???? ISSN 0219-4678.

Loia:2002:FRC

[LPS02] Vincenzo Loia, Witold Pedrycz, and Salvatore Sessa. Fuzzy relation calculus in the compression and decompression of fuzzy relations. International Journal of Image and Graphics (IJIG), 2(4):617–??, October 2002. CODEN ???? ISSN 0219-4678.

Latha:2015:LFI

[LPS15] Y. L. Malathi Latha, Munaga V. N. K. Prasad, and Banoth Sammulal. Local feature integration method using phase congruency for palm print authentication. International Journal of Image and Graphics (IJIG), 15(3): 1550008, July 2015. CODEN ???? ISSN 0219-4678.

Li:2006:PMW

[LPZ06] Li Li, Zhigeng Pan, and David Zhang. A public mesh watermarking algorithm based on addition property of Fourier Transform. International Journal of Image and Graphics (IJIG), 6(1): 35–??, January 2006. CO-DEN ???? ISSN 0219-4678.

Liu:2015:HIA

[LQDZ15] Hao Liu, Hongbo Qian, Ning Dai, and Jianning Zhao. Heuristic initialization for active contour models in CT/ MRI image processing. International Journal of Image and Graphics (IJIG), 15(3): 1550010, July 2015. CODEN ???? ISSN 0219-4678.

Lang:2003:FSS

[LS03] Christian A. Lang and Ambuj K. Singh. Faster similarity search for multimedia data via query transformations. International Journal of Image and Graphics (IJIG), 3(1):3–??, January 2003. CODEN ???? ISSN 0219-4678.

Livny:2007:DAP

[LSES07] Yotam Livny, Neta Sokolovsky, and Jihad El-Sana. Dual adaptive paths for multiresolution hierarchies. International Journal of Image and Graphics (IJIG), 7(2):273– 290, April 2007. CODEN ???? ISSN 0219-4678.

Lu:2005:BIW

[LSKC05] Haiping Lu, Yun Q. Shi, Alex C. Kot, and Lihui Chen. Binary image watermarking through blurring and biased binarization. International Journal of Image and Graphics (IJIG), 5(1):67–??, January 2005. CODEN ???? ISSN 0219-4678.

Ladha:2016:RPC

[LSMC16] Shamsuddin N. Ladha, Kate Smith-Miles, and Sharat Chandran. Realistic projection on casual dual-planar surfaces with global illumination compensation. International Journal of Image and Graphics (IJIG), 16(3): 1650014, July 2016. CODEN ???? ISSN 0219-4678.

Li:2012:HIR

[LSSH12] Ping Li, Hanqiu Sun, Jianbing Shen, and Chen Huang. HDR image rerendering using GPU-based processing. International Journal of Image and Graphics (IJIG), 12(1): 1250007, January 2012. CO-DEN ???? ISSN 0219-4678.

Li:2001:CIB

[LSZ01] J. Li, H.-Y. Shum, and Y.-Q.

Zhang. On the compression of image based rendering scene: a comparison among block, reference and wavelet coders. International Journal of Image and Graphics (IJIG), 1 (1):45–??, January 2001. CO-DEN ???? ISSN 0219-4678.

Liu:2005:MSM

[LTJY05] Yong-Jin Liu, Kai Tang, Ajay Joneja, and Matthew Ming-Fai Yuen. Multiresolution shape modeling and editing in reverse engineering. International Journal of Image and Graphics (IJIG), 5(4):765–??, October 2005. CODEN ???? ISSN 0219-4678.

Lukac:2004:PBO

[Luk04]  Rastislav Lukac.  Perfor-

mance boundaries of optimal weighted median filters. International Journal of Image and Graphics (IJIG), 4(2): 157–??, April 2004. CODEN ???? ISSN 0219-4678.

Li:2006:RTM

[LW06] Xiaoying Li and Enhua Wu.

Relief texture mapping on field programmable gate array. International Journal of Image and Graphics (IJIG), 6(4):641–655, October 2006. CODEN ???? ISSN 0219-4678.

Lei:2007:MRM

[LW07] Tianhu Lei and Felix W.

Wehrli. Magnetic resonance (MR) image analysis — a statistical approach. International Journal of Image and Graphics (IJIG), 7(1): 119–141, January 2007. CO-DEN ???? ISSN 0219-4678.

Lu:2011:NTB

[LWCC11] Huchuan Lu, Dong Wang, Yen-Wei Chen, and Hao Chen. A novel texturebased multi-linear analysis algorithm for face recognition. International Journal of Image and Graphics (IJIG), 11 (4):495–508, October 2011. CODEN ???? ISSN 0219-4678.

Lu:2008:NRM

[LWLY08] Jianming Lu, Ling Wang, Yeqiu Li, and Takashi Yahagi. Noise removal for medical Xray images in multiwavelet domain. International Journal of Image and Graphics (IJIG), 8(1):25–46, January 2008. CODEN ???? ISSN 0219-4678.

Liu:2002:PAB

[LWT02] Y. Liu, C.-K. Wu, and H.-T. Tsui. A practical approach for 3D building modeling from uncalibrated video sequences. International Journal of Image and Graphics (IJIG), 2(2):287–??, April 2002. CODEN ???? ISSN 0219-4678.

Liu:2014:AIA

[LWY14] Yongmei Liu, Tanakrit Wongwitit, and Linsen Yu. Automatic image annotation based on scene analysis. International Journal of Image and Graphics (IJIG), 14(3): 1450012, July 2014. CODEN ???? ISSN 0219-4678.

Li:2017:OAS

[LWZ17] Qianwen Li, Zhihua Wei, and Cairong Zhao. Optimized automatic seeded region growing algorithm with application to ROI extraction. International Journal of Image and Graphics (IJIG), 17(4): 1750024, October 2017. CO-DEN ???? ISSN 0219-4678.

Liu:2001:AIG

[LXX^{+}01] W. Liu, T. Xin, Y. Xu, H. Shum, and H. Zhong. Artistic image generation by deviation mapping. International Journal of Image and Graphics (IJIG), 1(4):565–??, October 2001. CODEN ???? ISSN 0219-4678.

Li:2007:KGR

[LY07a] Jing Li and Yuan Yuan. Kernel GBDA for relevance feedback in image retrieval. International Journal of Image and Graphics (IJIG), 7(4): 767–776, October 2007. CO-DEN ???? ISSN 0219-4678.

Li:2007:NBA

[LY07b] Weihai Li and Yuan Yuan. A new blind attack procedure for DCT-based image encryption with spectrum learning. International Journal of Image and Graphics (IJIG), 7 (3):481–496, July 2007. CO-DEN ???? ISSN 0219-4678.

Li:2012:HBS

[LYL12] Shifeng Li, Meng Yao, and Huchuan Lu. Human body segmentation in a static image with on-line AdaBoost at multiscale superpixels. International Journal of Image and Graphics (IJIG), 12(3): 1250018, July 2012. CODEN ???? ISSN 0219-4678.

Li:2008:NAS

[LYP^{+}08] Jing Li, Tao Yang, Quan Pan, Yong-Mei Cheng, and Jun Hou. A novel algorithm for speeding up keypoint detection and matching. International Journal of Image and Graphics (IJIG), 8(4): 643–661, October 2008. CO-DEN ???? ISSN 0219-4678.

Li:2006:AMS

[LYT06] Xuelong Li, Yuan Yuan, and Dacheng Tao. Artistic mosaic series generation. International Journal of Image and Graphics (IJIG), 6(1):139–??, January 2006. CODEN ???? ISSN 0219-4678.

Lao:2009:ORA

[LZ09] Yuanwei Lao and Yuan F. Zheng. Optimal rate allocation for logo watermarking. International Journal of Image and Graphics (IJIG), 9 (1):1–25, January 2009. CO-DEN ???? ISSN 0219-4678.

Li:2015:IFB

[LZF15] Shuiwang Li, Qijun Zhao, and Xiangdong Fei. An improved AM–FM-based approach for reconstructing fingerprints from minutiae. International Journal of Image and Graphics (IJIG), 15(1): 1550007, January 2015. CO-DEN ???? ISSN 0219-4678.

Lin:2001:VCR

[LZS01] T. Lin, H. J. Zhang, and Q.- Y. Shi. Video content representation for shot retrieval and scene extraction. International Journal of Image and Graphics (IJIG), 1(3): 507–??, July 2001. CODEN ???? ISSN 0219-4678.

Li:2006:OSV

[LZW06] Bin Li, David Zhang, and Kuanquan Wang. Online signature verification by combining shape contexts and local features. International Journal of Image and Graphics (IJIG), 6(3):407–??, July 2006. CODEN ???? ISSN 0219-4678.

Liu:2007:TDI

[LZZ07] Chunsheng Liu, Tianxu Zhang, and Biyin Zhang. Turbulence degraded images restoration based on improved multiframe iterative loops and data mining. International Journal of Image and Graphics (IJIG), 7(3): 515–527, July 2007. CODEN ???? ISSN 0219-4678.

Liu:2013:ROC

[LZZ^{+}13] Hao Liu, Guanhua Zhu, Jianning Zhao, Hongbo Qian, and Ning Dai. Recognition of occlusions in CT images using a curve-based parameterization method. International Journal of Image and Graphics (IJIG), 13(4):1350018, October 2013. CODEN ???? ISSN 0219-4678.

Miyamoto:2002:FMM

[MA02] Sadaaki Miyamoto and Arnold C. Alanzado. Fuzzy c-means and mixture distribution models in the presence of noise clusters. International Journal of Image and Graphics (IJIG), 2(4):573–??, October 2002. CODEN ???? ISSN 0219-4678.

Mahjoub:2012:MOR

[MA12] Mohamed Ali Mahjoub and Malek Abbassi. 3D mesh object retrieval by discrete and continuous hidden Markov models. International Journal of Image and Graphics (IJIG), 12(4):1250022, October 2012. CODEN ???? ISSN 0219-4678.

Mondal:2021:SGS

[MA21] Md. Abdul Mannan Mondal and Mohammad Haider Ali. Self-guided stereo correspondence estimation algorithm. International Journal of Image and Graphics (IJIG), 21 (03):??, July 2021. ISSN 0219-4678. URL https://

www.worldscientific.com/
doi/10.1142/S0219467821500285.

Mashaly:2019:PAS

[Mas19] Ahmed S. Mashaly. Performance assessment of sky segmentation approaches for UAVs. International Journal of Image and Graphics (IJIG), 19(4):??, October 2019. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467819500232.

Mamic:2004:ASS

[MB04] G. Mamic and M. Bennamoun. Automated spline surface modeling and matching for recognition of freeform objects. International Journal of Image and Graphics (IJIG), 4(1):51–??, January 2004. CODEN ???? ISSN 0219-4678.

Mollah:2013:HDB

[MBN13] Ayatullah Faruk Mollah, Subhadip Basu, and Mita Nasipuri. Handheld devicebased character recognition system for camera captured images. International Journal of Image and Graphics (IJIG), 13(4):1350016, October 2013. CODEN ???? ISSN 0219-4678.

Melkemi:2013:PAD

[MCS13] Mahmoud Melkemi, Frederic Cordier, and Nickolas S. Sapidis. A provable algorithm to detect weak symmetry in a polygon. International Journal of Image and Graphics (IJIG), 13(1):1350002:1– 1350002:28, January 2013. CODEN ???? ISSN 0219-4678.

Mondal:2023:DCW

[MCT23] Saorabh Kumar Mondal, Arpitam Chatterjee, and Bipan Tudu. DCT coefficients weighting (DCTCW)- based Gray Wolf Optimization (GWO) for brightness preserving image contrast enhancement. International Journal of Image and Graphics (IJIG), 23(02):??, March 2023. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467823500183.

Moradi:2022:IDM

[MF22] Hamid Moradi and Amir Hossein Foruzan. Integration of dynamic multi-atlas and deep learning techniques to improve segmentation of the prostate in MR images. International Journal of Image and Graphics (IJIG), 22(04): ??, July 2022. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467822500310.

Manchanda:2021:ICB

[MG21] Meenu Manchanda and Deepak Gambhir. Improvement in CNN-based multifocus image fusion algorithm with triangulated fuzzy filter. International Journal of Image and Graphics (IJIG), 21(02):

??, April 2021. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467821500200.

Malathi:2023:RRC

[MG23] V. Malathi and M. P.

Gopinath. A review on rice crop disease classification using computational approach. International Journal of Image and Graphics (IJIG), 23 (03):??, May 2023. ISSN 0219-4678. URL https://

www.worldscientific.com/
doi/10.1142/S0219467822400071.

Miyazaki:2010:CDM

[MH10] Ryuji Miyazaki and Koichi Harada. Creating the displacement mapped low-level mesh and its application for CG software. International Journal of Image and Graphics (IJIG), 10(3):467–480, July 2010. CODEN ???? ISSN 0219-4678.

Manjunath:2008:DSD

[MHN08] A. V. N. Manjunath, K. G.

Hemantha, and S. Noushath. Document skew detection — a novel approach. International Journal of Image and Graphics (IJIG), 8(1):47–59, January 2008. CODEN ???? ISSN 0219-4678.

Maehara:2017:DOR

[MII17] Seiichi Maehara, Kazuo Ikeshiro, and Hiroki Imamura. A 3-dimensional object recognition method using relationship of distances and angles in feature points. International Journal of Image and Graphics (IJIG), 17(4): 1750019, October 2017. CO-DEN ???? ISSN 0219-4678.

McGraw:2016:SNN

[MKH16] Tim McGraw, Jisun Kang, and Donald Herring. Sparse non-negative matrix factorization for mesh segmentation. International Journal of Image and Graphics (IJIG), 16(1):1650004, January 2016. CODEN ???? ISSN 0219-4678.

Miyamoto:2009:FAM

[MKS^{+}09] Kentaro Miyamoto, Tetsuo Kamina, Tetsuo Sugiyama, Keisuke Kameyama, Kazuo Toraichi, and Yasuhiro Ohmiya. A function approximation method for images with grading regions. International Journal of Image and Graphics (IJIG), 9(1):101–119, January 2009. CODEN ???? ISSN 0219-4678.

Miao:2003:SHF

[MLG^{+}03] Jun Miao, Hong Liu, Wen Gao, Hongming Zhang, Gang Deng, and Xilin Chen. A system for human face and facial feature location. International Journal of Image and Graphics (IJIG), 3(3):461–??, July 2003. CODEN ???? ISSN 0219-4678.

Matsushita:2004:LSI

[MLS^{+}04] Yasuyuki Matsushita, Stephen Lin, Heung-Yeung Shum, Xin Tong, and Sing Bing Kang. Lighting and shadow interpolation using intrinsic lumigraphs. International Journal of Image and Graphics (IJIG), 4(4):585–??, October 2004. CODEN ???? ISSN 0219-4678.

Maity:2017:ODC

[MM17] Santi P. Maity and Hirak Kumar Maity. Optimality in distortion control in reversible watermarking using genetic algorithms. International Journal of Image and Graphics (IJIG), 17(3):1750013, July 2017. CODEN ???? ISSN 0219-4678.

Ma:2014:EDV

[MMP^{+}14] Ji Ma, David Murphy, Gregory Provan, Cian O’Mathuna, and Michael Hayes. The evaluation of direct volume rendering-based uncertainty visualization techniques for 3D scalar data. International Journal of Image and Graphics (IJIG), 14(4):1450017, October 2014. CODEN ???? ISSN 0219-4678.

Massoptier:2011:CGC

[MMSC11] Laurent Massoptier, Avishkar Misra, Arcot Sowmya, and Sergio Casciaro. Combining graph-cut technique and anatomical knowledge for automatic segmentation of lungs affected by diffuse parenchymal disease in HRCT images. International Journal of Image and Graphics (IJIG), 11(4):509–529, October 2011. CODEN ???? ISSN 0219-4678.

Maitra:2013:MDE

[MNB13] Indra Kanta Maitra, Sanjay Nag, and Samir K. Bandyopadhyay. Mammographic density estimation and classification using segmentation and progressive elimination method. International Journal of Image and Graphics (IJIG), 13(3):??, July 2013. CODEN ???? ISSN 0219-4678.

Montazeri:2019:MAI

[Mon19] Mitra Montazeri. Memetic algorithm image enhancement for preserving mean brightness without losing image features. International Journal of Image and Graphics (IJIG), 19(4):??, October 2019. ISSN 0219-

                                        4678. URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0219467819500207.

Mondal:2020:COD

[Mon20] Ajoy Mondal. Camouflaged object detection and tracking: a survey. International Journal of Image and Graphics (IJIG), 20(04):??, October 2020. ISSN 0219-

                                        4678. URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S021946782050028X.

Maurer:2009:PVL

[MPR09] Mauricio Rafael Maurer, Helio Pedrini, and Marco Antonio Ferreira Randi. Processing and visualization of light microscope images. International Journal of Image and Graphics (IJIG), 9(3): 369–388, July 2009. CODEN ???? ISSN 0219-4678.

Meegama:2004:FAP

[MR04] Ravinda G. N. Meegama and Jagath C. Rajapakse. Fully automated peeling technique for T1-weighted, high-quality MR head scans. International Journal of Image and Graphics (IJIG), 4(2):141–??, April 2004. CODEN ???? ISSN 0219-4678.

Marcialis:2006:DLF

[MR06] Gian Luca Marcialis and Fabio Roli. Decision-level fusion of PCA and LDA-based face recognition algorithms. International Journal of Image and Graphics (IJIG), 6 (2):293–??, April 2006. CO-DEN ???? ISSN 0219-4678.

Mukherjee:2012:CIU

[MR12] Dipti Prasad Mukherjee and Nilanjan Ray. Contour interpolation using level-set analysis. International Journal of Image and Graphics (IJIG), 12(1):1250004, January 2012. CODEN ???? ISSN 0219-4678.

Manimehalai:2016:NRR

[MR16] P. Manimehalai and P. Arockia Jansi Rani. A new robust reversible blind watermarking in wavelet-domain for color images. International Journal of Image and Graphics (IJIG), 16(2):1650006, April 2016. CODEN ???? ISSN 0219-4678.

Marchesotti:2005:VPU

[MRBL05] Luca Marchesotti, Carlo Regazzoni, Carlo Bonamico, and Fabio Lavagetto. Video processing and understanding tools for augmented multisensor perception and mobile user interaction in smart spaces. International Journal of Image and Graphics (IJIG), 5(3):679–698, July 2005. CODEN ???? ISSN 0219-4678. URL http:/

/ejournals.wspc.com.sg/
ijig/05/preserved-docs/
0503/S021946780500194X.
pdf.

Madhusudhan:2023:FVR

[MRH23] M. V. Madhusudhan, V. Udaya Rani, and Chetana Hegde. Finger vein recognition model for biometric authentication using intelligent deep learning. International Journal of Image and Graphics (IJIG), 23(03):??, May 2023. ISSN 0219-4678. URL https://

www.worldscientific.com/
doi/10.1142/S0219467822400046.

Mari:2004:CFF

[MS04] Jean-Luc Mari and Jean Sequeira. Closed free-form surface geometrical modeling a new approach with global and local characterization. International Journal of Image and Graphics (IJIG), 4(2): 241–??, April 2004. CODEN ???? ISSN 0219-4678.

Manchanda:2017:FTB

[MS17] Meenu Manchanda and Rajiv Sharma. Fuzzy transformbased fusion of multiple images. International Journal of Image and Graphics (IJIG), 17(2):1750008, April 2017. CODEN ???? ISSN 0219-4678.

Mohanta:2013:NTS

[MSC13] Partha Pratim Mohanta, Sanjoy Kumar Saha, and Bhabatosh Chanda. A novel technique for size constrained video storyboard generation using statistical run test and spanning tree. International Journal of Image and Graphics (IJIG), 13(1):1350001:1– 1350001:24, January 2013. CODEN ???? ISSN 0219-4678.

Mohideen:2013:RBC

[MT13] Abubacker Kaja Mohideen and Kuttiannan Thangavel. Region-based contrast enhancement of digital mammograms using an improved watershed segmentation. International Journal of Image and Graphics (IJIG), 13(1): 1350007:1–1350007:25, January 2013. CODEN ???? ISSN 0219-4678.

Miyazaki:2019:EVD

[MTH19] Daisuke Miyazaki, Sayaka Taomoto, and Shinsaku Hiura. Extending the visibility of dichromats using histogram equalization of hue value defined for dichromats. International Journal of Image and Graphics (IJIG), 19 (3):??, 2019. ISSN 0219-

                                        4678. URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0219467819500165.

Magnenat-Thalmann:2001:DSC

[MTVM01] N. Magnenat-Thalmann, P. Volino, and L. Moccozet. Designing and simulating clothes. International Journal of Image and Graphics (IJIG), 1(1):1– ??, January 2001. CODEN ???? ISSN 0219-4678.

Matungka:2010:EII

[MZE10] Rittavee Matungka, Yuan F. Zheng, and Robert L. Ewing. Efficient invariant image registration utilizing pre-shifted logarithmic spiral. International Journal of Image and Graphics (IJIG), 10(3):395– 421, July 2010. CODEN ???? ISSN 0219-4678.

Nikam:2009:COP

[NA09] Shankar Bhausaheb Nikam and Suneeta Agarwal. Cooccurrence probabilities and wavelet-based spoof fingerprint detection. International Journal of Image and Graphics (IJIG), 9(2):171– 199, April 2009. CODEN ???? ISSN 0219-4678.

Nakamae:2001:OPR

[Nak01] E. Nakamae. An overview of photo-realism for outdoor scenes. International Journal of Image and Graphics (IJIG), 1(1):27–??, January 2001. CODEN ???? ISSN 0219-4678.

Noyer:2005:SME

[NBB05] Jean-Charles Noyer, Christophe Boucher, and Mohammed Benjelloun. 3D structure and motion estimation from range and intensity images using particle filtering. International Journal of Image and Graphics (IJIG), 5(3):639– 661, July 2005. CODEN ???? ISSN 0219-4678. URL http:

//ejournals.wspc.com.sg/
ijig/05/preserved-docs/
0503/S0219467805001926.
pdf.

Nigam:2010:FPB

[NBRR10] Chhabi Nigam, R. Venkatesh Babu, S. Kumar Raja, and K. R. Ramakrishnan. Fragmented particles-based robust object tracking with feature fusion. International Journal of Image and Graphics (IJIG), 10(1):93–112, January 2010. CODEN ???? ISSN 0219-4678.

Nivedita:2021:ICV

[NCM^{+}21] M. Nivedita, Priyanka Chandrashekar, Shibani Mahapatra, Y. Asnath Victy Phamila, and Sathish Kumar Selvaperumal. Image captioning for video surveillance system using neural networks. International Journal of Image and Graphics (IJIG), 21(04):??, October 2021. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467821500443.

Naveenkumar:2021:STJ

[ND21] M. Naveenkumar and S. Domnic. Spatio temporal joint distance maps for skeleton-based action recognition using convolutional neural networks. International Journal of Image and Graphics (IJIG), 21 (05):??, December 2021. CO-DEN ???? ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467821400015.

Ng:2006:IFN

[NESW06] Geok See Ng, Sevki Erdogan, Daming Shi, and Abdul Wahab. Insight of fuzzy neural systems in the application of handwritten digits classification. International Journal of Image and Graphics (IJIG), 6(4):511–532, October 2006. CODEN ???? ISSN 0219-4678.

Nassar:2010:NFE

[NETM10] Hamed Nassar, Ghada El-Taweel, and Eman Mahmoud. A novel feature extraction scheme for human gait recognition. International Journal of Image and Graphics (IJIG), 10(4):575– 587, October 2010. CODEN ???? ISSN 0219-4678.

Nesme:2010:AIA

[NFP10] Matthieu Nesme, François Faure, and Yohan Payan. Accurate interactive animation of deformable models at arbitrary resolution. International Journal of Image and Graphics (IJIG), 10(2):175– 202, April 2010. CODEN ???? ISSN 0219-4678.

Nagendraswamy:2007:NMR

[NG07] H. S. Nagendraswamy and D. S. Guru. A new method of representing and matching two dimensional shapes. International Journal of Image and Graphics (IJIG), 7 (2):377–405, April 2007. CO-DEN ???? ISSN 0219-4678.

Nigam:2013:FCI

[NIT13] Akriti Nigam, Ajay Indoria, and R. C. Tripathi. Fuzzy clustering of image trademark database and preprocessing using adaptive filter and Karhunen–Loève transform. International Journal of Image and Graphics (IJIG), 13(2):??, April 2013. CODEN ???? ISSN 0219-4678.

Nguyen:2011:OTV

[NK11] Thanh Binh Nguyen and Ashish Khare. Object tracking of video sequences in curvelet domain. International Journal of Image and Graphics (IJIG), 11(1):1–20, January 2011. CODEN ???? ISSN 0219-4678.

Nandal:2020:FOA

[NK20] Savita Nandal and Sanjeev Kumar. Fractional-order anisotropic diffusion for defogging of RGB images. International Journal of Image and Graphics (IJIG), 20 (01):??, January 2020. CO-DEN ???? ISSN 0219-

                                        4678. URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0219467820500011.

Ng:2009:RFM

[NLC^{+}09] Ada N. Y. Ng, Edmund Y. Lam, Ronald Chung, Kenneth S. M. Fung, and W. H. Leung. Reference-free machine vision inspection of semiconductor die images. International Journal of Image and Graphics (IJIG), 9(1): 133–152, January 2009. CO-DEN ???? ISSN 0219-4678.

Nair:2020:RCA

[NM20] Arun T. Nair and K. Muthuvel. Research contributions with algorithmic comparison on the diagnosis of diabetic retinopathy. International Journal of Image and Graphics (IJIG), 20(04):??, October 2020. ISSN 0219-

                                        4678. URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0219467820500308.

Nguyen:2010:TBR

[NMM10] Van-Hanh Nguyen, Frederic Merienne, and Jean-Luc Martinez. Training based on real-time motion evaluation for functional rehabilitation in virtual environment. International Journal of Image and Graphics (IJIG), 10(2): 235–250, April 2010. CODEN ???? ISSN 0219-4678.

Nasrollahi:2011:SSV

[NMR11] Kamal Nasrollahi, Thomas B.

Moeslund, and Mohammad Rahmati. Summarization of surveillance video sequences using face quality assessment. International Journal of Image and Graphics (IJIG), 11 (2):207–233, April 2011. CO-DEN ???? ISSN 0219-4678.

Ngom:2018:SDP

[NNNS18] Ndeye Fatou Ngom, Cheikh H. T. C. Ndiaye, Oumar Niang, and Samba Sidibe. Shape descriptors for porous media analysis using computed tomography images. International Journal of Image and Graphics (IJIG), 18 (2):1850011, April 2018. CO-DEN ???? ISSN 0219-4678.

Nnolim:2017:FBH

[Nno17a] Uche A. Nnolim. FPGAbased hardware architecture for fuzzy homomorphic enhancement based on partial differential equations. International Journal of Image and Graphics (IJIG), 17(4):

1750022, October 2017. CO-DEN ???? ISSN 0219-4678.

Nnolim:2017:FBM

[Nno17b] Uche A. Nnolim. FPGAbased multiplier-less logbased hardware architectures for hybrid color image enhancement system. International Journal of Image and Graphics (IJIG), 17(1): 1750004, January 2017. CO-DEN ???? ISSN 0219-4678.

Nnolim:2019:FFD

[Nno19a] Uche A. Nnolim. Formulation of fractional derivative-based de-hazing algorithm and implementation on mobileembedded devices. International Journal of Image and Graphics (IJIG), 19(1): ??, January 2019. ISSN 0219-

                                        4678. URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0219467819500037.

Nnolim:2019:ISI

[Nno19b] Uche A. Nnolim. Improved single image de-hazing via sky region detection, classification and illumination refinement. International Journal of Image and Graphics (IJIG), 19(4):??, October 2019. ISSN 0219-

                                        4678. URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0219467819500219.

Nnolim:2020:PMS

[Nno20] Uche A. Nnolim. Probabilistic, multi-scale fractional tonal correction bilateral filter-based hazy image enhancement. International Journal of Image and Graphics (IJIG), 20(02): ??, April 2020. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467820500102.

Nnolim:2021:SIH

[Nno21] Uche A. Nnolim. Single image de-hazing via multiscale wavelet decomposition and estimation with fractional gradient-anisotropic diffusion fusion. International Journal of Image and Graphics (IJIG), 21(03):??, July 2021. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467821500327.

Nnolim:2022:DSE

[Nno22]  Uche A. Nnolim.  Dy-

namic selective edge-based Integer/Fractional-Order partial differential equation for degraded document image binarization. International Journal of Image and Graphics (IJIG), 22(04):??, July 2022. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467822500309.

Nnolim:2023:FOP

[Nno23] Uche A. Nnolim. Fourthorder partial differential equation framelet fusionbased colour correction and contrast enhancement for underwater images. International Journal of Image and Graphics (IJIG), 23(05):??, September 2023. ISSN 0219-

                                        4678. URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0219467823500444.

Ngo:2001:RAC

[NPZ01] C.-W. Ngo, T.-C. Pong, and H.-J. Zhang. Recent advances in content-based video analysis. International Journal of Image and Graphics (IJIG), 1(3):445–??, July 2001. CO-DEN ???? ISSN 0219-4678.

Nikesh:2020:DVB

[NR20] P. Nikesh and G. Raju. Directional vector-based skin lesion segmentation — a novel approach to skin segmentation. International Journal of Image and Graphics (IJIG), 20(03):??, July 2020. ISSN 0219-

                                        4678. URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0219467820500217.

Nisha:2020:STP

[NRK20] S. Shajun Nisha, S. P. Raja, and A. Kasthuri. Static thresholded pulse coupled neural networks in contourlet domain — a new framework for medical image denoising. International Journal of Image and Graphics (IJIG), 20(03):??, July 2020. ISSN 0219-

                                        4678. URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0219467820500254.

Naveen:2023:FDA

[NSS23] J. Naveen, Sheba Selvam, and Blessy Selvam. FO-DPSO algorithm for segmentation and detection of diabetic mellitus for ulcers. International Journal of Image and Graphics (IJIG), 23(03): ??, May 2023. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467822400113.

Nejla:2012:BIR

[NVN^{+}12] Gribaa Nejla, Noblet Vincent, Khlifa Nawres, Faisan Sylvain, and Hamrouni Kamel. Binary image registration based on geometric moments: Application to the registration of 3D segmented CT head images. International Journal of Image and Graphics (IJIG), 12(2):1250009, April 2012. CODEN ???? ISSN 0219-4678.

Oh:2001:ESQ

[OCL^{+}01] K.-M. Oh, J.-D. Choi, C.-S.

Lee, C.-J. Park, and E.-T. Lee. An efficient and simple quad edge conversion of polygonal mainfold objects. International Journal of Image and Graphics (IJIG), 1 (2):251–??, April 2001. CO-DEN ???? ISSN 0219-4678.

Omari:2020:CPA

[OJD20] Mohammed Omari, Yamina Ouled Jaafri, and Rekia Dlim. Comparative performance analysis of enhancement methods applied to Arabic manuscripts. International Journal of Image and Graphics (IJIG), 20(02): ??, April 2020. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467820500138.

Oshiro:2003:KGO

[OKI^{+}03] Osamu Oshiro, Kumi Kamada, Masataka Imura, Kunihiro Chihara, Eiji Toyota, Yasuo Ogasawara, and Fumihiko Kajiya. Kidney glomerulus observation in interactive VR space. International Journal of Image and Graphics (IJIG), 3(4):629–??, October 2003. CODEN ???? ISSN 0219-4678.

Ou:2008:LIT

[OLHL08] Chien-Min Ou, Hui-Ya Li, Wen-Jyi Hwang, and Mei-Hwa Liu. Layered image transmission with quality pre-specifiable JPEG2000. International Journal of Image and Graphics (IJIG), 8 (4):629–641, October 2008. CODEN ???? ISSN 0219-4678.

Oria:2003:VPV

[OÖ03] Vincent Oria and M. Tamer Özsu. Views or points of view on images. International Journal of Image and Graphics (IJIG), 3(1):55–??, January 2003. CODEN ???? ISSN 0219-4678.

Owada:2003:ECC

[OSN03] Shigeru Owada, Yoshihisa Shinagawa, and Frank Nielsen. Enumeration of contour correspondence. International Journal of Image and Graphics (IJIG), 3(4):609–??, October 2003. CODEN ???? ISSN 0219-4678.

Periasamy:2011:ATB

[PAD11] P. S. Periasamy, S. Athinarayanan, and K. Duraiswamy. An adaptive thresholding-based color reduction algorithm and its applications. International Journal of Image and Graphics (IJIG), 11(1):83–101, January 2011. CODEN ???? ISSN 0219-4678.

Poonkuntran:2021:SIC

[PAG^{+}21] S. Poonkuntran, P. Alli, T. M. Senthil Ganesan, S. Manthira Moorthi, and M. P. Oza. Satellite image classification using cellular automata. International Journal of Image and Graphics (IJIG), 21(02):??, April 2021. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467821500145.

Pal:2001:FIP

[Pal01] S. K. Pal. Fuzzy image processing and recognition: Uncertainty handling and applications. International Journal of Image and Graphics (IJIG), 1(2):169–??, April 2001. CODEN ???? ISSN 0219-4678.

Pal:2021:RMD

[Pal21] Tannistha Pal. A robust method for dehazing of single image with sky region detection and segmentation. International Journal of Image and Graphics (IJIG), 21(04): ??, October 2021. ISSN 0219-

                                        4678. URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0219467821500455.

Papaleo:2007:ASR

[Pap07] Laura Papaleo. An approach to surface reconstruction using uncertain data. International Journal of Image and Graphics (IJIG), 7(1): 177–194, January 2007. CO-DEN ???? ISSN 0219-4678.

Pratihar:2016:FDP

[PB16] Sanjoy Pratihar and Partha Bhowmick. Fast and direct polygonization for gray-scale images using digital straightness and exponential averaging. International Journal of Image and Graphics (IJIG), 16(2):1650007, April 2016. CODEN ???? ISSN 0219-4678.

Preetham:2023:SMR

[PB23] Anusha Preetham and Vishnu Vardhan Battu. Soil moisture retrieval using sail squirrel search optimization-based deep convolutional neural network with Sentinel-1 images. International Journal of Image and Graphics (IJIG), 23(05):??, September 2023. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467823500481.

Piva:2005:SRA

[PBC05] Alessandro Piva, Franco Bartolini, and Roberto Caldelli. Self recovery authentication of images in the DWT domain. International Journal of Image and Graphics (IJIG), 5(1):149–??, January 2005. CODEN ???? ISSN 0219-4678.

Pratikakis:2003:RMD

[PBHM03] Ioannis Pratikakis, Christian Barillot, Pierre Hellier, and Etienne Memin. Robust multiscale deformable registration of 3D ultrasound images. International Journal of Image and Graphics (IJIG), 3(4):547–??, October 2003. CODEN ???? ISSN 0219-4678.

Pankaj:2023:NDO

[PBK23] Pankaj, P. K. Bharti, , and Brajesh Kumar. A new design of occlusioninvariant face recognition using optimal pattern extraction and CNN with GRUbased architecture. International Journal of Image and Graphics (IJIG), 23(04): ??, July 2023. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467823500298.

Park:2003:SBS

[PC03a] Sanghyun Park and Wesley W. Chu. Similarity-based subsequence search in image sequence databases. International Journal of Image and Graphics (IJIG), 3(1):31–??, January 2003. CODEN ???? ISSN 0219-4678.

Prabhakar:2003:MLJ

[PC03b] Sunil Prabhakar and Rahul Chari. Minimizing latency and jitter for large-scale multimedia repositories through prefix caching. International Journal of Image and Graphics (IJIG), 3(1):95–??, January 2003. CODEN ???? ISSN 0219-4678.

Pan:2021:SID

[PCLH21] Yongpeng Pan, Zhenxue Chen, Xianming Li, and Weikai He. Single-image dehazing via dark channel prior and adaptive threshold. International Journal of Image and Graphics (IJIG), 21(04): ??, October 2021. ISSN 0219-

                                        4678. URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0219467821500534.

Pissaloux:2010:IMO

[PCV10] Edwige Pissaloux, Yong Chen, and Ramiro Velazquez. Image matching optimization via vision and inertial data fusion: Application to navigation of the visually impaired. International Journal of Image and Graphics (IJIG), 10(4):545–558, October 2010. CODEN ???? ISSN 0219-4678.

Prabhanjan:2017:DLA

[PD17] S. Prabhanjan and R. Dinesh. Deep learning approach for Devanagari script recognition. International Journal of Image and Graphics (IJIG), 17(3):1750016, July 2017. CODEN ???? ISSN 0219-4678.

Pal:2012:CAS

[PDB12] Shyamosree Pal, Rahul Dutta, and Partha Bhowmick. Circular arc segmentation by curvature estimation and geometric validation. International Journal of Image and Graphics (IJIG), 12(4): 1250024, October 2012. CO-DEN ???? ISSN 0219-4678.

Pervej:2021:RTC

[PDH^{+}21] Masud Pervej, Sabuj Das, Md. Parvez Hossain, Md. Atikuzzaman, Md. Mahin, and Muhammad Aminur Rahaman. Real-time computer vision-based Bangla vehicle license plate recognition using contour analysis and prediction algorithm. International Journal of Image and Graphics (IJIG), 21(04):??, October 2021. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S021946782150042X.

Pratt:2002:SPC

[PF02] K. B. Pratt and E. Fink. Search for patterns in compressed time series. International Journal of Image and Graphics (IJIG), 2(1):89–??, January 2002. CODEN ???? ISSN 0219-4678.

Peng:2015:UBT

[PFQ^{+}15] Chao Peng, Bing Fang, Francis Quek, Yong Cao, Seung In Park, and Liguang Xie. Upper body tracking and 3D gesture reconstruction using agent-based architecture. International Journal of Image and Graphics (IJIG), 15(4): 1550016, October 2015. CO-DEN ???? ISSN 0219-4678.

Panigrahi:2021:JBF

[PG21] Susant Kumar Panigrahi and Supratim Gupta. Joint bilateral filter for signal recovery from phase preserved curvelet coefficients for image denoising. International Journal of Image and Graphics (IJIG), 21(04):??, October 2021. ISSN 0219-

                                        4678. URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0219467821500492.

Palacios:2004:HBC

[PGW04] Rafael Palacios, Amar Gupta, and Patrick S. Wang. Handwritten bank check recognition of courtesy amounts. International Journal of Image and Graphics (IJIG), 4(2):

203–??, April 2004. CODEN ???? ISSN 0219-4678.

Padmavathi:2023:WCE

[PH23] P. Padmavathi and J. Harikiran. Wireless capsule endoscopy infected images detection and classification using MobileNetV2-BiLSTM model. International Journal of Image and Graphics (IJIG), 23(05):??, September 2023. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467823500419.

Patel:2021:ELB

[PIG21] Krina Patel, Dippal Israni, and Dweepna Garg. An efficient local block Sobolev gradient and Laplacian approach for elimination of atmospheric turbulence. International Journal of Image and Graphics (IJIG), 21(04): ??, October 2021. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467821500431.

Pereira:2001:MSM

[PK01] F. Pereira and R. Koenen.

MPEG-7: a standard for multimedia content description. International Journal of Image and Graphics (IJIG), 1 (3):527–??, July 2001. CO-DEN ???? ISSN 0219-4678.

Prakash:2014:MID

[PK14] Om Prakash and Ashish Khare. Medical image denoising based on soft thresholding using biorthogonal multiscale wavelet transform. International Journal of Image and Graphics (IJIG), 14(1– 2):1450002, 2014. CODEN ???? ISSN 0219-4678.

Pathade:2021:SMC

[PK21] Manasi Pathade and Madhuri Khambete. Supervised method for congestion detection at entry and exit corridors of public places. International Journal of Image and Graphics (IJIG), 21(02): ??, April 2021. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467821500194.

Padalkar:2022:FBS

[PK22] Ganesh R. Padalkar and Madhuri B. Khambete. Fusionbased semantic segmentation using deep learning architecture in case of very small training dataset. International Journal of Image and Graphics (IJIG), 22(05):??, October 2022. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467822500437.

Pemula:2023:GRF

[PKN23] Rambabu Pemula, Sagenela Vijaya Kumar, and C. Nagaraju. Generation of random fields for image segmentation techniques: a review. International Journal of Image and Graphics (IJIG), 23(02): ??, March 2023. ISSN 0219-

4678. URL https://www.

worldscientific.com/doi/
10.1142/S0219467823500225.

Pu:2002:NNB

[PL02] Her-Chang Pu and Chin-Teng Lin. A neural-network-based image resolution enhancement scheme for image resizing. International Journal of Image and Graphics (IJIG), 2(4):503–??, October 2002. CODEN ???? ISSN 0219-4678.

Parseh:2017:NCF

[PM17] Mohammad Javad Parseh and Mojtaba Meftahi. A new combined feature extraction method for Persian handwritten digit recognition. International Journal of Image and Graphics (IJIG), 17(2): 1750012, April 2017. CODEN ???? ISSN 0219-4678.

Porta:2009:NVM

[Por09] Marco Porta. New visualization modes for effective image presentation. International Journal of Image and Graphics (IJIG), 9(1):27–49, January 2009. CODEN ???? ISSN 0219-4678.

Poulet:2002:FVV

[Pou02] F. Poulet. Full-view: a visual data-mining environment. International Journal of Image and Graphics (IJIG), 2(1): 127–??, January 2002. CO-DEN ???? ISSN 0219-4678.

Parker:2001:RSD

[PP01] J. R. Parker and J. Pivovarov. Recognizing symbols by drawing them. International Journal of Image and Graphics (IJIG), 1(4):605–??, October 2001. CODEN ???? ISSN 0219-4678.

Pradhan:2021:ALE

[PP21] Ashis Pradhan and Mohan P. Pradhan. Automatic localization of elevation values in a poor quality topographic map. International Journal of Image and Graphics (IJIG), 21(01):??, January 2021. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467821500091.

Panwar:2021:FES

[PPS21] Kirtee Panwar, Ravindra Kumar Purwar, and Garima Srivastava. A fast encryption scheme suitable for video surveillance applications using SHA-256 hash function and 1D sine–sine chaotic map. International Journal of Image and Graphics (IJIG), 21(02):??, April 2021. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467821500224.

Pujol:2004:TSS

[PR04] Oriol Pujol and Petia Radeva. Texture segmentation by statistical deformable models.

International Journal of Image and Graphics (IJIG), 4 (3):433–??, July 2004. CO-DEN ???? ISSN 0219-4678.

Prasad:2001:TTC

[PS01] M. V. N. K. Prasad and K. K.

Shukla. Tree triangular coding image compression algorithms. International Journal of Image and Graphics (IJIG), 1(4):591–??, October 2001. CODEN ???? ISSN 0219-4678.

Prasath:2012:ADS

[PS12] V. B. Surya Prasath and Arindama Singh. An adaptive diffusion scheme for image restoration and selective smoothing. International Journal of Image and Graphics (IJIG), 12(1):1250003, January 2012. CODEN ???? ISSN 0219-4678.

Patel:2022:AET

[PS22] Alpesh M. Patel and Anil Suthar. AdaBoosted extra trees classifier for objectbased multispectral image classification of urban fringe area. International Journal of Image and Graphics (IJIG), 22(03):??, May 2022. CODEN ???? ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467821400064.

Padhy:2022:CLR

[PSDM22] Rajalaxmi Padhy, Shashwat Sourav Swain, Sanjit Kumar Dash, and Jibitesh Mishra. Classification of low-resolution satellite images using fractal augmented descriptors. International Journal of Image and Graphics (IJIG), 22(01):??, January 2022. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467822500024.

Prashar:2020:NCA

[PSJ20] Navdeep Prashar, Meenakshi Sood, and Shruti Jain. Novel cardiac arrhythmia processing using machine learning techniques. International Journal of Image and Graphics (IJIG), 20(03):??, July 2020. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467820500230.

Pajarola:2004:DFD

[PSM04] Renato Pajarola, Miguel Sainz, and Yu Meng. DMesh: Fast depth-image meshing and warping. International Journal of Image and Graphics (IJIG), 4(4):653–??, October 2004. CODEN ???? ISSN 0219-4678.

Pan:2005:FRR

[PW05] Gang Pan and Zhaohui Wu. 3D face recognition from range data. International Journal of Image and Graphics (IJIG), 5(3):573–593, July 2005. CODEN ???? ISSN 0219-4678. URL http:/

/ejournals.wspc.com.sg/
ijig/05/preserved-docs/

0503/S0219467805001884.
pdf.

Park:2007:FMO

[PW07] Chan Jong Park and Kwang Yun Wohn. Fusion of the magnetic and optical information for motion capturing. International Journal of Image and Graphics (IJIG), 7(4): 641–662, October 2007. CO-DEN ???? ISSN 0219-4678.

Peng:2008:RNP

[PXLS08] Haoyu Peng, Hua Xiong, Zhen Liu, and Jiaoying Shi. Research of nested parallel pipelines on parallel graphics rendering system. International Journal of Image and Graphics (IJIG), 8(2): 209–222, April 2008. CODEN ???? ISSN 0219-4678.

Pan:2001:LDM

[PZZ^{+}01] Z. Pan, M. Zhang, K. Zhou, C. Cheng, and J. Shi. Level of detail and multi-resolution modeling for virtual prototyping. International Journal of Image and Graphics (IJIG), 1(2):329–??, April 2001. CODEN ???? ISSN 0219-4678.

Rautaray:2013:HGR

[RA13] Siddharth Swarup Rautaray and Anupam Agrawal. Hand gesture recognition towards vocabulary and application independency. International Journal of Image and Graphics (IJIG), 13(2):??, April 2013. CODEN ???? ISSN 0219-4678.

Rajalakshmi:2022:EVP

[RA22] M. Rajalakshmi and K. Annapurani. Enhancement of vascular patterns in palm images using various image enhancement techniques for person identification. International Journal of Image and Graphics (IJIG), 22(04): ??, July 2022. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467822500322.

Raja:2019:LPC

[Raj19] S. P. Raja. Line and polygon clipping techniques on natural images — a mathematical solution and performance evaluation. International Journal of Image and Graphics (IJIG), 19 (2):??, 2019. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467819500128.

Raja:2020:WBI

[Raj20] S. P. Raja. Wavelet-based image compression encoding techniques — a complete performance analysis. International Journal of Image and Graphics (IJIG), 20(02): ??, April 2020. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467820500084.

Rajkumar:2023:DLF

[Raj23]  Rajeev Rajkumar.  Deep

learning feature extraction using attention-based DenseNet 121 for copy move forgery detection. International Journal of Image and Graphics (IJIG), 23(05):??, September 2023. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467823500420.

Rani:2012:FRU

[Ran12] J. Sheeba Rani. Face recognition using hybrid approach. International Journal of Image and Graphics (IJIG), 12 (1):1250005, January 2012. CODEN ???? ISSN 0219-4678.

Ray:2014:PRB

[Ray14] Kumar S. Ray. Pattern recognition based on fuzzy set and genetic algorithm. International Journal of Image and Graphics (IJIG), 14(3): 1450009, July 2014. CODEN ???? ISSN 0219-4678.

Ren:2002:MDT

[RCH02] Y. Ren, C. S. Chua, and Y. K. Ho. Motion detection from time-varied background. International Journal of Image and Graphics (IJIG), 2 (2):163–??, April 2002. CO-DEN ???? ISSN 0219-4678.

Rani:2022:SIP

[RDKS22] Rajneesh Rani, Renu Dhir, Deepti Kakkar, and Nonita Sharma. Script identification for printed and handwritten Indian documents: an empirical study of different feature classifier combinations. International Journal of Image and Graphics (IJIG), 22(03): ??, May 2022. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467821400118.

Ramwala:2021:RNC

[RDPP21] Ojas A. Ramwala, Smeet A. Dhakecha, Chirag N. Paunwala, and Mita C. Paunwala. Reminiscent net: Conditional GAN-based old image de-creasing. International Journal of Image and Graphics (IJIG), 21(04):??, October 2021. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467821500509.

Raj:2023:DRV

[RDS23] R. Jisha Raj, Smitha Dharan, and T. T. Sunil. Dimensionality reduction and visualization of Bharatanatyam Mudras. International Journal of Image and Graphics (IJIG), 23(01):??, January 2023. CO-DEN ???? ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467823500018.

Refaey:2015:BRL

[Ref15] Mohammed A. A. Refaey. Background ruled-lines detection and removal in fullcolored handwritten image documents. International Journal of Image and Graphics (IJIG), 15(2):1540006, April 2015. CODEN ???? ISSN 0219-4678.

Roy:2004:MCU

[RFT04] Michaël Roy, Sebti Foufou, and Frédéric Truchetet. Mesh comparison using attribute deviation metric. International Journal of Image and Graphics (IJIG), 4(1):127–??, January 2004. CODEN ???? ISSN 0219-4678.

Rahman:2005:HPR

[RI05a] M. Masudur Rahman and Seiji Ishikawa. Human posture recognition: Eigenspace tuning by a mean eigenspace. International Journal of Image and Graphics (IJIG), 5(4):825–??, October 2005. CODEN ???? ISSN 0219-4678.

Rahman:2005:ODE

[RI05b] M. Masudur Rahman and Seiji Ishikawa. Overcoming dress effect in eigenspace. International Journal of Image and Graphics (IJIG), 5(4): 811–??, October 2005. CO-DEN ???? ISSN 0219-4678.

Ravikumar:2022:MLT

[RK22] S. Ravikumar and E. Kannan. Machine learning techniques for identifying fetal risk during pregnancy. International Journal of Image and Graphics (IJIG), 22(05):

??, October 2022. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467822500450.

Ravikumar:2023:AMS

[RK23] S. Ravikumar and E. Kannan. Analysis on mental stress of professionals and pregnant women using machine learning techniques. International Journal of Image and Graphics (IJIG), 23(05): ??, September 2023. CO-DEN ???? ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467823500389.

Rodriguez:2001:GAD

[RLKB01] W. Rodriguez, M. Last, A. Kandel, and H. Bunke. Geometric approach to data mining. International Journal of Image and Graphics (IJIG), 1(2):363–??, April 2001. CODEN ???? ISSN 0219-4678.

Rema:2021:EHC

[RM21] N. R. Rema and P. Mythili. Extremely high compression and identification of fingerprint images using SA4 multiwavelet transform. International Journal of Image and Graphics (IJIG), 21(03): ??, July 2021. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467821500376.

Roy:2018:HIR

[RMG18] Aniket Roy, Arpan Kumar Maiti, and Kuntal Ghosh. An HVS inspired robust nonblind watermarking scheme in YCbCr color space. International Journal of Image and Graphics (IJIG), 18(3): ??, July 2018. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467818500158.

Roy:2014:MSU

[ROAK14] Kaushik Roy, Brian O’Connor, Foysal Ahmad, and Mohamed S. Kamel. Multibiometric system using level set, modified LBP and random forest. International Journal of Image and Graphics (IJIG), 14(3):1450013, July 2014. CODEN ???? ISSN 0219-4678.

Rao:2023:DLB

[RP23] E. Srinivasa Rao and Ch. Raghava Prasad. Deep learning-based medical image fusion using integrated joint slope analysis with probabilistic parametric steered image filter. International Journal of Image and Graphics (IJIG), 23(03): ??, May 2023. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467822400137.

Raju:2023:PAC

[RPR23] Ayalapogu Ratna Raju, Suresh Pabboju, and Rajeswara Rao Ramisetty. Performance analysis and critical review on segmentation techniques for brain tumor classification. International Journal of Image and Graphics (IJIG), 23(02):??, March 2023. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467823500237.

Ratschek:2005:SHM

[RR05] Helmut Ratschek and Jon Rokne. SCCI-hybrid methods for 2D curve tracing. International Journal of Image and Graphics (IJIG), 5(3):447– 479, July 2005. CODEN ???? ISSN 0219-4678. URL http:

//ejournals.wspc.com.sg/
ijig/05/preserved-docs/
0503/S0219467805001859.
pdf.

Rawat:2011:CBR

[RR11] Sanjay Rawat and Balasubramanian Raman. A chaosbased robust watermarking algorithm for rightful ownership protection. International Journal of Image and Graphics (IJIG), 11(4):471– 493, October 2011. CODEN ???? ISSN 0219-4678.

Ray:2013:PAD

[RR13] Kumar S. Ray and Bimal Kumar Ray. Polygonal approximation of digital curve based on reverse engineering concept. International Journal of Image and Graphics (IJIG), 13(4):1350017, October 2013.

CODEN ???? ISSN 0219-4678.

Ravikumar:2021:PPB

[RR21] K. P. Ravikumar and H. S. Manjunatha Reddy. Pixel prediction-based image steganography using crow search algorithm-based deep belief network approach. International Journal of Image and Graphics (IJIG), 21(01):??, January 2021. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467821500029.

Reddy:2023:MMM

[RRR23] Mummadi Gowthami Reddy, Palagiri Veera Narayana Reddy, and Patil Ramana Reddy. Multi-modal medical image fusion using 3stage multiscale decomposition and PCNN with adaptive arguments. International Journal of Image and Graphics (IJIG), 23(03):??, May 2023. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467822400101.

Raja:2016:EMT

[RS16] S. P. Raja and A. Suruliandi. Evaluating multiscale transform based image compression using encoding techniques. International Journal of Image and Graphics (IJIG), 16(4):1650018, October 2016. CODEN ???? ISSN 0219-4678.

Rawal:2021:DMW

[RS21] Kirti Rawal and Gaurav Sethi. Design of matched wavelet using improved genetic algorithm for heart rate variability analysis of the menstrual cycle. International Journal of Image and Graphics (IJIG), 21(03): ??, July 2021. ISSN 0219-

                                        4678. URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0219467821500303.

Rashwan:2022:MFW

[RS22] Shaheera Rashwan and Walaa Sheta. A metaheuristics framework for weighted multi-band image fusion. International Journal of Image and Graphics (IJIG), 22(02): ??, April 2022. ISSN 0219-

                                        4678. URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0219467822500164.

Ravikumar:2022:EMB

[RSG22] M. Ravikumar, B. J. Shivaprasad, and D. S. Guru. Enhancement of MRI brain images using notch filter based on discrete wavelet transform. International Journal of Image and Graphics (IJIG), 22(01):??, January 2022. ISSN 0219-

                                        4678. URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0219467822500103.

Rahul:2023:ESP

[RTP^{+}23] Vaddadi Sai Rahul, M. Tejas, N. Narayanan Prasanth, , and S. P. Raja. Early success prediction of Indian movies using subtitles: a document vector approach. International Journal of Image and Graphics (IJIG), 23(04): ??, July 2023. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467823500304.

Rukundo:2020:NEP

[Ruk20] Olivier Rukundo. Non-extra pixel interpolation. International Journal of Image and Graphics (IJIG), 20(04):??, October 2020. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S021946782050031X.

Regentova:2006:ISU

[RYLZ06] Emma Regentova, Dongsheng Yao, Shahram Latifi, and Jun Zheng. Image segmentation using ncut in the wavelet domain. International Journal of Image and Graphics (IJIG), 6(4):569– 582, October 2006. CODEN ???? ISSN 0219-4678.

Sanchez:2007:CFT

[SAA^{+}07] Danmary Sanchez, Malek Adjouadi, Nolan R. Altman, Daniel Sanchez, and Byron Bernal. Comprehensive 3D fiber tracking as a new visualization system in brain studies. International Journal of Image and Graphics (IJIG), 7(4):749–765, October 2007. CODEN ???? ISSN 0219-

Sagayam:2022:RHG

[SAG^{+}22] K. Martin Sagayam, A. Diana Andrushia, Ahona Ghosh, Omer Deperlioglu, and Ahmed A. Elngar. Recognition of hand gesture image using deep convolutional neural network. International Journal of Image and Graphics (IJIG), 22(03):??, May 2022. ISSN 0219-

                                        4678. URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0219467821400088.

Sallee:2005:MBM

[Sal05] Phil Sallee. Model-based methods for steganography and steganalysis. International Journal of Image and Graphics (IJIG), 5(1):167–??, January 2005. CODEN ???? ISSN 0219-4678.

Salehi:2022:ISB

[Sal22]  Hadi Salehi.  Image de-

speckling based on the coefficient of variation, improved guided filter, and fast bilateral filter. International Journal of Image and Graphics (IJIG), 22(04): ??, July 2022. ISSN 0219-

                                        4678. URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S021946782250036X.

Sarfraz:2004:SAC

[Sar04] Muhammad Sarfraz. Some algorithms for curve design and automatic outline capturing of images. International Journal of Image and Graphics (IJIG), 4(2):301–??, April 2004. CODEN ???? ISSN 0219-4678.

Shahrokhi:2022:ICM

[SAS22] Marziye Shahrokhi, Alireza Akoushideh, and Asadollah Shahbahrami. Image copy– move forgery detection using combination of scaleinvariant feature transform and local binary pattern features. International Journal of Image and Graphics (IJIG), 22(05):??, October 2022. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467822500486.

Sidram:2015:ENS

[SB15] M. H. Sidram and Nagappa U. Bhajantri. An exploration with novel shape signature of GMSC distance function to track the object. International Journal of Image and Graphics (IJIG), 15 (4):1550014, October 2015. CODEN ???? ISSN 0219-4678.

Shrivastava:2020:ASR

[SB20] Neeraj Shrivastava and Jyoti Bharti. Automatic seeded region growing image segmentation for medical image segmentation: a brief review. International Journal of Image and Graphics (IJIG), 20(03): ??, July 2020. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467820500187.

Shrivastava:2022:BTD

[SB22] Neeraj Shrivastava and Jyoti Bharti. Breast tumor detection in MR images based on density. International Journal of Image and Graphics (IJIG), 22 (01):??, January 2022. CO-DEN ???? ISSN 0219-

                                        4678. URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0219467822500012.

Saha:2022:TBF

[SBDN22] Priya Saha, Debotosh Bhattacharjee, Barin Kumar De, and Mita Nasipuri. A thermal blended facial expression analysis and recognition system using deformed thermal facial areas. International Journal of Image and Graphics (IJIG), 22(05):??, October 2022. ISSN 0219-

                                        4678. URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0219467822500498.

Scherff:2008:IPI

[SBH08]  Phillip-Christoph  Scherff,

George Baciu, and Jinlian Hu. Intuitive parameterized input interface for proportional reshaping of human bodies. International Journal of Image and Graphics (IJIG), 8(2):299–325, April 2008. CODEN ???? ISSN 0219-4678.

Sun:2004:IBT

[SBYG04] Jing Sun, George Baciu, Xiaobo Yu, and Mark Green.

Image-based template generation of road networks for virtual maps. International Journal of Image and Graphics (IJIG), 4(4):701–??, October 2004. CODEN ???? ISSN 0219-4678.

Singh:2016:IGZ

[SC16] Geetika Singh and Indu Chhabra. Integrating global Zernike and local discriminative HOG features for face recognition. International Journal of Image and Graphics (IJIG), 16(4):1650021, October 2016. CODEN ???? ISSN 0219-4678.

Schuckers:2003:UBB

[Sch03] Michael E. Schuckers. Using the beta-binomial distribution to assess performance of a biometric identification device. International Journal of Image and Graphics (IJIG), 3(3):523–??, July 2003. CO-DEN ???? ISSN 0219-4678.

Sudha:2023:CES

[SCM^{+}23] K. Antony Sudha, V. Cibi Castro, G. Muthulakshmi, T. Ilam Parithi, and S. P. Raja. A chaotic encryption system based on DNA coding using a deep neural network. International Journal of Image and Graphics (IJIG), 23(02):??, March 2023. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467823500201.

Slamu:2012:SFE

[SCY12] Wushour Slamu, Juming Cao, and Xinhui Yao. Sharp features extraction from point clouds. International Journal of Image and Graphics (IJIG), 12(4):1250029, October 2012. CODEN ???? ISSN 0219-4678.

Stetten:2001:AFC

[SD01] G. D. Stetten and R. Drezek. Active Fourier contour applied to real time 3D ultrasound of the heart. International Journal of Image and Graphics (IJIG), 1(4):647–??, October 2001. CODEN ???? ISSN 0219-4678.

Shreekanth:2018:NDI

[SDK18] T. Shreekanth, M. R. Deeksha, and Karthikeya R. Kaushik. A novel data independent approach for conversion of hand punched Kannada Braille script to text and speech. International Journal of Image and Graphics (IJIG), 18(2):1850010, April 2018. CODEN ???? ISSN 0219-4678.

Suryanarayana:2019:SIS

[SDY19] Gunnam Suryanarayana, Ravindra Dhuli, and Jie Yang. Single image super-resolution algorithm possessing edge and contrast preservation. International Journal of Image and Graphics (IJIG), 19(4): ??, October 2019. ISSN 0219-

                                        4678. URL https://www.

worldscientific.com/doi/
10.1142/S0219467819500244.

Snidaro:2008:THM

[SFC08] Lauro Snidaro, Gian Luca Foresti, and Luca Chittaro. Tracking human motion from monocular sequences. International Journal of Image and Graphics (IJIG), 8(3): 455–471, July 2008. CODEN ???? ISSN 0219-4678.

Su:2011:ISR

[SG11] Ya Su and Xinbo Gao. Iterative shape refinement in AAM. International Journal of Image and Graphics (IJIG), 11(1):137–151, January 2011. CODEN ???? ISSN 0219-4678.

Singh:2013:HAI

[SG13] Pankaj Pratap Singh and R. D. Garg. A hybrid approach for information extraction from high resolution satellite imagery. International Journal of Image and Graphics (IJIG), 13(2): ??, April 2013. CODEN ???? ISSN 0219-4678.

Shivsharan:2021:DRD

[SG21] Nitin Shivsharan and Sanjay Ganorkar. Diabetic retinopathy detection using optimization assisted deep learning model: Outlook on improved grey wolf algorithm. International Journal of Image and Graphics (IJIG), 21(03): ??, July 2021. ISSN 0219-

4678. URL https://www.

                                        worldscientific.com/doi/
                                        10.1142/S0219467821500352.

Sakib:2020:RDB

[SGB20] Mohammad Nazmus Sakib, Shuvashis Das Gupta, and Satyendra N. Biswas. A robust DWT-based compressed domain video watermarking technique. International Journal of Image and Graphics (IJIG), 20(01):??, January 2020. ISSN 0219-

                                        4678. URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0219467820500047.

Sharma:2023:PAS

[SGKA23] Sandhya Sharma, Sheifali Gupta, Neeraj Kumar, and Tanvi Arora. Postal automation system in Gurmukhi script using deep learning. International Journal of Image and Graphics (IJIG), 23(01): ??, January 2023. ISSN 0219-

                                        4678. URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0219467823500055.

Suraj:2011:RPC

[SGM11] M. G. Suraj, D. S. Guru, and S. Manjunath. Recognition of postal codes from fingerspelling video sequence. International Journal of Image and Graphics (IJIG), 11(1): 21–41, January 2011. CO-DEN ???? ISSN 0219-4678.

Singh:2022:EMD

[SGTG22] Swati Singh, Sheifali Gupta, Ankush Tanta, and Rupesh Gupta. Extraction of multiple diseases in apple leaf using machine learning. International Journal of Image and Graphics (IJIG), 22(03): ??, May 2022. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S021946782140009X.

Savakar:2020:ERU

[SH20] Dayanand G. Savakar and Ravi Hosur. The 3D emotion recognition using SVM and HoG features. International Journal of Image and Graphics (IJIG), 20(03): ??, July 2020. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467820500199.

Shah:2017:NMI

[Sha17] Said Khalid Shah. Nonrigid medical image registration based on curves. International Journal of Image and Graphics (IJIG), 17(2): 1750011, April 2017. CODEN ???? ISSN 0219-4678.

Shah:2022:NRI

[Sha22] Said Khalid Shah. Non-rigid image registration based on parameterized surfaces: Application to 3D cardiac motion image analysis. International Journal of Image and Graphics (IJIG), 22(04): ??, July 2022. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467822500280.

Suresh:2012:STP

[SJ12] R. M. Suresh and N. Jayalakshmi. Segmentation and tracking of progenitor cells in time lapse microscopy. International Journal of Image and Graphics (IJIG), 12(1): 1250008, January 2012. CO-DEN ???? ISSN 0219-4678.

Silva:2021:CMA

[SJ21] Rodrigo Dalvit C. Silva and Thomas R. Jenkyn. Classification of mammogram abnormalities using Legendre moments. International Journal of Image and Graphics (IJIG), 21(01):??, January 2021. ISSN 0219-

                                        4678. URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0219467821500108.

Sharma:2021:DDS

[SJK^{+}21] Moolchand Sharma, Bhanu Jain, Chetan Kargeti, Vinayak Gupta, and Deepak Gupta. Detection and diagnosis of skin diseases using residual neural networks (RESNET). International Journal of Image and Graphics (IJIG), 21(05):??, December 2021. ISSN 0219-4678. URL https:

                                        //www.worldscientific.com/
                                        doi/10.1142/S0219467821400027.

Singh:2009:IIC

[SK09] Satish Kumar Singh and Shishir Kumar. Improved image compression based on feed-forward adaptive downsampling algorithm. International Journal of Image and Graphics (IJIG), 9(4):575– 589, October 2009. CODEN ???? ISSN 0219-4678.

Saini:2015:SVB

[SK15] Deepika Saini and Sanjeev Kumar. Stereo vision-based conic reconstruction using a ray-quadric intersection. International Journal of Image and Graphics (IJIG), 15(4): 1550019, October 2015. CO-DEN ???? ISSN 0219-4678.

Siddiqui:2021:CBV

[SK21] Tanveer J. Siddiqui and Ashish Khare. Chaosbased video steganography method in discrete cosine transform domain. International Journal of Image and Graphics (IJIG), 21(02): ??, April 2021. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467821500157.

Sridhar:2022:PAT

[SK22] S. Sridhar and A. Kalaivani.

Performance analysis of twostage iterative ensemble method over random oversampling methods on multiclass imbalanced datasets. International Journal of Image and Graphics (IJIG), 22(02): ??, April 2022. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467822500255.

Skala:2006:LAV

[Ska06] Vaclav Skala. Length, area and volume computation in homogeneous coordinates. International Journal of Image and Graphics (IJIG), 6 (4):625–639, October 2006. CODEN ???? ISSN 0219-4678.

Skala:2008:ICP

[Ska08] Vaclav Skala. Intersection computation in projective space using homogeneous coordinates. International Journal of Image and Graphics (IJIG), 8(4):615–628, October 2008. CODEN ???? ISSN 0219-4678.

Soni:2020:TRE

[SKC20] Rituraj Soni, Bijendra Kumar, and Satish Chand. Text region extraction from scene images using AGF and MSER. International Journal of Image and Graphics (IJIG), 20(02):??, April 2020. ISSN 0219-

                                        4678. URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0219467820500096.

Syamsuddin:2015:FFB

[SKL15] Muhammad Rusdi Syamsuddin, Jimwook Kim, and Sung-Hee Lee. Force field-based control of dynamic particles with user-specified paths. International Journal of Image and Graphics (IJIG), 15(3): 1550009, July 2015. CODEN ???? ISSN 0219-4678.

Siri:2021:ALB

[SKL21] Sangeeta K. Siri, S. Pramod Kumar, and Mrityunjaya V. Latte. Accurate liver border identification model in CT scan images. International Journal of Image and Graphics (IJIG), 21(03): ??, July 2021. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S021946782150039X.

Soodamani:2002:FHT

[SL02] R. Soodamani and Z. Q.

Liu. A fuzzy Hough transform approach to shape description. International Journal of Image and Graphics (IJIG), 2(4):603–??, October 2002. CODEN ???? ISSN 0219-4678.

Stylianou:2009:IBF

[SL09] Georgios Stylianou and Andreas Lanitis. Image based 3D face reconstruction: a survey. International Journal of Image and Graphics (IJIG), 9 (2):217–250, April 2009. CO-DEN ???? ISSN 0219-4678.

Soderstrom:2010:RBH

[SL10] Ulrik Söderström and Haibo Li. Representation bound for human facial mimic with the aid of principal component analysis. International Journal of Image and Graphics (IJIG), 10(3):343–363, July 2010. CODEN ???? ISSN 0219-4678.

Su:2011:BIR

[SL11] Liyun Su and Ruihua Liu. Blind image restoration with modified CMA. International Journal of Image and Graphics (IJIG), 11(3):403– 413, July 2011. CODEN ???? ISSN 0219-4678.

Schenker:2002:FCG

[SLBK02] Adam Schenker, Mark Last, Horst Bunke, and Abraham Kandel. Fuzzy clustering with genetically adaptive scaling. International Journal of Image and Graphics (IJIG), 2(4):557–??, October 2002. CODEN ???? ISSN 0219-4678.

Shao:2004:APS

[SLN04] Fan Shao, Keck Voon Ling, and Wan Sing Ng. Automatic 3D prostate surface detection from TRUS with level sets. International Journal of Image and Graphics (IJIG), 4 (3):385–??, July 2004. CO-DEN ???? ISSN 0219-4678.

Shi:2009:SBI

[SLSZ09] Weiren Shi, Zuojin Li, Xin Shi, and Zhi Zhong. A survey of biologically inspired image processing for objects recognition. International Journal of Image and Graphics (IJIG), 9(4):495–510, October 2009. CODEN ???? ISSN 0219-4678.

Senapati:2014:ILE

[SM14] Ranjan Kumar Senapati and Prasanth Mankar. Improved listless embedded block partitioning algorithms for image compression. International Journal of Image and Graphics (IJIG), 14(4): 1450020, October 2014. CO-DEN ???? ISSN 0219-4678.

Sowmyayani:2021:MTC

[SM21] S. Sowmyayani and V. Murugan. Multi-type classification comparison of mammogram abnormalities. International Journal of Image and Graphics (IJIG), 21(03): ??, July 2021. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467821500273.

Spoorthi:2023:FCS

[SM23] B. Spoorthi and Shanthi Mahesh. Firefly competitive swarm optimization based hierarchical attention network for lung cancer detection. International Journal of Image and Graphics (IJIG), 23(02): ??, March 2023. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467823500171.

Samanta:2018:LTB

[SMA^{+}18] Sourav Samanta, Amartya Mukherjee, Amira S. Ashour, Nilanjan Dey, João Manuel R. S. Tavares, Wahiba Ben Abdessalem Karâa, Redha Taiar, Ahmad Taher Azar, and Aboul Ella Hassanien. Log transform based optimal image enhancement using firefly algorithm for autonomous mini unmanned aerial vehicle: An application of aerial photography. International Journal of Image and Graphics (IJIG), 18 (4):??, October 2018. CO-DEN ???? ISSN 0219-

                                        4678. URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0219467818500195.

Suruliandi:2015:EEG

[SMR15] A. Suruliandi, G. Murugeswari, and P. Arockia Jansi Rani. Empirical evaluation of generic weighted cubicle pattern and LBP derivatives for abnormality detection in mammogram images. International Journal of Image and Graphics (IJIG), 15(1): 1550001, January 2015. CO-DEN ???? ISSN 0219-4678.

Sharma:2018:CSS

[SMSK18] Himani Sharma, D. C. Mishra, R. K. Sharma, and Naveen Kumar. Cryptostego system for securing text and image data. International Journal of Image and Graphics (IJIG), 18(4):??, October 2018. ISSN 0219-

                                        4678. URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0219467818500201.

Sacht:2018:RTC

[SNdL18] Leonardo Sacht, Diego Nehab, and Rodolfo Schulz de Lima. Real-time continuous image processing. International Journal of Image and Graphics (IJIG), 18(3):

??, July 2018. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S021946781850016X.

Santosh:2012:RPS

[SNL12] K. C. Santosh, Cholwich Nattee, and Bart Lamiroy. Relative positioning of strokebased clustering: a new approach to online handwritten Devanagari character recognition. International Journal of Image and Graphics (IJIG), 12(2):1250016, April 2012. CODEN ???? ISSN 0219-4678.

Schwartz:2011:IFI

[SP11] William Robson Schwartz and Helio Pedrini. Improved fractal image compression based on robust feature descriptors. International Journal of Image and Graphics (IJIG), 11(4):571–587, October 2011. CODEN ???? ISSN 0219-4678.

Surajkanta:2020:RIA

[SP20] Yumnam Surajkanta and Shyamosree Pal. Recognition of isothetic arc using number theoretic properties. International Journal of Image and Graphics (IJIG), 20(02): ??, April 2020. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467820500114.

Shaikh:2021:STB

[SP21] Ayesha S. Shaikh and Vibha D.

Patel. Significance of the transition to biometric template protection: Explore the future. International Journal of Image and Graphics (IJIG), 21(02):??, April 2021. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S021946782150025X.

Sammeta:2023:DOS

[SP23] Naresh Sammeta and Latha Parthiban. Data ownership and secure medical data transmission using optimal multiple key-based homomorphic encryption with hyperledger blockchain. International Journal of Image and Graphics (IJIG), 23(03): ??, May 2023. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467822400034.

Solka:2002:VFA

[SPC02] J. L. Solka, C. E. Priebe, and B. T. Clark. A visualization framework for the analysis of hyperdimensional data. International Journal of Image and Graphics (IJIG), 2(1): 145–??, January 2002. CO-DEN ???? ISSN 0219-4678.

Sun:2008:BWN

[SPK08] Shusen Sun, Zhigeng Pan, and Tae-Wan Kim. Blind watermarking of non-uniform Bspline surfaces. International Journal of Image and Graphics (IJIG), 8(3):439–454, July 2008. CODEN ???? ISSN 0219-4678.

Senapati:2013:LBR

[SPM13] Ranjan K. Senapati, Umesh C.

Pati, and Kamala K. Mahapatra. Low bit rate image compression using hierarchical listless block-tree DTT algorithm. International Journal of Image and Graphics (IJIG), 13(1):1350005:1– 1350005:23, January 2013. CODEN ???? ISSN 0219-4678.

Sowmyayani:2016:ETR

[SR16] S. Sowmyayani and P. Arockia Jansi Rani. An efficient temporal redundancy transformation for wavelet based video compression. International Journal of Image and Graphics (IJIG), 16(3):1650015, July 2016. CODEN ???? ISSN 0219-4678.

Sasikaladevi:2023:CBD

[SR23a] N. Sasikaladevi and A. Revathi. Certainty-based deep fused neural network using transfer learning and adaptive movement estimation for the diagnosis of cardiomegaly. International Journal of Image and Graphics (IJIG), 23(04):??, July 2023. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S021946782350033X.

Suresh:2023:DLC

[SR23b] Gulivindala Suresh and Chanamallu Srinivasa Rao. Detection and localization of copy–move forgery in digital images: Review and challenges. International Journal of Image and Graphics (IJIG), 23(04):??, July 2023. CODEN ???? ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467823500250.

Shu:2001:APV

[SRBZ01] W. Shu, G. Rong, Z. Bian, and D. Zhang. Automatic palmprint verification. International Journal of Image and Graphics (IJIG), 1(1): 135–??, January 2001. CO-DEN ???? ISSN 0219-4678.

Shivaprasad:2022:ABT

[SRG22] B. J. Shivaprasad, M. Ravikumar, and D. S. Guru. Analysis of brain tumor using MR images: a brief survey. International Journal of Image and Graphics (IJIG), 22(02): ??, April 2022. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467822500231.

Singh:2022:OAI

[SRK22a] Rishipal Singh, Rajneesh Rani, and Aman Kamboj. An optimized approach for intra-class fruit classification using deep convolutional neural network. International Journal of Image and Graphics (IJIG), 22(03):??, May 2022. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467821400143.

Suresha:2022:KDB

[SRK22b] M. Suresha, D. S. Raghukumar, and S. Kuppa. Kumaraswamy distribution based bi-histogram equalization for enhancement of microscopic images. International Journal of Image and Graphics (IJIG), 22(01):??, January 2022. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467822500036.

Singh:2009:NWM

[SRM09] Vipula Singh, Navin Rajpal, and K. Srikanta Murthy. A neuro-wavelet model using fuzzy vector quantization for efficient image compression. International Journal of Image and Graphics (IJIG), 9 (2):299–320, April 2009. CO-DEN ???? ISSN 0219-4678.

Suryanarayana:2021:SRB

[SRY21] Gunnam Suryanarayana, Kandala N. V. P. S. Rajesh, and Jie Yang. Super-resolution based on residual learning and optimized phase stretch transform. International Journal of Image and Graphics (IJIG), 21(01):??, January 2021. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S021946782150008X.

Shang:2002:RFS

[SS02] Changjing Shang and Qiang Shen. Rough feature selection for neural network based image classification. International Journal of Image and Graphics (IJIG), 2(4):541–??, October 2002. CODEN ???? ISSN 0219-4678.

Sethi:2016:CAD

[SS16] Gaurav Sethi and B. S. Saini. Computer aided diagnosis of abdomen diseases using curvelet transform. International Journal of Image and Graphics (IJIG), 16(3): 1650013, July 2016. CODEN ???? ISSN 0219-4678.

Swamy:2021:HIC

[SS21] A. S. Anand Swamy and N. Shylashree. HDR image compression by multiscale down sampling of intensity levels. International Journal of Image and Graphics (IJIG), 21(04):??, October 2021. ISSN 0219-

                                        4678. URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0219467821500480.

Singh:2013:QBF

[SSA13] Durgesh Singh, Shivendra Shivani, and Suneeta Agarwal. Quantization-based fragile watermarking using block-wise authentication and pixel-wise recovery scheme for tampered image. International Journal of Image and Graphics (IJIG), 13 (2):??, April 2013. CODEN ???? ISSN 0219-4678.

Shashikala:2023:SSM

[SSBD23] T. D. Shashikala, B. L.

Sunitha, S. Basavarajappa, and J. P. Davim. Some studies on measurement of worn surface by digital image processing. International Journal of Image and Graphics (IJIG), 23(02):??, March 2023. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S021946782350016X.

Sadaghiani:2023:IIB

[SSF23] Abdol Vahab Khalili Sadaghiani, Samad Sheikhaei, , and Behjat Forouzandeh. Image interpolation based on 2D-DWT with novel regularitypreserving algorithm using RLS adaptive filters. International Journal of Image and Graphics (IJIG), 23(05):??, September 2023. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467823500390.

Seddik:2009:IWB

[SSFC09] Hassen Seddik, Mounir Sayadi, Farhat Fnaiech, and Mohamed Cheriet. Image watermarking based on the Hessenberg transform. International Journal of Image and Graphics (IJIG), 9(3):411–433, July 2009. CODEN ???? ISSN 0219-4678.

Singh:2015:MOT

[SSGM15] Brij Mohan Singh, Rahul Sharma, Debashis Ghosh, and Ankush Mittal. Multioriented text extraction in stylistic documents. International Journal of Image and Graphics (IJIG), 15(1): 1550002, January 2015. CO-DEN ???? ISSN 0219-4678.

Sarfraz:2017:QTN

[SSH17] Muhammad Sarfraz, Shamaila Samreen, and Malik Zawwar Hussain. A quadratic trigonometric nu spline with shape control. International Journal of Image and Graphics (IJIG), 17(3):1750015, July 2017. CODEN ???? ISSN 0219-4678.

Shen:2007:ESQ

[SSN07] Jialie Shen, John Shepherd, and Anne H. H. Ngu. An empirical study of query effectiveness improvement via multiple visual feature integration. International Journal of Image and Graphics (IJIG), 7(3):551–581, July 2007. CODEN ???? ISSN 0219-4678.

Sharma:2021:RIB

[SSP21] Urvashi Sharma, Meenakshi Sood, and Emjee Puthooran. Region of interest-based coding technique of medical images using varying grading of compression. International Journal of Image and Graphics (IJIG), 21(02):??, April 2021. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467821500212.

Sinduja:2020:EET

[SSR20] A. Sinduja, A. Suruliandi, and S. P. Raja. Empirical evaluation of texture features and classifiers for liver disease diagnosis. International Journal of Image and Graphics (IJIG), 20(02):??, April 2020. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467820500151.

Sarkar:2004:GAB

[SSS04] Biswajit Sarkar, Lokendra Kumar Singh, and Debranjan Sarkar. A genetic algorithmbased approach for detection of significant vertices for polygonal approximation of digital curves. International Journal of Image and Graphics (IJIG), 4(2):223–??, April 2004. CODEN ???? ISSN 0219-4678.

Sherstyuk:2010:SAS

[STB10] Andrei Sherstyuk, Anton Treskunov, and Benjamin Berg. Semi-automatic surface scanner for medical tangible user interfaces. International Journal of Image and Graphics (IJIG), 10(2):219– 233, April 2010. CODEN ???? ISSN 0219-4678.

Sangani:2023:PSS

[STPG23] Dhara J. Sangani, Rajesh A.

Thakker, S. D. Panchal, and Rajesh Gogineni. Pansharpening for spectral details preservation via convolutional sparse coding in nonsubsampled shearlet space. International Journal of Image and Graphics (IJIG), 23 (02):??, March 2023. CO-DEN ???? ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467823500134.

Stando:2015:P

[STSI15] Jacek Stando, Ali Dehghan Tanha, Waralak V. Siricharoen, and Yoshiro Imai. Preface. International Journal of Image and Graphics (IJIG), 15(2):1502001, April 2015. CODEN ???? ISSN 0219-4678.

Shreekanth:2015:HBT

[SU15] T. Shreekanth and V. Udayashankara. A histogram-based two-stage adaptive character segmentation for transcription of interpoint Hindi Braille to text. International Journal of Image and Graphics (IJIG), 15 (3):1550012, July 2015. CO-DEN ???? ISSN 0219-4678.

Salehi:2020:UID

[SV20] Hadi Salehi and Javad Vahidi. An ultrasound image despeckling method based on weighted adaptive bilateral filter. International Journal of Image and Graphics (IJIG), 20(03): ??, July 2020. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467820500205.

Salehi:2021:NHF

[SV21] Hadi Salehi and Javad Vahidi. A novel hybrid filter for image despeckling based on improved adaptive Wiener filter, bilateral filter and wavelet filter. International Journal of Image and Graphics (IJIG), 21(03): ??, July 2021. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467821500364.

Sasikumar:2022:CAD

[SV22] K. Sasikumar and B. Vijayakumar. Comparative analysis of different data replication strategies in cloud environment. International Journal of Image and Graphics (IJIG), 22(05):??, October 2022. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467822500425.

Salehi:2018:RHF

[SVM18] Hadi Salehi, Javad Vahidi, and Homayun Motameni. A robust hybrid filter based on evolutionary intelligence and fuzzy evaluation. International Journal of Image and Graphics (IJIG), 18(4): ??, October 2018. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467818500237.

Singh:2005:FHS

[SW05] Chandan Singh and Ekta Walia. Fast hybrid shading: an application of finite element methods in 3D rendering. International Journal of Image and Graphics (IJIG), 5(4):789–??, October 2005. CODEN ???? ISSN 0219-4678.

Song:2007:LLD

[SWC07] Mingli Song, Huiqiong Wang, and Chun Chen. Local Laplacian detail learning for face aging manipulation. International Journal of Image and Graphics (IJIG), 7(3): 463–480, July 2007. CODEN ???? ISSN 0219-4678.

Suri:2001:MSG

[SWR^{+}01] J. Suri, D. Wu, L. Reden, J. Gao, S. Singh, and S. Laxminarayan. Modeling segmentation via geometric deformable regularizers, PDE and level sets in still and motion imagery: a revisit. International Journal of Image and Graphics (IJIG), 1(4): 681–??, October 2001. CO-DEN ???? ISSN 0219-4678.

Sun:2003:DSC

[SWZM03] Jun Sun, Wenyuan Wang, Qing Zhuo, and Chengyuan Ma. Discriminatory sparse coding and its application to face recognition. International Journal of Image and Graphics (IJIG), 3(3):503–??, July 2003. CODEN ???? ISSN 0219-4678.

Sun:2005:CSS

[SYLC05] Qibin Sun, Shuiming Ye, Ching-Yung Lin, and Shih-Fu Chang. A crypto signature scheme for image authentication over wireless channel. International Journal of Image and Graphics (IJIG), 5(1): 135–??, January 2005. CO-DEN ???? ISSN 0219-4678.

Tomar:2016:LRP

[TA16] Divya Tomar and Sonali Agarwal. Leaf recognition for plant classification using direct acyclic graph based multi-class least squares twin support vector machine. International Journal of Image and Graphics (IJIG), 16(3): 1650012, July 2016. CODEN ???? ISSN 0219-4678.

Thushara:2023:GTB

[TAJ23] A. Thushara, C. Ushadevi Amma, and Ansamma John. Graph theory-based brain network connectivity analysis and classification of Alzheimer’s disease. International Journal of Image and Graphics (IJIG), 23(03): ??, May 2023. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S021946782240006X.

Talbi:2018:SIW

[TB18] Mourad Talbi and Med Salim Bouhlel. Secure image watermarking based on LWT and SVD. International Journal of Image and Graphics (IJIG), 18(4):??, October 2018. ISSN 0219-

                                        4678. URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0219467818500213.

Tan:2014:SVI

[TC14] Cheen-Hau Tan and Lap-Pui Chau. Single viewpoint image-driven simplification. International Journal of Image and Graphics (IJIG), 14(3):1450008, July 2014. CODEN ???? ISSN 0219-4678.

Tan:2006:IVE

[TCB06] Jiacheng Tan, Gordon J. Clapworthy, and Igor R. Belousov. The integration of a virtual environment and 3D modeling tools in a networked robot system. International Journal of Image and Graphics (IJIG), 6(1):65–??, January 2006. CODEN ???? ISSN 0219-4678.

Thakoor:2008:AVO

[TG08] Ninad Thakoor and Jean X. Gao. Automatic video object extraction with camera in motion. International Journal of Image and Graphics (IJIG), 8(4):573–600, October 2008. CODEN ???? ISSN 0219-4678.

Taylor-Hell:2005:SAR

[THBR05] Julia F. Taylor-Hell, Gladimir V. G. Baranoski, and Jon G. Rokne. State of the art in the realistic modeling of plant venation systems. International Journal of Image and Graphics (IJIG), 5(3):663–678, July 2005. CODEN ???? ISSN 0219-4678. URL http:/

/ejournals.wspc.com.sg/
ijig/05/preserved-docs/
0503/S0219467805001938.
pdf.

Tian:2007:DEV

[THW^{+}07] Yun Tian, Chongyang Hao, Yi Wang, Guiqing He, Jun Wei, Haitiao Zhao, and Benhua Zhao. Dynamic extraction for VOI from CT images based on volume rendering. International Journal of Image and Graphics (IJIG), 7(1):195–209, January 2007. CODEN ???? ISSN 0219-4678.

Tico:2003:RAS

[TK03] Marius Tico and Pauli Kuosmanen. A remote authentication system using fingerprints. International Journal of Image and Graphics (IJIG), 3(3):425–??, July 2003. CODEN ???? ISSN 0219-4678.

Tohka:2004:DMA

[TM04] Jussi Tohka and Jouni M.

Mykkänen. Deformable mesh for automated surface extraction from noisy images. International Journal of Image and Graphics (IJIG), 4(3): 405–??, July 2004. CODEN ???? ISSN 0219-4678.

Toulminet:2004:FAS

[TMB04] Gwenaëlle Toulminet, Stéphane Mousset, and Abdelaziz Bensrhair. Fast and accurate stereo vision-based estimation of 3D position and axial motion of road obstacles. International Journal of Image and Graphics (IJIG), 4 (1):99–??, January 2004. CO-DEN ???? ISSN 0219-4678.

Theobalt:2004:CFT

[TMSS04] Christian Theobalt, Marcus A. Magnor, Pascal Schüler, and Hans-Peter Seidel. Combining 2D feature tracking and volume reconstruction for online videobased human motion capture. International Journal of Image and Graphics (IJIG), 4(4):563–??, October 2004. CODEN ???? ISSN 0219-4678.

Tadamura:2001:FRW

[TQJN01] K. Tadamura, X. Qin, G. Jiao, and E. Nakamae. Fast rendering water surface for outdoor scenes. International Journal of Image and Graphics (IJIG), 1(2):313–??, April 2001. CODEN ???? ISSN 0219-4678.

Treuillet:2010:OIV

[TR10] Sylvie Treuillet and Eric Royer. Outdoor/indoor vision-based localization for blind pedestrian navigation assistance. International Journal of Image and Graphics (IJIG), 10(4):481–496, October 2010. CODEN ???? ISSN 0219-4678.

Tamburo:2001:GOP

[TS01] R. J. Tamburo and G. D.

Stetten. Gradient-oriented profiles for boundary parameterization and their application to core atoms towards shape analysis. International Journal of Image and Graphics (IJIG), 1(4):659–??, October 2001. CODEN ???? ISSN 0219-4678.

Tissainayagam:2002:PMA

[TS02] P. Tissainayagam and D. Suter.

Performance measures for assessing contour trackers. International Journal of Image and Graphics (IJIG), 2(2): 343–??, April 2002. CODEN ???? ISSN 0219-4678.

Tunga:2021:UNM

[TSAS21] P. Prakash Tunga, Vipula Singh, V. Sri Aditya, and N. Subramanya. U-Net model-based classification and description of brain tumor in MRI images. International Journal of Image and Graphics (IJIG), 21(05):??, December 2021. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467821400052.

Tanveer:2022:EIP

[TSAS22] Muhammad Tanveer, Tariq Shah, Asif Ali, and Dawood Shah. An efficient image privacy-preserving scheme based on mixed chaotic map and compression. International Journal of Image and Graphics (IJIG), 22(02):??, April 2022. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467822500206.

Tripathi:2020:SAH

[TSJ20] Kirthi Tripathi, Harsh Sohal, and Shruti Jain. Statistical analysis of HRV parameters for the detection of arrhythmia. International Journal of Image and Graphics (IJIG), 20(04):??, October 2020. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467820500369.

Tudavekar:2023:STI

[TSP23] Gajanan Tudavekar, Santosh S. Saraf, and Sanjay R. Patil. Spatio-temporal inference transformer network for video inpainting. International Journal of Image and Graphics (IJIG), 23(01):??, January 2023. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467823500079.

Takizawa:2003:RML

[TSY^{+}03] Hotaka Takizawa, Kanae Shigemoto, Shinji Yamamoto, Tohru Matsumoto, Yukio Tateno, Takeshi Iinuma, and Mitsuomi Matsumoto. A recognition method of lung nodule shadows in X-ray CT images using 3D object models. International Journal of Image and Graphics (IJIG), 3(4):533–??, October 2003. CODEN ???? ISSN 0219-4678.

Tian:2017:ISC

[TSZ^{+}17] Chunwei Tian, Guanglu Sun, Qi Zhang, Weibing Wang, Teng Chen, and Yuan Sun. Integrating sparse and collaborative representation classifications for image classification. International Journal of Image and Graphics (IJIG), 17(2):1750007, April 2017. CODEN ???? ISSN 0219-4678.

Tizhoosh:2006:RCA

[TT06] Hamid R. Tizhoosh and Graham W. Taylor. Reinforced contrast adaptation. International Journal of Image and Graphics (IJIG), 6(3):377–??, July 2006. CODEN ???? ISSN 0219-4678.

Tsechpenakis:2004:PBB

[TTK04] Gabriel Tsechpenakis, Nicolas Tsapatsoulis, and Stefanos Kollias. Probabilistic boundary-based contour tracking with snakes in natural cluttered video sequences. International Journal of Image and Graphics (IJIG), 4 (3):469–??, July 2004. CO-DEN ???? ISSN 0219-4678.

Tao:2007:RCO

[TTL07] Ji Tao, Yap-Peng Tan, and Wenmiao Lu. Robust color object tracking with application to people monitoring. International Journal of Image and Graphics (IJIG), 7 (2):227–254, April 2007. CO-DEN ???? ISSN 0219-4678.

Tangelder:2003:PMR

[TV03] Johan W. H. Tangelder and Remco C. Veltkamp. Polyhedral model retrieval using weighted point sets. International Journal of Image and Graphics (IJIG), 3(1):209–??, January 2003. CODEN ???? ISSN 0219-4678.

Takimoto:2014:FOE

[TYKO14] Hironori Takimoto, Hitoshi Yamauchi, Mitsuyoshi Kishihara, and Kensuke Okubo. Foreground object extraction based on interactive color saliency map. International Journal of Image and Graphics (IJIG), 14(1–2):1450007, 2014. CODEN ???? ISSN 0219-4678.

Takimoto:2012:ICP

[TYM12] Hironori Takimoto, Seiki Yoshimori, and Yasue Mitsukura. Invisible calibration pattern for print-and-scan data hiding based on human visual perception. International Journal of Image and Graphics (IJIG), 12(4): 1250026, October 2012. CO-DEN ???? ISSN 0219-4678.

Tamaki:2002:CDI

[TYO02] T. Tamaki, T. Yamamura, and N. Ohnishi. Correcting distortion of image by image registration with the implicit function theorem. International Journal of Image and Graphics (IJIG), 2(2):309–??, April 2002. CODEN ???? ISSN 0219-4678.

Unwin:2002:DMG

[UHW02] A. R. Unwin, H. Hofmann, and A. F. X. Wilhelm. Direct manipulation graphics for data mining. International Journal of Image and Graphics (IJIG), 2(1):49–??, January 2002. CODEN ???? ISSN 0219-4678.

Urolagin:2011:KAR

[UPR11] Siddhaling Urolagin, K. V.

Prema, and N. V. Subba Reddy. Kannada alphabets recognition with application to Braille translation. International Journal of Image and Graphics (IJIG), 11(3): 293–314, July 2011. CODEN ???? ISSN 0219-4678.

Verma:2021:CBD

[VDBK21] Parag Verma, Ankur Dumka, Anuj Bhardwaj, and Mukesh Chandra Kestwal. Classifying breast density in mammographic images using waveletbased and fine-tuned sensory neural networks. International Journal of Image and Graphics (IJIG), 21(05):??, December 2021. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467821400040.

Voloshynovskiy:2005:ITD

[VDKP05] Sviatoslav Voloshynovskiy, Frederic Deguillaume, Oleksiy Koval, and Thierry Pun. Information-theoretic datahiding: Recent achievements and open problems. International Journal of Image and Graphics (IJIG), 5(1):5–??, January 2005. CODEN ???? ISSN 0219-4678.

Veinidis:2019:EDH

[VDPT19] Christos Veinidis, Antonios Danelakis, Ioannis Pratikakis, and Theoharis Theoharis. Effective descriptors for human action retrieval from 3D mesh sequences. International Journal of Image and Graphics (IJIG), 19 (3):??, 2019. ISSN 0219-

                                        4678. URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0219467819500189.

Vasuki:2009:EAC

[VG09] S. Vasuki and L. Ganesan. An efficient approach to color image segmentation using intermediate features of maximum overlap wavelet transform in peak finding algorithm. International Journal of Image and Graphics (IJIG), 9(1): 67–76, January 2009. CO-DEN ???? ISSN 0219-4678.

Verma:2007:CSI

[VH07] Nishchal K. Verma and M. Hanmandlu. Color segmentation via improved mountain clustering technique. International Journal of Image and Graphics (IJIG), 7(2):407–426, April 2007. CODEN ???? ISSN 0219-4678.

Vishnuvardhan:2023:MIF

[VJ23] Veruva Vishnuvardhan and T. Jaya. Medical image fusion using ECNN- and OMBObased adaptive weighted fusion rule. International Journal of Image and Graphics (IJIG), 23(05):??, September 2023. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467823500468.

Vandana:2022:ARB

[VK22] Vandana and Navdeep Kaur.

Analytical review of biometric technology employing vivid modalities. International Journal of Image and Graphics (IJIG), 22(01):??, January 2022. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467822500048.

Vaidya:2022:HEM

[VP22] Bhaumik Vaidya and Chirag Paunwala. Hardware efficient modified CNN architecture for traffic sign detection and recognition. International Journal of Image and Graphics (IJIG), 22(02): ??, April 2022. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467822500176.

Vyas:2009:GTI

[VR09] Vibha S. Vyas and Priti P. Rege. Geometric transform invariant texture analysis with modified Chebyshev moments based algorithm. International Journal of Image and Graphics (IJIG), 9(4):559–574, October 2009. CODEN ???? ISSN 0219-4678.

Vasantharaj:2023:ABI

[VRH^{+}23] A. Vasantharaj, Pacha Shoba Rani, Sirajul Huque, K. S. Raghuram, R. Ganeshkumar, and Sebahadin Nasir Shafi. Automated brain imaging diagnosis and classification model using rat swarm optimization with deep learning based capsule network. International Journal of Image and Graphics (IJIG), 23 (03):??, May 2023. CO-DEN ???? ISSN 0219-

                                        4678. URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0219467822400010.

Verma:2013:ISP

[VS13] Nishchal K. Verma and Shikha Singh. Image sequence prediction using ANN and RBFNN. International Journal of Image and Graphics (IJIG), 13(2):??, April 2013. CODEN ???? ISSN 0219-4678.

Varghese:2022:DLB

[VS22] Prathibha Varghese and G. Arockia Selva Saroja.

Deep learning-based hexrep neural network for convergence free with operator’s efficacy in hexagonal image processing. International Journal of Image and Graphics (IJIG), 22(04):??, July 2022. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467823500328.

Verma:2018:IDU

[VSK18] Atul Kumar Verma, Barjinder Singh Saini, and Taranjit Kaur. Image denoising using Alexander fractional hybrid filter. International Journal of Image and Graphics (IJIG), 18(1):1850003, January 2018. CODEN ???? ISSN 0219-4678.

Wagdy:2021:DDI

[WAI21] Marian Wagdy, Khaild Amin, and Mina Ibrahim. Dewarping document image techniques: Survey and comparative study. International Journal of Image and Graphics (IJIG), 21(03):??, July 2021. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467821500315.

Wagdy:2022:DCM

[WAI22] Marian Wagdy, Khaild Amin, and Mina Ibrahim. Detection and correction of multiwarping document image. International Journal of Image and Graphics (IJIG), 22(04):

??, July 2022. ISSN 0219-

                                        4678. URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0219467822500346.

Wayman:2001:FBA

[Way01] J. L. Wayman. Fundamentals of biometric authentication technologies. International Journal of Image and Graphics (IJIG), 1(1):93–??, January 2001. CODEN ???? ISSN 0219-4678.

Wang:2002:ROM

[WB02] Yaming Wang and George Baciu. Robust object matching using a modified version of the Hausdorff measure. International Journal of Image and Graphics (IJIG), 2(3): 361–??, July 2002. CODEN ???? ISSN 0219-4678.

Wang:2011:ISC

[WC11] Wei Wang and Chi-Kit Ronald Chung. Image segmentation with complementary use of edge and region information. International Journal of Image and Graphics (IJIG), 11 (4):549–570, October 2011. CODEN ???? ISSN 0219-4678.

Wang:2012:IED

[WCB12] Xingyuan Wang, Zhifeng Chen, and Xuemei Bao. An improved edge-directed image interpolation algorithm. International Journal of Image and Graphics (IJIG), 12(4): 1250023, October 2012. CO-DEN ???? ISSN 0219-4678.

Wang:2006:DLL

[WCC^{+}06] Zhenlan Wang, Chee-Kong Chui, Yiyu Cai, Chuan-Heng Ang, and Swee-Hin Teoh. Dynamic linear level octree-based volume rendering methods for interactive microsurgical simulation. International Journal of Image and Graphics (IJIG), 6(2): 155–??, April 2006. CODEN ???? ISSN 0219-4678.

Wang:2003:CBI

[WCFT03] Zhiyong Wang, Zheru Chi, Dagan Feng, and Ah Chung Tsoi. Content-based image retrieval with relevance feedback using adaptive processing of tree-structure image representation. International Journal of Image and Graphics (IJIG), 3(1):119–??, January 2003. CODEN ???? ISSN 0219-4678.

Wu:2001:MTD

[WCRW01] P. Wu, Y. Choi, Y. M. Ro, and C. S. Won. MPEG-7 texture descriptors. International Journal of Image and Graphics (IJIG), 1(3):547–??, July 2001. CODEN ???? ISSN 0219-4678.

Wang:2002:SVW

[WCX02] L. Wang, K. L. Chan, and X.-J. Xiong. A sub-vector weighting scheme for image retrieval with relevance feedback. International Journal of Image and Graphics (IJIG), 2(2):199–??, April 2002. CODEN ???? ISSN 0219-4678.

Wang:2005:NMA

[WCXL05] Qiang Wang, Hongbo Chen, Xiaorong Xu, and Haiyan Liu. A newly modified algorithm of Hough Transform for line detection. International Journal of Image and Graphics (IJIG), 5(4):715–??, October 2005. CODEN ???? ISSN 0219-4678.

Wang:2011:NTL

[WCY11] Xing-Yuan Wang, Zhi-Feng Chen, and Jiao-Jiao Yun. A novel two-level color image retrieval method. International Journal of Image and Graphics (IJIG), 11(3):339– 353, July 2011. CODEN ???? ISSN 0219-4678.

Wu:2010:SBM

[WCZC10] Jie Wu, Jiabi Chen, Xuelong Zhang, and Jinghai Chen. The segmentation of brain MR images using reformative expectation–maximization algorithm. International Journal of Image and Graphics (IJIG), 10(2):289–297, April 2010. CODEN ???? ISSN 0219-4678.

Wang:2023:TIN

[WD23] Xin Wang and Xiaogang Dong. Time image de-noising method based on sparse regularization. International Journal of Image and Graphics (IJIG), 23(05):??, September 2023. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467825500093.

Wei:2013:SMO

[Wei13] Jie Wei. Small moving object detection from infra-red sequences. International Journal of Image and Graphics (IJIG), 13(3):??, July 2013. CODEN ???? ISSN 0219-4678.

Wei:2014:MEM

[Wei14] Jie Wei. On Markov Earth Mover’s distance. International Journal of Image and Graphics (IJIG), 14(4): 1450016, October 2014. CO-DEN ???? ISSN 0219-4678.

Wang:2005:AEM

[WF05] Xiuying Wang and David Dagan Feng. Automatic elastic medical image registration based on image intensity. International Journal of Image and Graphics (IJIG), 5(2): 351–??, April 2005. CODEN ???? ISSN 0219-4678.

Worring:2001:IRC

[WG01] M. Worring and T. Gevers. Interactive retrieval of color images. International Journal of Image and Graphics (IJIG), 1(3):387–??, July 2001. CODEN ???? ISSN 0219-4678.

Wan:2003:OTA

[WGPC03] Huagen Wan, Shuming Gao, Qunsheng Peng, and Yiyu Cai. Optimization techniques for assembly planning of complex models in large-scale virtual environments. International Journal of Image and Graphics (IJIG), 3(2):379–??, April 2003. CODEN ???? ISSN 0219-4678.

Wang:2013:VBA

[WHC13] Shuenn-Jyi Wang, Chung-Kai Hsieh, and Tsorng-Lin Chia. Video-based approach for detecting prohibited activities on sporting courts. International Journal of Image and Graphics (IJIG), 13(4): 1350019, October 2013. CO-DEN ???? ISSN 0219-4678.

Wang:2012:MIS

[WL12] Haijun Wang and Ming Liu. Medical images segmentation using active contours driven by global and local image fitting energy. International Journal of Image and Graphics (IJIG), 12(2):1250015, April 2012. CODEN ???? ISSN 0219-4678.

Wang:2007:ISE

[WLW^{+}07] Zhenguo Wang, Christopher S. D. Lee, Wayne C. Waltzer, Zhijia Yuan, and Yingtian Pan. Interpixel-shifted endoscopic optical coherence tomography for in vivo bladder cancer diagnosis. International Journal of Image and Graphics (IJIG), 7(1): 105–117, January 2007. CO-DEN ???? ISSN 0219-4678.

Wang:2010:CSF

[WLZ10] Xiaoping Wang, Shenglan Liu, and Liyan Zhang. Constructing surface features through deformation. International Journal of Image and Graphics (IJIG), 10(1): 41–56, January 2010. CO-DEN ???? ISSN 0219-4678.

Wong:2009:PPP

[Won09] Alexander Wong. PECSI:

a practical perceptuallyenhanced compression framework for still images. International Journal of Image and Graphics (IJIG), 9(4): 511–529, October 2009. CO-DEN ???? ISSN 0219-4678.

Wang:2009:MCM

[WPGF09] J. Wang, N. V. Patel, W. I.

Grosky, and F. Fotouhi. Moving camera moving object segmentation in compressed video sequences. International Journal of Image and Graphics (IJIG), 9(4):609– 627, October 2009. CODEN ???? ISSN 0219-4678.

Wang:2002:OOB

[WS02] Taehyung Wang and Phillip C. Y. Sheu. An objectoriented BSP tree algorithm for hidden surface removal. International Journal of Image and Graphics (IJIG), 2 (3):395–??, July 2002. CO-DEN ???? ISSN 0219-4678.

Wachs:2002:CFS

[WSL02] Juan Wachs, Helman Stern, and Mark Last. Color face segmentation using a fuzzy min-max neural network. International Journal of Image and Graphics (IJIG), 2(4): 587–??, October 2002. CO-DEN ???? ISSN 0219-4678.

Wang:2003:EDE

[WSO^{+}03] Jiahui Wang, Hideo Saito, Shinji Ozawa, Tomohiro Kuwahara, Toyonobu Yamashita, and Motoji Takahashi. Extraction of dermoepidermal surface from 3D volumetric images of human skin. International Journal of Image and Graphics (IJIG), 3(4):589–??, October 2003. CODEN ???? ISSN 0219-4678.

Wang:2005:PVR

[WSW05] Wencheng Wang, Hanqiu Sun, and Enhua Wu. Projective volume rendering by excluding occluded voxels. International Journal of Image and Graphics (IJIG), 5(2): 413–??, April 2005. CODEN ???? ISSN 0219-4678.

Wang:2009:JSN

[WSZ09] Xingyuan Wang, Wenjing Song, and Lixian Zou. Julia set of the Newton method for solving some complex exponential equation. International Journal of Image and Graphics (IJIG), 9(2):153– 169, April 2009. CODEN ???? ISSN 0219-4678.

Werghi:2007:LTD

[WXS07] Naoufel Werghi, Yijun Xiao, and Paul Siebert. Labelling of three dimensional human body scans: a topological approach. International Journal of Image and Graphics (IJIG), 7(2):255–272, April 2007. CODEN ???? ISSN 0219-4678.

Wu:2018:IAB

[WXS^{+}18] Zhaoqi Wu, Reziwanguli Xiamixiding, Atul Sajjanhar, Juan Chen, and Quan Wen. Image appearance-based facial expression recognition. International Journal of Image and Graphics (IJIG), 18 (2):1850012, April 2018. CO-DEN ???? ISSN 0219-4678.

Wang:2012:TTC

[WYL12] Dong Wang, Gang Yang, and Huchuan Lu. Tri-tracking: Combining three independent views for robust visual tracking. International Journal of Image and Graphics (IJIG), 12(3):1250021, July 2012. CODEN ???? ISSN 0219-4678.

Wang:2018:IST

[WYZR18] Huan Wang, Fei Yang, Congcong Zhang, and Mingwu Ren. Infrared small target detection based on patch image model with local and global analysis. International Journal of Image and Graphics (IJIG), 18(1):1850002, January 2018. CODEN ???? ISSN 0219-4678.

Wei:2018:AQE

[WZF18] Jie Wei, Lin Zhang, and Bingmei M. Fu. Automatic quantification of endothelial nitric oxide levels in a microvessel with and without tumor cell adhesion. International Journal of Image and Graphics (IJIG), 18(1):1850001, January 2018. CODEN ???? ISSN 0219-4678.

Wang:2011:QBC

[WZPW11] Yuqing Wang, Ming Zhu, Haochen Pang, and Yong Wang. Quaternion based color image quality assessment index. International Journal of Image and Graphics (IJIG), 11(2):195–206, April 2011. CODEN ???? ISSN 0219-4678.

Wang:2007:BCR

[WZW07] Kongqiao Wang, Yanming Zou, and Hao Wang. 1D bar code reading on camera phones. International Journal of Image and Graphics (IJIG), 7(3):529–550, July 2007. CODEN ???? ISSN 0219-4678.

Wang:2004:DCI

[WZZG04] Yongmei Michelle Wang, Jingdan Zhang, Zhunping Zhang, and Baining Guo. Directional coherence interpolation for three-dimensional gray-level images. International Journal of Image and Graphics (IJIG), 4(4):535–??, October 2004. CODEN ???? ISSN 0219-4678.

Xin:2008:IST

[XHB08] Binjie Xin, Jinlian Hu, and George Baciu. An imaging system for textile surface profile based on silhouette image analysis. International Journal of Image and Graphics (IJIG), 8(4):601–613, October 2008. CODEN ???? ISSN 0219-4678.

Xu:2007:MCM

[XL07] Xinyu Xu and Baoxin Li.

Multiple class multiple-instance learning and its application to image categorization. International Journal of Image and Graphics (IJIG), 7(3): 427–444, July 2007. CODEN ???? ISSN 0219-4678.

Xu:2012:FRB

[XLW12] Gang Xu, Huchuan Lu, and Zunyi Wang. Face recognition based on GPPBTF and LBP with classifier fusion. International Journal of Image and Graphics (IJIG), 12(2): 1250011, April 2012. CODEN ???? ISSN 0219-4678.

Xu:2003:FSB

[XPZ03] Weiwei Xu, Zhigeng Pan, and Mingmin Zhang. Footprint sampling-based motion editing. International Journal of Image and Graphics (IJIG), 3 (2):311–??, April 2003. CO-DEN ???? ISSN 0219-4678.

Xu:2007:FRG

[XTLY07] Dong Xu, Dacheng Tao, Xuelong Li, and Shuicheng Yan. Face recognition — a generalized marginal Fisher analysis approach. International Journal of Image and Graphics (IJIG), 7(3):583–591, July 2007. CODEN ???? ISSN 0219-4678.

Xu:2006:NCW

[XWB06] Qing Xu, Wei Wang, and Shiqiang Bao. A new computational way to Monte Carlo global illumination. International Journal of Image and Graphics (IJIG), 6(1):23–??, January 2006. CODEN ???? ISSN 0219-4678.

Yan:2001:HFI

[Yan01] H. Yan. Human face image processing techniques. International Journal of Image and Graphics (IJIG), 1(2): 197–??, April 2001. CODEN ???? ISSN 0219-4678.

Yang:2008:VBT

[Yan08] Shiueng-Bien Yang. Variablebranch tree-structured residual vector quantization. International Journal of Image and Graphics (IJIG), 8(1): 61–80, January 2008. CO-DEN ???? ISSN 0219-4678.

Yang:2013:SHB

[YCL13] Jia Yang, Chee Kooi Chan, and Ameersing Luximon. A survey on 3D human body modeling for interactive fashion design. International Journal of Image and Graphics (IJIG), 13(4):1350021, October 2013. CODEN ???? ISSN 0219-4678.

Yamada:2017:SBW

[YCVdS17] F. A. A. Yamada, L. W. X.

Cejnog, M. B. Vieira, and R. L. S. da Silva. A shapebased weighting strategy applied to the covariance estimation on ICP. International Journal of Image and Graphics (IJIG), 17(4): 1750021, October 2017. CO-DEN ???? ISSN 0219-4678.

Yang:2009:DEO

[YCWZ09] Gaobo Yang, Weiwei Chen, Xiao Jing Wang, and Zhaoyang Zhang. Dense estimation of optical flow field within the MPEG-2 compressed domain. International Journal of Image and Graphics (IJIG), 9 (3):435–448, July 2009. CO-DEN ???? ISSN 0219-4678.

Yang:2011:MHC

[YDZ11] Wenjia Yang, Lihua Dou, and Juan Zhan. A multihistogram clustering approach toward Markov random field for foreground segmentation. International Journal of Image and Graphics (IJIG), 11(1):65–81, January 2011. CODEN ???? ISSN 0219-4678.

Ye:2020:MFM

[YF20] Dan Ye and Chiou-Shann Fuh. 3D morphable face model for face animation. International Journal of Image and Graphics (IJIG), 20(01): ??, January 2020. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467820500035.

Yadav:2020:EIA

[YG20] Navneet Yadav and Navdeep Goel. An effective imageadaptive hybrid watermarking scheme with transform coefficients. International Journal of Image and Graphics (IJIG), 20(01):??, January 2020. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467820500023.

Yu:2006:CIR

[YHZ06] Shengsheng Yu, Chaobing Huang, and Jingli Zhou. Color image retrieval based on color-texture-edge feature histograms. International Journal of Image and Graphics (IJIG), 6(4):583–598, October 2006. CODEN ???? ISSN 0219-4678.

Yau:2008:VSR

[YKA08] Wai Chee Yau, Dinesh Kant Kumar, and Sridhar Poosapadi Arjunan. Visual speech recognition using dynamic features and support vector machines. International Journal of Image and Graphics (IJIG), 8(3):419–437, July 2008. CODEN ???? ISSN 0219-4678.

Yang:2010:DTC

[YL10] Zhaohui Yang and Naimin Li. Detection of tongue crack based on distant gradient and prior knowledge. International Journal of Image and Graphics (IJIG), 10(2):273– 288, April 2010. CODEN ???? ISSN 0219-4678.

Yang:2012:IRB

[YL12] Shiueng-Bien Yang and Ting-Wen Liang. Image restoration based on smooth gray-level detection and line prediction method for large missing regions. International Journal of Image and Graphics (IJIG), 12(2):1250013, April 2012. CODEN ???? ISSN 0219-4678.

Yuan:2016:ADM

[YL16] Jianjun Yuan and Lipei Liu.

Anisotropic diffusion model based on a new diffusion coefficient and fractional order differential for image denoising. International Journal of Image and Graphics (IJIG), 16(1):1650003, January 2016. CODEN ???? ISSN 0219-4678.

Yang:2008:VTA

[YLPC08] Tao Yang, Jing Li, Quan Pan, and Yong-Mei Cheng. Visual tracking with automatic confident region extraction. International Journal of Image and Graphics (IJIG), 8(3): 369–381, July 2008. CODEN ???? ISSN 0219-4678.

Yilmaz:2003:IBI

[YMA03] Ulaş Yilmaz, Adem Yasar Mülayim, and Volkan Atalay. An image-based inexpensive 3D scanner. International Journal of Image and Graphics (IJIG), 3(2):235–??, April 2003. CODEN ???? ISSN 0219-4678.

Yu:2004:SCS

[YMG04] Jingyi Yu, Leonard McMillan, and Steven Gortler. Surface camera (SCAM) light field rendering. International Journal of Image and Graphics (IJIG), 4(4):605–??, October 2004. CODEN ???? ISSN 0219-4678.

Yang:2004:UAR

[YPYW04] Ruigang Yang, Marc Pollefeys, Hua Yang, and Greg Welch. A unified approach to real-time, multi-resolution, multi-baseline 2D view synthesis and 3D depth estimation using commodity graphics hardware. International Journal of Image and Graphics (IJIG), 4(4):627–??, October 2004. CODEN ???? ISSN 0219-4678.

Yang:2011:IMV

[YTW11] J. X. Yang, D. M. Tan, and H. R. Wu. An impairment metric for video temporal fluctuation measure. International Journal of Image and Graphics (IJIG), 11(2): 251–264, April 2011. CODEN ???? ISSN 0219-4678.

Yang:2011:STV

[YW11] Fuzheng Yang and Shuai Wan. Spatial-temporal video quality assessment based on two-level temporal pooling. International Journal of Image and Graphics (IJIG), 11 (2):235–249, April 2011. CO-DEN ???? ISSN 0219-4678.

Yang:2012:LTI

[YW12] Shiueng-Bien Yang and Chi-Feng Wu. Locating text in images based on the smooth gray-level detection. International Journal of Image and Graphics (IJIG), 12(4): 1250027, October 2012. CO-DEN ???? ISSN 0219-4678.

Yuan:2014:MGD

[YW14] Yongfeng Yuan and Kuanquan Wang. A mixed Gauss and directional distance filter for fiber direction tracking. International Journal of Image and Graphics (IJIG), 14(1–2):1450001, 2014. CO-DEN ???? ISSN 0219-4678.

Yuan:2016:PMM

[YW16] Jianjun Yuan and Jianjun Wang. Perona–Malik model with a new diffusion coefficient for image denoising. International Journal of Image and Graphics (IJIG), 16(2): 1650011, April 2016. CODEN ???? ISSN 0219-4678.

You:2002:SSB

[YZ02] Jane You and David Zhang.

Smart sensor: an on-board image processing system for real-time remote sensing. International Journal of Image and Graphics (IJIG), 2(3): 481–??, July 2002. CODEN ???? ISSN 0219-4678.

Yazdi:2004:IFP

[YZ04] Mehran Yazdi and Andre Zaccarin. Inter-frame prediction of medical and videophone sequences: a deformable triangle-based approach. International Journal of Image and Graphics (IJIG), 4(3): 453–??, July 2004. CODEN ???? ISSN 0219-4678.

Zhou:2005:COV

[ZDT05] Jianlong Zhou, Andreas Döring, and Klaus D. Tönnies. Control of object visibility in volume rendering — a distance-based approach. International Journal of Image and Graphics (IJIG), 5(4): 699–??, October 2005. CO-DEN ???? ISSN 0219-4678.

Zaghloul:2021:FSI

[ZH21] Rawan I. Zaghloul and Hazem Hiary. A fast single image fog removal method using geometric mean histogram equalization. International Journal of Image and Graphics (IJIG), 21 (01):??, January 2021. CO-DEN ???? ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467821500017.

Zhang:2002:IER

[Zha02] Y. J. Zhang. Image engineering and related publications. International Journal of Image and Graphics (IJIG), 2 (3):441–??, July 2002. CO-DEN ???? ISSN 0219-4678.

Zhang:2022:MDM

[Zha22] Qi Zhang. Medical data and mathematically modeled implicit surface real-rime visualization in Web browsers. International Journal of Image and Graphics (IJIG), 22 (04):??, July 2022. CO-DEN ???? ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467822500279.

Zarif:2021:VIC

[ZI21] Sameh Zarif and Mina Ibrahim. Video inpainting: A complete framework. International Journal of Image and Graphics (IJIG), 21(03): ??, July 2021. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467821500388.

Zhang:2012:ICB

[ZL12] Lihe Zhang and Zhenzhen Liu. Image cosegmentation based on local and global level set methods. International Journal of Image and Graphics (IJIG), 12(3): 1250019, July 2012. CODEN ???? ISSN 0219-4678.

Zhang:2007:ADM

[ZLG^{+}07] Liang Zhang, Qingping Lin, Robert Gay, Guangbin Huang, and Norman Neo. An autonomous decentralized multi-server framework for large scale collaborative virtual environments. International Journal of Image and Graphics (IJIG), 7(2):353– 375, April 2007. CODEN ???? ISSN 0219-4678.

Zhao:2005:HPR

[ZLK05] Jianhui Zhao, Ling Li, and Kwoh Chee Keong. Human posture reconstruction and animation from monocular images based on genetic algorithms. International Journal of Image and Graphics (IJIG), 5(2):371–??, April 2005. CODEN ???? ISSN 0219-4678.

Zhou:2007:IRD

[ZLL^{+}07] Huiyu Zhou, Tangwei Liu, Faquan Lin, Yusheng Pang, and Ji Wu. Image restoration and detail preservation by Bayesian estimation. International Journal of Image and Graphics (IJIG), 7(3): 497–514, July 2007. CODEN ???? ISSN 0219-4678.

Zheng:2009:MTM

[ZLY09] Liying Zheng, Kuifeng Liu, and Lei Yu. Multilevel thresholding method based on normalized cut. International Journal of Image and Graphics (IJIG), 9(4):531–540, October 2009. CODEN ???? ISSN 0219-4678.

Zou:2008:RCM

[Zou08] Jie Zou. Rose curve model and an analytical solution for estimating its parameters. International Journal of Image and Graphics (IJIG), 8(1): 99–108, January 2008. CO-DEN ???? ISSN 0219-4678.

Zongqing:2009:NFB

[ZQJ09] Lu Zongqing, Liao Qingmin, and Pei Jihong. A nonlinear filtering based optical flow computation. International Journal of Image and Graphics (IJIG), 9(1):121–132, January 2009. CODEN ???? ISSN 0219-4678.

Zhou:2006:SAR

[ZS06] Hong Zhou and Ray Seyfarth. Semi automatic registration of partially overlapped aerial images via pattern search method. International Journal of Image and Graphics (IJIG), 6(3):393–??, July 2006. CODEN ???? ISSN 0219-4678.

Zhi:2012:IIA

[ZS12] Zhanjiang Zhi and Yi Sun.

An image inpainting algorithm based on energy minimization. International Journal of Image and Graphics (IJIG), 12(3):1250017, July 2012. CODEN ???? ISSN 0219-4678.

Zhong:2006:HDS

[ZSAS06] Yongmin Zhong, Bijan Shirinzadeh, Gursel Alici, and Julian Smith. Haptic deformation simulation with Poisson equation. International Journal of Image and Graphics (IJIG), 6(3):445–??, July 2006. CODEN ???? ISSN 0219-4678.

Zhang:2007:LGB

[ZSCG07] Wenchao Zhang, Shiguang Shan, Xilin Chen, and Wen Gao. Local Gabor binary patterns based on mutual information for face recognition. International Journal of Image and Graphics (IJIG), 7(4):777–793, October 2007. CODEN ???? ISSN 0219-4678.

Zhong:2008:RDB

[ZSS08] Yongmin Zhong, Bijan Shirinzadeh, and Julian Smith. Reaction-diffusion based deformable object simulation. International Journal of Image and Graphics (IJIG), 8 (2):265–280, April 2008. CO-DEN ???? ISSN 0219-4678.

Zhang:2008:HAP

[ZSXL08] Fan Zhang, Hanqiu Sun, Leilei Xu, and Kitlun Lee. Hardware-accelerated parallelsplit shadow maps. International Journal of Image and Graphics (IJIG), 8(2):223– 241, April 2008. CODEN ???? ISSN 0219-4678.

Zhang:2009:GOM

[ZW09] Shixue Zhang and Enhua Wu.

Generation of optimal multiresolution models for deforming mesh sequence. International Journal of Image and Graphics (IJIG), 9(2): 201–215, April 2009. CODEN ???? ISSN 0219-4678.

Zhiyu:2023:IIA

[ZWM23] Wang Zhiyu, Ding Weili, , and Wang Mingkui. Illumination invariance adaptive sidewalk detection based on unsupervised feature learning. International Journal of Image and Graphics (IJIG), 23(04): ??, July 2023. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467823500274.

Zhu:2008:MSS

[ZWZ08] Dengming Zhu, Zhaoqi Wang, and Yingping Zhang. Motion synthesis from the semantic signals. International Journal of Image and Graphics (IJIG), 8(4):535–550, October 2008. CODEN ???? ISSN 0219-4678.

Zhan:2006:FER

[ZYNC06] Yongzhao Zhan, Jingfu Ye, Dejiao Niu, and Peng Cao. Facial expression recognition based on Gabor wavelet transformation and elastic templates matching. International Journal of Image and Graphics (IJIG), 6(1):125–??, January 2006. CODEN ???? ISSN 0219-4678.

Zhu:2005:FMR

[ZYZH05] En Zhu, Jian-Ping Yin, Guo-Min Zhang, and Chun-Feng Hu. Fingerprint minutiae relationship representation and matching based on curve coordinate system. International Journal of Image and Graphics (IJIG), 5(4):729–??, October 2005. CODEN ???? ISSN 0219-4678.

Zhang:2018:ACF

[ZZ18] Jinpeng Zhang and Jinming Zhang. An analysis of CNN feature extractor based on KL divergence. International Journal of Image and Graphics (IJIG), 18(3): ??, July 2018. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467818500171.

Zheng:2006:SBA

[ZZWG06] Qing-Fang Zheng, Wei Zeng, Wei-Qiang Wang, and Wen Gao. Shape-based adult image detection. International Journal of Image and Graphics (IJIG), 6(1):115–??, January 2006. CODEN ???? ISSN 0219-4678.

Zhang:2022:FBC

[ZZY^{+}22] Geng Zhang, Qi Zhu, Jing Yang, Ruting Xu, Zhiqiang Zhang, and Daoqiang Zhang. Functional brain connectivity hyper-network embedded with structural information for epilepsy diagnosis. International Journal of Image and Graphics (IJIG), 22(04): ??, July 2022. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467822500292.